### v2

# Equilibrium Investment with Random Risk Aversion: (Non-)uniqueness, Optimality, and Comparative Statics

Abstract
This paper investigates infinite-dimensional portfolio selection problem under a general distribution of the risk aversion parameter. We provide a complete characterization of all deterministic equilibrium investment strategies. Our results reveal that the solution structure depends critically on the distribution of risk aversion: the equilibrium is unique whenever it exists in the case of finite expected risk aversion, whereas an infinite expectation can lead to infinitely many equilibria or to a unique trivial one ($\bar{\pi}\equiv\mathbf{0}$). To address this multiplicity, we introduce three optimality criteria—optimal, uniformly optimal, and uniformly strictly optimal—and explicitly characterize the existence and uniqueness of the corresponding equilibria. Under the same necessary and sufficient condition, the optimal and uniformly optimal equilibria exist uniquely and coincide. Furthermore, assuming that the market price of risk is non-zero near the terminal time, we show that the optimal (and hence uniformly optimal) equilibrium is also uniformly strictly optimal. Finally, we perform comparative statics to demonstrate that a risk aversion distribution dominating another in the reverse hazard rate order leads to a less aggressive equilibrium strategy.
Keywords: Random Risk Aversion; Time-Inconsistency; Portfolio Selection; Intra-Personal Equilibrium; Multiple Equilibria; Comparative Statics
1 Introduction
The classical framework for portfolio selection, pioneered by merton1969lifetime; merton1971 and samuelson1969lifetime, is based on the assumption that an investor’s risk aversion can be captured by a single known constant. In practice, however, this coefficient is very difficult for an agent to specify with certainty, and a growing body of empirical evidence suggests that it varies dynamically with market conditions, often increasing during financial crises and bear markets, as shown in studies by gordon2000preference; chetty2006new and guiso2018time. These challenges have motivated a move towards models that incorporate random risk aversion. Within this paradigm, a more economically coherent approach, as advanced by Desmettre2023, is to aggregate certainty equivalents over the distribution of the risk aversion parameter. While the random risk aversion in Desmettre2023 provides a consistent performance measure in monetary units, it introduces a nonlinearity in expectation that naturally leads to the time inconsistency of the preference.
The first systematic treatment of time inconsistency was given by Strotz1955, who proposed the notion of consistent planning, laying the foundation for the intra-personal game-theoretic framework. Building on this idea, ekeland2006being later developed a rigorous continuous-time formalization, which in turn inspired extensive research on continuous-time control with time-inconsistent preferences. For example, bjork2010general (published version Bjork2017) established a general theoretical framework by deriving an extended HJB equation and hu2012time; Hu2017 investigated a time-inconsistent stochastic linear–quadratic control problem. For further discussions and related developments, see, e.g., yan2019time, he2021equilibrium, Hernandez2023, and the references therein.
Within the Black-Scholes market model, Desmettre2023 provide a verification theorem on the intra-personal equilibrium strategies for a CRRA utility function with a general random risk aversion and, in the case of a binomial random risk aversion, characterize the equilibrium by a three-dimensional ODE system without establishing the existence and uniqueness of the solution. [^1] This work has spurred further research extending the model to more complex settings. For example, chen2025equilibrium incorporate a regime-switching framework in which market dynamics and preferences co-evolve; WANG2025103140 study the management of the DC pension plan, providing semi-explicit solutions for a CRRA utility with a binomial random risk aversion; he2025dynamic establish the existence and uniqueness of the solution to an infinite-dimensional Riccati equation which characterizes the equilibrium strategies under Heston’s SV model for a CRRA or CARA utility function with a general bounded random risk aversion. For an investigation on the corresponding pre-commitment problem, see, e.g., xia2024optimal. This technical gap is filled by liang2025short. For general random risk aversion, liang2025short characterize intra-personal equilibrium strategies using an integral equation. They establish the existence and uniqueness of the solution to the integral equation under the assumptions that the variance of the random risk aversion is finite and that the function $h$ defined in (2.6) below is bounded.
This paper aims to offer more insight into the portfolio selection problem with random risk aversion within the simplest model, in line with Desmettre2023 and liang2025short. Our main contributions are summarized as follows:
  • Complete characterization of all deterministic equilibria. We relax the assumptions in the existing literature and derive a full characterization of all deterministic equilibrium strategies. An important finding is that there may be infinitely many equilibria when the expectation of the random risk aversion is infinite. The uniqueness of the equilibrium strategy for the continuous-time time-inconsistent control problem is a very challenging problem (see, e.g., hjz2021). This paper sheds light on this issue by demonstrating that equilibrium strategies, even within the class of deterministic strategies, may lack uniqueness. Furthermore, we provide a necessary and sufficient condition for the uniqueness of the deterministic equilibrium.
  • Resolution of equilibrium multiplicity via the optimality criterion. A question then arises: which one should the agent choose from these infinitely many equilibrium strategies? We suggest that the agent should maximize the objective functional at the initial time over all available equilibria. This leads to the discussion of the so-called optimal equilibria. To our best knowledge, there is only one existing paper, wei2024time, on optimal equilibria for a time-inconsistent control problem, and there are three papers, hz2019; hz2020mf and hw2021, on optimal equilibria for time-inconsistent stopping problems. We give a necessary and sufficient condition for the existence of optimal equilibria and observe that the optimal equilibrium, if exists, is unique and is also uniformly optimal. Moreover, if the market price of risk remains non-zero near the terminal time, the optimal equilibrium is also uniformly strictly optimal.
  • Comparative statics on equilibrium strategies. We examine how the distribution of the random risk aversion influences the equilibrium strategies. It is natural to conjecture that a “larger” random risk aversion would lead to less risky investments. We observe by counterexamples that the “largeness” in the sense of the first-order stochastic dominance is insufficient to validate the conjecture. However, a stronger stochastic order, the reverse hazard rate dominance, is sufficient. Thus we get an analogy with Borell2007 and xia2011risk, where the comparative statics on the optimal portfolio with respect to risk aversion are carried out for the standard expected utility maximization with a general utility function and within the Black-Scholes market model (The corresponding comparative statics for a static model with a risk-free asset and a risky asset goes back to Arrow1970 and Pratt1964:RiskAversion).
The paper is organized as follows. Section 2 presents the model and derives the integral equation characterizing the equilibrium. Section 3 analyzes both the finite and infinite expectation cases, providing a complete characterization of all equilibria. Section 4 introduces three optimality criteria for equilibrium selection and establishes the necessary and sufficient conditions for their existence and uniqueness. Section 5 performs comparative statics to investigate the impact of risk aversion distributions on equilibrium strategies, establishing monotonicity under the reverse hazard rate order, and provides numerical illustrations that verify the theoretical results and present a counterexample under first-order stochastic dominance. [^2] Herein, “increasing” means “non-decreasing” and “decreasing” means “non-increasing.”
2 Problem Formulation
Let $T>0$ be a finite time horizon and $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ a filtered complete probability space, where $\mathbb{F}=\{\mathcal{F}_{t},\ 0\leq t\leq T\}$ is the augmented natural filtration generated by a $d$-dimensional standard Brownian motion $\{B(t),\ 0\leq t\leq T\}$ and $\mathcal{F}=\mathcal{F}_{T}$. The market consists of one risk-free asset (bond) and $d$ risky assets (stocks). For simplicity, we assume that the interest rate of the bond is zero. The dynamics of the stock price processes $S_{i}$, $i=1,\cdots,d$, are given by
$dS_{i}(t)=S_{i}(t)[\mu_{i}(t)dt+\sigma_{i}(t)dB(t)],\quad t\in[0,T],$
where the market coefficients $\mu:[0,T]\rightarrow\mathbb{R}^{d}$ and $\sigma:[0,T]\rightarrow\mathbb{R}^{d\times d}$ are right-continuous and deterministic (each $\sigma_{i}$ denotes the $i$-th row of $\sigma$). We always assume that
$\int_{0}^{T}|\mu(t)|dt+\sum_{i=1}^{d}\int_{0}^{T}|\sigma_{i}(t)|^{2}dt<\infty$
and $\sigma(t)$ is invertible for every $t\in[0,T]$. The market price of risk is $\lambda(t)\triangleq(\sigma(t))^{-1}\mu(t)$, $t\in[0,T]$. We also assume that $\int_{0}^{T}|\lambda(t)|^{2}\mathrm{d}t<\infty$.
For any $m\geq 1$ and $\mathbb{S}\subset\mathbb{R}^{m}$, $L^{0}(\mathbb{F},\mathbb{S})$ is the space of $\mathbb{S}$-valued, $\mathbb{F}$-progressively measurable processes. For each $t\in[0,T]$, $p\in[1,\infty]$, $L^{p}(\mathcal{F}_{t},\mathbb{S})$ is the set of all $L^{p}$-integrable, $\mathbb{S}$-valued, and $\mathcal{F}_{t}$-measurable random variables.
A trading strategy is a process $\pi=\{\pi_{t},\ t\in[0,T)\}\in L^{0}(\mathbb{F},\mathbb{R}^{d})$ such that
$\int_{0}^{T}|\pi^{\top}_{t}\mu(t)|\mathrm{d}t+\int_{0}^{T}|\sigma^{\top}(t)\pi_{t}|^{2}\mathrm{d}t<\infty\text{ a.s.,}$
where $\pi_{t}$ represents the vector of portfolio weights determining the investment of wealth in the stocks at time $t$. For any $t\in[0,T)$ and any initial wealth level $w>0$, let $W^{t,w,\pi}=\{W^{t,w,\pi}_{s},\ t\leq s\leq T\}$ denote the self-financing wealth process starting from $w$ at time $t$ and driven by a trading strategy $\pi$. It evolves according to the stochastic differential equation (SDE):
$\left\{\begin{aligned} &\mathrm{d}W^{t,w,\pi}_{s}=W^{t,w,\pi}_{s}\pi^{\top}_{s}\mu(s)\mathrm{d}s+W^{t,w,\pi}_{s}\pi^{\top}_{s}\sigma(s)\mathrm{d}B(s),\quad s\geq t,\\ &W^{t,w,\pi}_{t}=w.\end{aligned}\right$. (2.1)
Consider the constant relative risk aversion (CRRA) utility functions:
$u^{\gamma}(x)=\begin{cases}\frac{x^{1-\gamma}}{1-\gamma},&\gamma>0,\gamma\neq 1,x>0,\\ \log x,&\gamma=1,x>0.\end{cases}$
Under the expected utility theory, for each CRRA utility function $u^{\gamma}$, at time $t$, the certainty equivalent of $W^{t,w,\pi}_{T}$ is $\left(u^{\gamma}\right)^{-1}\left(\mathbb{E}_{t}\left[u^{\gamma}\left(W^{t,w,\pi}_{T}\right)\right]\right)$. Following Desmettre2023, we assume that the objective functional of the agent at time $t$ is the weighted average of the certainty equivalents:
$J\left(\pi;t,w\right)=\int_{[0,\infty)}\left(u^{\gamma}\right)^{-1}\left(\mathbb{E}_{t}\left[u^{\gamma}\left(W^{t,w,\pi}_{T}\right)\right]\right)\mathrm{d}F_{\bm{R}}(\gamma),$ (2.2)
[…]
A trading strategy $\pi$ is called admissible if for all $t\in[0,T)$ and $w>0$,
$\begin{cases}\mathbb{E}_{t}\left[\left|u^{\gamma}\left(W^{t,w,\pi}_{T}\right)\right|\right]<\infty\,\ a.s.\quad\text{ for all }\gamma\in\mathrm{supp}(F_{\bm{R}}),\\ \int_{[0,\infty)}\left|\left(u^{\gamma}\right)^{-1}\left(\mathbb{E}_{t}\left[u^{\gamma}\left(W^{t,w,\pi}_{T}\right)\right]\right)\right|\mathrm{d}F_{\bm{R}}(\gamma)<\infty\,\ \ a.s..\end{cases}$
Let $\Pi$ denote the set of all admissible strategies.
Hereafter, we consider a fixed $\bar{\pi}\in\Pi$, which is a candidate equilibrium strategy. For any $t\in[0,T)$, $\varepsilon\in(0,T-t)$, and $k\in L^{\infty}(\mathcal{F}_{t},\mathbb{R}^{d})$, let $\pi^{t,\varepsilon,k}$ be defined by
$\pi^{t,\varepsilon,k}_{s}\triangleq\left\{\begin{aligned} &{\bar{\pi}_{s}+k},&s\in[t,t+\varepsilon),\\ &\bar{\pi}_{s},&s\notin[t,t+\varepsilon).\end{aligned}\right$.
The strategy $\pi^{t,\varepsilon,k}$ serves as a perturbation of $\bar{\pi}$. Following hu2012time; Hu2017 and yan2019time, we introduce the definition of equilibrium strategies as follows.
Definition 2.1.
A strategy $\bar{\pi}$ is called an equilibrium strategy if for any $t\in[0,T)$, $w>0$, and $k\in L^{\infty}(\mathcal{F}_{t},\mathbb{R}^{d})$, whenever $\pi^{t,\varepsilon,k}\in\Pi$ for all sufficiently small $\varepsilon>0$, and for any positive sequence $\{\varepsilon_{n},n\geq 1\}$ satisfying $\lim\limits_{n\to\infty}\varepsilon_{n}=0$ we have
$\limsup_{n\to\infty}\frac{J(\pi^{t,\varepsilon_{n},k};t,w)-J(\bar{\pi};t,w)}{\varepsilon_{n}}\leq 0\quad\text{a.s.}.$ (2.3)
Remark 2.2.
In literature, the inequality is usually
$\limsup_{\varepsilon\downarrow 0}\frac{J(\pi^{t,\varepsilon,k};t,w)-J(\bar{\pi};t,w)}{\varepsilon}\leq 0\quad\text{a.s.}.$
However, the left-hand side of the inequality might be unmeasurable. To guarantee the measurability, we modify the definition of equilibrium accordingly.
Remark 2.3.
Note that
$J(\pi;t,w)$ $=w\int_{[0,\infty)}\left(u^{\gamma}\right)^{-1}\left(\mathbb{E}_{t}\left[u^{\gamma}\left({W^{t,w,\pi}_{T}}/w\right)\right]\right)\mathrm{d}F_{\bm{R}}(\gamma)=wJ_{0}(t,\pi),$
where
$J_{0}(t,\pi)\triangleq\int_{[0,\infty)}\left(u^{\gamma}\right)^{-1}\left(\tilde{\mathbb{E}}_{t}\left[u^{\gamma}\left({W^{\pi}_{T}/W^{\pi}_{t}}\right)\right]\right)\mathrm{d}F_{\bm{R}}(\gamma)$
is the objective functional used in liang2025short and $W^{\pi}$ evolves according to
$\mathrm{d}W^{\pi}_{t}=W^{\pi}_{t}\pi^{\top}_{t}\mu(t)\mathrm{d}t+W^{\pi}_{t}\pi^{\top}_{t}\sigma(t)\mathrm{d}B(t),\quad W^{\pi}_{0}=w_{0}>0.$
Thus, under Definition 2.1, an equilibrium strategy for $J$ is also an equilibrium strategy for $J_{0}$, and vice versa. We will frequently use $J_{0}$ in the subsequent analysis.
In our analysis, we follow the approach of liang2025short and focus on the equilibrium strategies given by $\bar{\pi}=(\sigma^{\top})^{-1}a$, where $a$, referred to as the risk exposure vector, is a deterministic and right-continuous funcion in the $L^{2}$ space. Let $\Pi_{d}\subset\Pi$ denote the set of all such strategies. Because $k\in L^{\infty}(\mathcal{F}_{t},\mathbb{R}^{d})$, the perturbed strategy $\pi^{t,\varepsilon,k}$ can be random, even though $\bar{\pi}\in\Pi_{d}$ itself is deterministic. The wealth process $W^{t,w,\bar{\pi}}$ of the portfolio $\bar{\pi}$ satisfies the following SDE
$\mathrm{d}W^{t,w,\bar{\pi}}_{s}=W^{t,w,\bar{\pi}}_{s}[a(s)^{\top}\lambda(s)\mathrm{d}s+a(s)^{\top}\mathrm{d}B(s)],\quad s\geq t,$
and $|a|$ is referred to as the risk exposure magnitude (or local volatility).
For any $\bar{\pi}=(\sigma^{\top})^{-1}a\in\Pi_{d}$, we introduce the following notations for ease of analysis:
$v_{a}(t)\triangleq\int_{t}^{T}|a(s)|^{2}ds,\quad y_{a}(t)\triangleq\int_{t}^{T}a^{\top}(s)\lambda(s)ds,\quad t\in[0,T].$
Because the strategy is deterministic, the conditional distribution of the relative wealth $W^{t,w,\bar{\pi}}_{T}/w$ depends only on the deterministic quantities $v_{a}(t)$ and $y_{a}(t)$ and we have
$J_{0}(t,\bar{\pi})=\tilde{\mathbb{E}}\left[\mathrm{e}^{-\frac{1}{2}\left(\bm{R}v_{a}(t)-2y_{a}(t)\right)}\right].$ (2.4)
The following theorem shows that $\pi\equiv\mathbf{0}$ is an equilibrium if $\tilde{\mathbb{E}}[\bm{R}]=\infty$. (This result is natural: if the random risk aversion is very large, then the investor puts no money into the risky asset. However, in Theorem 3.2(2) we will see that $\pi\equiv\mathbf{0}$ is not necessarily the unique equilibrium.)
Theorem 2.4.
$\bar{\pi}\equiv\mathbf{0}$ is an equilibrium if $\tilde{\mathbb{E}}[\bm{R}]=\infty$.
Proof.
Let $\bar{\pi}=\mathbf{0}$. For any $t\in[0,T)$, $\varepsilon\in(0,T-t)$, and $k\in L^{\infty}(\mathcal{F}_{t},\mathbb{R}^{d})$, we have $J_{0}(t,\bar{\pi})=1$ and
$J_{0}(t,\pi^{t,\varepsilon,k})=\mathrm{e}^{\int_{t}^{t+\varepsilon}\mu^{\top}(s)k\mathrm{d}s}\int_{[0,\infty)}\mathrm{e}^{-\frac{\gamma}{2}\int_{t}^{t+\varepsilon}|\sigma^{\top}(s)k|^{2}\mathrm{d}s}\mathrm{d}F_{\bm{R}}(\gamma)\triangleq\phi(\varepsilon)\psi(\varepsilon).$
Noting that
$\lim_{\varepsilon\to 0}\frac{\psi(\varepsilon)-1}{\varepsilon}$ $=-\frac{1}{2}|\sigma^{\top}(t)k|^{2}\lim_{\varepsilon\to 0}\int_{[0,\infty)}\gamma\mathrm{e}^{-\frac{\gamma}{2}\int_{t}^{t+\varepsilon}|\sigma^{\top}(s)k|^{2}\mathrm{d}s}\mathrm{d}F_{\bm{R}}(\gamma)$
$=-\frac{1}{2}|\sigma^{\top}(t)k|^{2}\int_{[0,\infty)}\gamma\mathrm{d}F_{\bm{R}}(\gamma),$
where the first equality uses the mean value theorem and the second the monotone convergence theorem. Thus, for any positive sequence $\{\varepsilon_{n},n\geq 1\}$ satisfying $\lim\limits_{n\to\infty}\varepsilon_{n}=0$, we have
$\lim_{n\to\infty}\frac{J_{0}(t,\pi^{t,\varepsilon_{n},k})-J_{0}(t,\bar{\pi})}{\varepsilon_{n}}$ $=\lim_{n\to\infty}\frac{\phi(\varepsilon_{n})\psi(\varepsilon_{n})-1}{\varepsilon_{n}}=\lim_{n\to\infty}\frac{(\phi(\varepsilon_{n})-1)\psi(\varepsilon_{n})+\psi(\varepsilon_{n})-1}{\varepsilon_{n}}$
$=\left(\mu^{\top}(t)k-\frac{1}{2}|\sigma^{\top}(t)k|^{2}\int_{[0,\infty)}\gamma\mathrm{d}F_{\bm{R}}(\gamma)\right).$
Therefore, $\mathbf{0}$ is an equilibrium if $\tilde{\mathbb{E}}[\bm{R}]=\infty$. ∎
The next theorem provides a necessary and sufficient condition for $\bar{\pi}=(\sigma^{\top})^{-1}a\in\Pi_{d}$ to be an equilibrium.
Theorem 2.5.
$\bar{\pi}=(\sigma^{\top})^{-1}a\in\Pi_{d}$ is an equilibrium if and only if
$a(t)=h(v_{a}(t))\lambda(t)$ (2.5)
holds for any $t\in[0,T)$, where [^3] This function $h$ is slightly different from that in liang2025short: $h(x)$ there amounts to $h(x^{2})$ here. Nevertheless, the relevant properties, such as the boundedness of $h$, remain equivalent under the two formulations.
$h(x)\triangleq{\tilde{\mathbb{E}}\left[\mathrm{e}^{-\frac{1}{2}\bm{R}x}\right]}\left/{\tilde{\mathbb{E}}\left[\bm{R}\mathrm{e}^{-\frac{1}{2}\bm{R}x}\right]}\right.,\quad x\in[0,\infty).$ (2.6)
(Occasionally, we write $h_{\bm{R}}$ for $h$ to emphasize its dependence on $\bm{R}$.)
Proof.
It suffices to show that for any $t\in[0,T)$, (2.3) is equivalent to (2.5).
When $\tilde{\mathbb{E}}[\bm{R}]<\infty$, we have $h(0)={1}/{\tilde{\mathbb{E}}[\bm{R}]}\in(0,\infty)$ and $h\in C([0,\infty))$. In this case, an argument similar to the proof of liang2025short yields the equivalence.
When $\tilde{\mathbb{E}}[\bm{R}]=\infty$, we have $h(0)=0$. In this case, similar to the case $\tilde{\mathbb{E}}[\bm{R}]<\infty$, (2.3) and (2.5) are equivalent for $t\in[0,T_{0})$, where $T_{0}=\inf\{t\in[0,T]:v_{a}(t)=0\}$. Moreover, they are automatically satisfied for $t\in[T_{0},T)$ based on Theorem 2.4 and $h(0)=0$. ∎
Now we introduce a function $l:[0,\infty)\to(0,\infty)$, which is defined by
$l(y)\triangleq\tilde{\mathbb{E}}[e^{-\bm{R}y}],\quad y\geq 0.$
The function $h$ defined in (2.6) plays a central role in our analysis. Here, we present two examples for which $h$ can be explicitly worked out.
Example 2.6.
Consider the following two distributions for the random risk aversion $\bm{R}$.
  1. (1)
    Suppose that $\bm{R}$ follows a Poisson distribution with parameter $\theta>0$. In this case, $l(y)=e^{\theta(e^{-y}-1)}$ for any $y\geq 0$. According to (2.6), we obtain
    $h(x)=-{l({x}/{2})}/{l^{\prime}({x}/{2})}=\frac{1}{\theta}e^{{x}/{2}},\quad x\geq 0.$
  2. (2)
    Suppose that $\bm{R}$ follows a Gamma distribution with shape parameter $\alpha>0$ and scale parameter $\beta>0$. Its probability density function is
    $f_{\bm{R}}(r)=\frac{r^{\alpha-1}e^{-{r}/{\beta}}}{\Gamma(\alpha)\beta^{\alpha}},\quad r>0.$
    In this case, $l(y)=\tilde{\mathbb{E}}[e^{-\bm{R}y}]=(1+\beta y)^{-\alpha}$ ($y\geq 0$), which gives
    $h(x)=-\frac{l(\frac{x}{2})}{l^{\prime}(\frac{x}{2})}=\frac{1}{\alpha\beta}\left(1+\frac{\beta}{2}x\right),\quad x\geq 0.$
In the two cases of the above example, the functions $h$ are unbounded, violating the boundedness assumption on $h$ made by liang2025short. Moreover, as we will see in Example 3.5 below, the equilibrium strategies can be explicitly obtained for both cases. To accommodate such interesting cases, we will conduct a thorough analysis on Condition (2.5) and the equilibrium strategies under a general distribution of $\bm{R}$.
3 Equilibrium Strategy
We first transform Condition (2.5) into an ordinary differential equation (ODE) for $v_{a}$. Differentiating $v_{a}(t)=\int_{t}^{T}|a(s)|^{2}ds$ with respect to $t$, we get from (2.5) that $v_{a}$ satisfies the following ODE:
$\begin{cases}v^{\prime}(t)=-h^{2}(v(t))|\lambda(t)|^{2},&t\in\left[0,T\right),\\ v(T)=0.\end{cases}$ (3.1)
Conversely, assume that $v$ solves the ODE (3.1) and let $a(\cdot)=h(v(\cdot))\lambda(\cdot)$. Then $a$ satisfies (2.5).
The ODE (3.1) is an equation with separated variables. We can obtain its solutions in closed form.
Now we introduce two functions. The first function $\Lambda$ is defined by
$\Lambda(t)\triangleq\int_{t}^{T}|\lambda(s)|^{2}\mathrm{d}s,\quad t\in[0,T].$
The second function $H$ is defined by
$H(y)\triangleq\int_{0}^{y}\frac{1}{h^{2}(x)}\mathrm{d}x=\int_{0}^{y}\left({\tilde{\mathbb{E}}\left[\bm{R}\mathrm{e}^{-\frac{1}{2}\bm{R}x}\right]}\left/{\tilde{\mathbb{E}}\left[\mathrm{e}^{-\frac{1}{2}\bm{R}x}\right]}\right.\right)^{2}\mathrm{d}x,\quad y\in[0,\infty].$
Occasionally, we write $H_{\bm{R}}$ for $H$ to emphasize its dependence on $\bm{R}$.
If $\Lambda(0)=0$, then $\lambda(\cdot)\equiv 0$ on $[0,T)$ and, in view of (2.5), there is a unique equilibrium $\pi\equiv\mathbf{0}$. This trivial case is excluded by making the following standing assumption.
Assumption 3.1.
$\Lambda(0)\in(0,\infty)$.
The next theorem provides, in closed form, the solutions to the ODE (3.1) and the equilibrium strategies in $\Pi_{d}$ .
Theorem 3.2.
Under Assumption 3.1, we have the following assertions.
  1. (1)
[…]
  2. (2)
    If $\tilde{\mathbb{E}}[\bm{R}]=\infty$ and $H(\varepsilon)<\infty$ for all $\varepsilon>0$, the ODE (3.1) admits a one-parameter family of solutions, indexed by $T_{0}\in\mathcal{T}$:
    $v^{(T_{0})}(t)\triangleq\begin{cases}H^{-1}\left(\int_{t}^{T_{0}}|\lambda(s)|^{2}\mathrm{d}s\right),&0\leq t<T_{0},\\ 0,&T_{0}\leq t\leq T,\end{cases}$ (3.2)
    where
    $\mathcal{T}\triangleq\left\{T_{0}\in[0,T]\;\bigg|\;H(\infty)>\Lambda(0)-\Lambda(T_{0})\text{ and }\Lambda(t)>\Lambda(T_{0})\text{ for all }t\in[0,T_{0})\right\}.$
    Then the set of all equilibrium strategies in $\Pi_{d}$ is :
    $\Pi^{e}_{d}=\left\{\pi^{(T_{0})}(\cdot)\triangleq(\sigma^{\top}(\cdot))^{-1}h(v^{(T_{0})}(\cdot))\lambda(\cdot):T_{0}\in\mathcal{T}\right\}.$
  3. (3)
[…]
Proof.
Assume $\tilde{\mathbb{E}}[\bm{R}]<\infty$. In this case, $h(x)>0$ for all $x\geq 0$. Suppose that $v(\cdot)$ is a solution to the ODE (3.1). Separating variables in (3.1) yields
$-|\lambda(t)|^{2}=\frac{v’(t)}{h^{2}(v(t))},\quad t\in[0,T).$ (3.3)
Integrating (3.3) from $t$ to $T$ yields
$-\Lambda(t)=\int_{t}^{T}\frac{v’(s)}{h^{2}(v(s))}\mathrm{d}s=\int_{v(t)}^{0}\frac{\mathrm{d}y}{h^{2}(y)}=-\int_{0}^{v(t)}\frac{\mathrm{d}y}{h^{2}(y)}=-H(v(t)),$
where the second equality follows from the change of variables $y=v(s)$. Thus, $H(v(t))=\Lambda(t)$ for all $t\in[0,T]$. As $h(0)={1}/{\tilde{\mathbb{E}}[\bm{R}]}>0$, it follows that $H(0)=0$ and $H(\cdot)\in{C}^{1}\left([0,+\infty)\right)$ is strictly increasing. Thus, the inverse function $H^{-1}$ is well-defined, continuous, and strictly increasing in $\left[0,H(\infty)\right)$. Thus, $H(\infty)>H(v(0))=\Lambda(0)$ and $v(t)=H^{-1}(\Lambda(t))$ for all $t\in[0,T]$, which yields the uniqueness of the solution. Conversely, if $H(\infty)>\Lambda(0)$, it is straightforward to verify that $v(\cdot)=H^{-1}(\Lambda(\cdot))$ indeed solves uniquely the ODE.
Assume that $\tilde{\mathbb{E}}[\bm{R}]=\infty$ and $H(\varepsilon)<\infty$ for all $\varepsilon>0$. We can see that $H(0)=0$, $H\in C^{0}([0,+\infty])\cap C^{1}((0,+\infty))$, and $H$ is strictly increasing. Suppose that $v(\cdot)$ is a solution to the ODE (3.1). Let $T_{0}=\inf\{t\in[0,T]:v(t)=0\}$. Then $v(t)=0$ for any $t\in[T_{0},T]$ and $v(t)>0$ for any $t\in[0,T_{0})$. For any $t\in[0,T_{0})$, integrating (3.3) from $t$ to $T_{0}$ yields
$\Lambda(T_{0})-\Lambda(t)=\int_{t}^{T_{0}}-|\lambda(s)|^{2}\mathrm{d}s=\int_{v(t)}^{v(T_{0})}\frac{\mathrm{d}y}{h^{2}(y)}=-H(v(t)).$
Thus, $H(v(t))=\Lambda(t)-\Lambda(T_{0})$ for all $t\in[0,T_{0})$. Therefore, $H(\infty)>H(v(0))=\Lambda(0)-\Lambda(T_{0})$ and $\Lambda(t)-\Lambda(T_{0})=H(v(t))>H(0)=0$ for all $t\in[0,T_{0})$. Consequently, $T_{0}\in\mathcal{T}$. Moreover, for any $t<T_{0}$, we have
$v(t)=H^{-1}(\Lambda(t)-\Lambda(T_{0}))=H^{-1}\left(\int_{t}^{T_{0}}|\lambda(s)|^{2}\mathrm{d}s\right),$
which implies $v=v^{(T_{0})}$. Conversely, a direct verification shows that (3.2) indeed solves the ODE.
Assume that $\tilde{\mathbb{E}}[\bm{R}]=\infty$ and $H(\varepsilon)=\infty$ for some $\varepsilon>0$. This implies that $\int_{0}^{y}\frac{1}{h^{2}(x)}\mathrm{d}x=\infty$ for any $y>0$. Clearly, $v\equiv 0$ is a solution (recall that $h(0)={1}/{\tilde{\mathbb{E}}[\bm{R}]}=0$). Suppose, on the contrary, that there exists another solution $v(\cdot)$ with $v(t)>0$ for some $t<T$. Let $t^{\prime}\triangleq\inf\{s\in(t,T]:v(s)=0\}$. By continuity, we have $v(t^{\prime})=0$ and $v(s)>0$ for any $s\in[t,t^{\prime})$. Integrating the ODE (3.3) from $t$ to $t^{\prime}$ yields
$\Lambda(t^{\prime})-\Lambda(t)=\int_{t}^{t^{\prime}}-|\lambda(s)|^{2}\mathrm{d}s=\int_{v(t)}^{0}\frac{\mathrm{d}y}{h^{2}(y)}=-\infty,$
which is impossible. Therefore, the ODE (3.1) admits the unique solution $v\equiv 0$. ∎
Assertion (2) of Theorem 3.2 characterizes the solutions to the ODE (3.1) via the parameter $T_{0}\in\mathcal{T}$. The following lemma provides an alternative explicit characterization of the set $\mathcal{T}$ via an auxiliary mapping.
Lemma 3.3.
Let $\mathcal{A}\triangleq[0,\Lambda(0)]\cap(\Lambda(0)-H(\infty),\Lambda(0)]$ and
$\quad\varphi(\eta)\triangleq\min\{t\in[0,T]\mid\Lambda(t)=\eta\},\quad\eta\in[0,\Lambda(0)].$
Then we have $\mathcal{T}=\{\varphi(\eta)\mid\eta\in\mathcal{A}\}$.
Proof.
From the definition of $\mathcal{T}$, for any $T_{0}\in\mathcal{T}$, we have $\Lambda(T_{0})\in\mathcal{A}$ and
$T_{0}=\inf\{t\in[0,T]:\Lambda(t)=\Lambda(T_{0})\}=\varphi(\Lambda(T_{0})).$
Thus, $\mathcal{T}\subset\{\varphi(\eta)\mid\eta\in\mathcal{A}\}$. Conversely, for any $\eta\in\mathcal{A}$, the continuity of $\Lambda$ implies $\Lambda(\varphi(\eta))=\eta\in\mathcal{A}$. Moreover, for any $t\in[0,\varphi(\eta))$, we have $\Lambda(t)>\Lambda(\varphi(\eta))$ by the definition of $\varphi$. Therefore, $\varphi(\eta)\in\mathcal{T}$ and hence $\{\varphi(\eta)\mid\eta\in\mathcal{A}\}\subset\mathcal{T}$. ∎
Remark 3.4.
Because $\varphi$ is strictly decreasing, the sets $\mathcal{T}$ and $\mathcal{A}$ are in one-to-one correspondence. It is easily seen that $\mathcal{A}$ contains infinitely many elements. Consequently, under the conditions $\tilde{\mathbb{E}}[\bm{R}]=\infty$ and $H(\varepsilon)<\infty$ for all $\varepsilon>0$, we obtain infinitely many equilibria.
It should be noted that the existing studies, such as Desmettre2023 and liang2025short, focus mainly on the specific cases covered by Theorem 3.2(1), under the additional assumptions that $h(\cdot)$ is bounded and $\bm{R}$ is square-integrable. Our result covers any general random risk aversion $\bm{R}$. In the following two examples, the functions $h$ are unbounded and the equilibrium strategies are explicitly worked out.
Example 3.5.
This example derives the explicit equilibrium strategies for the two distributions discussed in Example 2.6.
  1. (1)
    Poisson distribution: Recall that $h(x)=\dfrac{1}{\theta}e^{\frac{x}{2}}$. Then $h$ is unbounded and
    $H(y)=\int_{0}^{y}\frac{1}{h^{2}(x)}dx=\theta^{2}(1-e^{-y}),\quad y\in[0,\infty),$
    with $H(\infty)=\theta^{2}$. If $\Lambda(0)<\theta^{2}$, then $v(t)=H^{-1}(\Lambda(t))=-\ln\left(1-\dfrac{\Lambda(t)}{\theta^{2}}\right)$. In this case, the uinque equilibrium strategy is explicitly given by
    $\bar{\pi}(t)=(\sigma^{\top})(t)^{-1}a(t)=(\sigma^{\top}(t))^{-1}h(v(t))\lambda(t)=\dfrac{(\sigma^{\top}(t))^{-1}\lambda(t)}{\sqrt{\theta^{2}-\Lambda(t)}},\quad t\in[0,T).$
  2. (2)
    Gamma distribution: Recall that $h(x)=\dfrac{1}{\alpha\beta}(1+\dfrac{\beta}{2}x)$. Then $h$ is unbounded and
    $H(y)=\int_{0}^{y}\frac{\alpha^{2}\beta^{2}}{(1+\frac{\beta}{2}x)^{2}}dx=2\alpha^{2}\beta\left(1-\dfrac{1}{1+\frac{\beta}{2}y}\right),\quad y\in[0,\infty),$
    with $H(\infty)=2\alpha^{2}\beta$. Provided that $\Lambda(0)<2\alpha^{2}\beta$, we obtain $v(t)=\dfrac{2}{\beta}\left[\left(1-\dfrac{\Lambda(t)}{2\alpha^{2}\beta}\right)^{-1}-1\right]$. In this case, the unique equilibrium strategy is explicitly given by
    $\bar{\pi}(t)=(\sigma^{\top}(t))^{-1}a(t)=\dfrac{(\sigma^{\top}(t))^{-1}\lambda(t)}{\alpha\beta-\frac{\Lambda(t)}{2\alpha}},\quad t\in[0,T).$
Example 3.6.
This example is provided for assertions (2) and (3) in Theorem 3.2. Let $\bm{R}_{\alpha}\sim G_{\alpha}$, where $\alpha\in(0,1)$ is a fixed parameter and $G_{\alpha}$ denotes the distribution function whose Laplace transform satisfies [^4] See Pollard1946 and Feller1971.
$\int_{0}^{\infty}\mathrm{e}^{-vx}\mathrm{d}G_{\alpha}(x)=\mathrm{e}^{-v^{\alpha}},\quad v\geq 0.$
Note that $\tilde{\mathbb{E}}[\bm{R}_{\alpha}]=\infty$ for $\alpha\in(0,1)$. Moreover, $h(x)=\frac{1}{\alpha}\left(\frac{x}{2}\right)^{1-\alpha}$, $x\geq 0$. Obviously, $h$ is unbounded. For simplicity, we assume that $d=1$ and $\sigma(\cdot)\equiv\sigma\neq 0$ and $\lambda(\cdot)\equiv\lambda\neq 0$ are constants in $\mathbb{R}$.
  • Let $\alpha\in(0.5,1)$. By direct calculation, the solutions to the ODE (3.1) take the form
    $v^{(T_{0})}(t)=\begin{cases}\left(2^{2\alpha-2}\dfrac{1}{\alpha^{2}}(2\alpha-1)\lambda^{2}(T_{0}-t)\right)^{\frac{1}{2\alpha-1}},&t\in[0,T_{0}),\\ 0,&t\in[T_{0},T],\end{cases}$
    where $T_{0}\in[0,T]$. Consequently, the explicit expression for the equilibrium strategy $\pi^{(T_{0})}(t)$ on $[0,T)$ is
[…]
    $=\begin{cases}(\sigma)^{-1}\left(\frac{\lambda}{\alpha}\right)^{\frac{1}{2\alpha-1}}\left[\frac{2\alpha-1}{2}(T_{0}-t)\right]^{\frac{1-\alpha}{2\alpha-1}},&0\leq t<T_{0},\\ 0,&T_{0}\leq t<T,\end{cases}$
    where $T_{0}\in[0,T]$.
  • Let $\alpha\in(0,0.5]$. Direct calculation shows that the ODE (3.1) admits only the trivial solution $v\equiv 0$ and therefore $\pi=0$ is the unique equilibrium.
Assume $\tilde{\mathbb{E}}[\bm{R}]<\infty$ and $H(\infty)>\Lambda(0)$. Assertion (1) of Theorem 3.2 shows that the ODE (3.1) admits a unique solution $v$ and therefore there exists a unique function $a$ such that (2.5) holds for all $t\in[0,T)$. To emphasize the dependence on $\bm{R}$, we occasionally write $v_{\bm{R}}$ and $a_{\bm{R}}$ for $v$ and $a$, respectively.
Remark 3.7.
The explicit characterization of the equilibrium strategy in Theorem 3.2(1) yields a clear structural decomposition of the risk exposure vector, given by
$a(t)=h(H^{-1}(\Lambda(t)))\lambda(t),\quad t\in[0,T).$ (3.4)
Note that the definitions of the functions $h(\cdot)$ and $H(\cdot)$ depend solely on the distribution of the random risk aversion $\bm{R}$}. The representation (3.4) shows that, at time $t$, the effects of three distinct factors on the risk exposure vector are cleanly separated: the instantaneous market opportunity $\lambda(t)$, the aggregate future market opportunities $\Lambda(t)$ over the remaining interval $(t,T)$, and the random risk aversion preference itself.
More specifically, at time $t$, if $\Lambda(t)$ and $\bm{R}$} are fixed, a larger current risk premium $|\lambda(t)|$ leads to a larger $|a(t)|$. If $|\lambda(t)|$ and $\bm{R}$} are fixed, then a larger future market opportunity $\Lambda(t)$ implies a larger value of $H^{-1}(\Lambda(t))$, and hence a larger value of $h\big(H^{-1}(\Lambda(t))\big)$ (see Proposition 3.8 for a detailed discussion; see alos Lemma A.1 for the monotonicity of $h$). Finally, if both $\lambda(t)$ and $\Lambda(t)$ are fixed, one can study how $\bm{R}$} affects $h$ and $H$, and thereby its impact on $a(t)$; see the comparative statics analysis in Section 5.
We emphasize that, in contrast to the classical CRRA utility, which generates a “myopic” strategy reacting primarily to the instantaneous market price of risk $\lambda(t)$, the equilibrium under random risk aversion is inherently forward-looking due to the dependence on $\Lambda(t)$. Finally, we refer to liang2024dynamic and liang2025dynamic for similar structural representations, although monotonicity properties and comparative statics are not investigated therein.
Next, we study the impact of changes in future market opportunities on the risk exposure vector.
Proposition 3.8.
Let $\bm{R}$ be a non-negative random variable with $\tilde{\mathbb{E}}[\bm{R}]\in(0,+\infty)$. Fix a vector $\lambda_{0}\in\mathbb{R}^{d}$. Consider a sequence of market settings indexed by $n\in\mathbb{N}$, defined by the pairs $\{(\lambda_{n}(\cdot),T_{n})\}_{n\geq 1}$, where $T_{n}>0$ and $\lambda_{n}:[0,T_{n}]\to\mathbb{R}^{d}$ satisfying $\lambda_{n}(0)=\lambda_{0}$. Let $\Lambda_{n}\triangleq\int_{0}^{T_{n}}|\lambda_{n}(s)|^{2}ds$. Let $a_{n}(0)$ denote the risk exposure vector at time 0 corresponding to the $n$-th market setting.
  1. (1)
    If the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly increasing and $\Lambda_{n}\nearrow H(\infty)$, then
    $\lim_{n\to\infty}a_{n}(0)=\frac{\lambda_{0}}{r_{0}}\quad\text{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(if $r_{0}>0$)},$
    where $r_{0}\triangleq\operatorname{essinf}\bm{R}$. Moreover, $|a_{n}(0)|$ converges monotonically increasing, i.e.,
    $|a_{n}(0)|\nearrow\frac{|\lambda_{0}|}{r_{0}}.$
  2. (2)
    If the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly decreasing and $\Lambda_{n}\searrow 0$, then
    $\lim_{n\to\infty}a_{n}(0)=\frac{\lambda_{0}}{\tilde{\mathbb{E}}[\bm{R}]}.$
    Moreover, $|a_{n}(0)|$ converges monotonically decreasing, i.e.,
    $|a_{n}(0)|\searrow\frac{|\lambda_{0}|}{\tilde{\mathbb{E}}[\bm{R}]}.$
Here, we adopt the convention that $\frac{1}{0}=\infty$.
Proof.
See Appendix B.1
Remark 3.9.
The convergence results in Proposition 3.8 depend solely on the sequence of integral values $\{\Lambda_{n}\}_{n\geq 1}$, and no convergence of the parameter sequence $\{(\lambda_{n}(\cdot),T_{n})\}_{n\geq 1}$ itself is required. Moreover, focusing on time 0 entails no loss of generality. For any arbitrary time $t$, analogous conclusions follow by a simple time translation, with $t$ treated as the initial time.
Remark 3.10.
In the single-stock case, assuming that $\lambda>0$ and $\sigma>0$ are constants, one can use $\hat{\pi}(\cdot)=\frac{1}{\sigma}h(v(\cdot))\lambda$, together with the monotonicity of $h$ given in Lemma A.1, to establish a favorable property: $\pi(\cdot)$ is decreasing. See Section 4.3 of Desmettre2023 for a detailed discussion. However, the results in Desmettre2023 do not yield Proposition 3.8. The proof of Proposition 3.8 crucially relies on our explicit characterization of the risk exposure vector, given in (3.4), which is not available in their framework.
As the final part of this section, we use the explicit characterization of the equilibria in Theorem 3.2 to investigate the continuity properties of $a(\cdot)$ with respect to the distribution of the random risk aversion. More specifically, the following theorem shows that if a sequence of random risk aversion $\{R_{n}\}_{n\geq 1}$ converges in distribution to $R$, then the corresponding sequence $\{a_{n}\}_{n\geq 1}$ of volatility converges to $a$.
Theorem 3.11.
Let $\{\bm{R}_{n}\}_{n\geq 1}$ be a sequence of random risk aversions that converges in distribution to $\bm{R}$ and $\lim\limits_{n\to\infty}\tilde{\mathbb{E}}[\bm{R}_{n}]=\tilde{\mathbb{E}}[\bm{R}]\in(0,\infty)$. Assume that $H_{n}(\infty)>\Lambda(0)$ for all $n\geq 1$ and $H(\infty)>\Lambda(0)$. Let $a_{n}=a_{\bm{R}_{n}}$ and $a=a_{\bm{R}}$. Then $\lim\limits_{n\to\infty}a_{n}(t)=a(t)$ for all $t\in[0,T)$. Moreover, if $\lambda\in L^{\infty}([0,T])$, then the convergence is uniform on $[0,T)$.
Proof.
See Appendix B.2. ∎
4 Optimal Equilibrium
Remark 3.4 shows that, if $\tilde{\mathbb{E}}[\bm{R}]=\infty$ and $H(\varepsilon)<\infty$ for all $\varepsilon>0$, then the set $\Pi^{e}_{d}$ of the equilibria contains infinitely many elements. This multiplicity of equilibria raises the natural question: which element of $\Pi^{e}_{d}$ should the agent choose? In this section, we propose that the agent choose the equilibrium $\pi$ that maximizes objective functional $J(\pi;0,w)$ at the initial time over all $\pi\in\Pi^{e}_{d}$. This leads to the following definition.
Definition 4.1.
An equilibrium $\hat{\pi}\in\Pi^{e}_{d}$ is called optimal if $J(\hat{\pi};0,w)\geq J(\pi;0,w)$ for all $\pi\in\Pi^{e}_{d}$ and for all $w>0$.
Recalling the relationship $J(\pi;t,w)=wJ_{0}(t,\pi)$ in Remark 2.3, finding an optimal equilibrium in $\Pi^{e}_{d}$ is equivalent to finding a $\hat{\pi}\in\Pi^{e}_{d}$ that maximizes $J_{0}(t,\pi)$.
Let $\pi\in\Pi^{e}_{d}$. We begin by computing $J_{0}(t,\pi)$. Using Theorem 3.2 (2), we have $\pi=\pi^{(T_{0})}$ for some $T_{0}\in\mathcal{T}$. Let $a^{(T_{0})}\triangleq\sigma^{\top}\pi^{(T_{0})}$ and $y^{(T_{0})}(t)\triangleq\int\limits_{t}^{T}(a_{s}^{(T_{0})})^{\top}\lambda(s)\mathrm{d}s$. Then
$-(\bm{R}v^{(T_{0})}(t)-2y^{(T_{0})}(t))$ $=-\bm{R}v^{(T_{0})}(t)+2\int_{t}^{T}(a_{s}^{(T_{0})})^{\top}\lambda(s)\mathrm{d}s$
$=-\bm{R}v^{(T_{0})}(t)+2\int_{t}^{T}h(v^{(T_{0})}(s))|\lambda(s)|^{2}\mathrm{d}s$
$=-\bm{R}v^{(T_{0})}(t)+2\int_{0}^{v^{(T_{0})}(t)}\frac{1}{h(y)}\mathrm{d}y,$
where the third equality arises from the change of variable $y=v^{(T_{0})}(s)$. Thus, we obtain
$J_{0}(t,\pi^{(T_{0})})=\exp\left\{\int_{0}^{v^{(T_{0})}(t)}\frac{1}{h(y)}\mathrm{d}y\right\}l\left(\frac{v^{(T_{0})}(t)}{2}\right).$
Let $\mathcal{L}(z):[0,\infty)\to\mathbb{R}$ be defined by $\mathcal{L}(z)=\int_{0}^{z}\frac{1}{h(y)}\mathrm{d}y+\log l\left(\frac{z}{2}\right).$ Then we have
$J_{0}(t,\pi^{(T_{0})})=\exp\left(\mathcal{L}(v^{(T_{0})}(t))\right).$ (4.1)
According to (4.1), in order to study the maximization of $J_{0}(t,\pi^{(T_{0})})$ over $T_{0}\in\mathcal{T}$, we need to analyze the monotonicity of $\mathcal{L}$ and the monotonicity of $v^{(T_{0})}$ with respect to $T_{0}$. These properties are characterized in Lemma 4.2 and Lemma 4.3, respectively.
Lemma 4.2.
$\mathcal{L}(z)$ is strictly increasing in $z$.
Proof.
By direct differentiation, we have
$\frac{d\mathcal{L}}{dz}=\frac{1}{h(z)}-\frac{1}{2}\frac{\tilde{\mathbb{E}}\left[\bm{R}e^{-\frac{1}{2}\bm{R}z}\right]}{\tilde{\mathbb{E}}\left[e^{-\frac{1}{2}\bm{R}z}\right]}=\frac{1}{2h(z)}>0\quad\text{for all }z>0.$
As such, $\mathcal{L}(z)$ is strictly increasing in $z$. ∎
Lemma 4.3.
Under the conditions of Assertion (2) of Theorem 3.2, let $T_{0,1},T_{0,2}\in\mathcal{T}$ with $T_{0,1}<T_{0,2}$. Then
$v^{(T_{0,1})}(t)\begin{cases}<v^{(T_{0,2})}(t),\quad t\in[0,T_{0,2}),\\ =v^{(T_{0,2})}(t),\quad t\in[T_{0,2},T].\end{cases}$ (4.2)
In particular, if $H(\infty)>\Lambda(0)$ and $\Lambda(t)>0$ for all $t\in[0,T)$, then $T\in\mathcal{T}$ and $v^{(T)}(t)>v^{(T_{0})}(t)$ for any $T_{0}\in\mathcal{T}\setminus\{T\}$ and any $t\in[0,T)$.
Proof.
Let $T_{0,1},T_{0,2}\in\mathcal{T}$ with $T_{0,1}<T_{0,2}$. For $t\in[T_{0,2},T]$, we have $v^{(T_{0,1})}(t)=v^{(T_{0,2})}(t)=0$ based on Definition of $\mathcal{T}$. For $t\in[T_{0,1},T_{0,2})$, we have $v^{(T_{0,2})}(t)>0=v^{(T_{0,1})}(t)$. For $t\in[0,T_{0,1})$, we have
$v^{(T_{0,i})}(t)=H^{-1}\left(\int_{t}^{T_{0,i}}|\lambda(s)|^{2}\mathrm{d}s\right)=H^{-1}(\Lambda(t)-\Lambda(T_{0,i})),\quad i=1,2.$
By the definition of $\mathcal{T}$, $\Lambda(T_{0,1})>\Lambda(T_{0,2})$. Because $H^{-1}$ is strictly increasing, it follows that $v^{(T_{0,1})}(t)<v^{(T_{0,2})}(t)$ for $t\in[0,T_{0,1})$. Thus, (4.2) is proved.
Assume $H(\infty)>\Lambda(0)$ and $\Lambda(t)>0$ for all $t\in[0,T)$. By the definition of $\mathcal{T}$, $T\in\mathcal{T}$. Plugging $T_{0,2}=T$ and $T_{0,1}=T_{0}$ into (4.2) yields $v^{(T)}(t)>v^{(T_{0})}(t)$ for any $T_{0}\in\mathcal{T}\setminus\{T\}$ and any $t\in[0,T)$. ∎
We are now ready to present the main result on the optimal equilibria.
Theorem 4.4.
Under the conditions of Assertion (2) of Theorem 3.2, we have the followings:
  1. (1)
    The optimal strategy exists if and only if $H(\infty)>\Lambda(0)$. If it exists, it is unique and given by $\pi^{(\varphi(0))}(\cdot)=(\sigma^{\top}(\cdot))^{-1}h(v^{(\varphi(0))}(\cdot))\lambda(\cdot)$.
  2. (2)
    $\pi^{(\varphi(0))}$ is also uniformly optimal:
    $J(\pi^{(\varphi(0))};t,w)\geq J(\pi;t,w)\text{ for all }t\in[0,T),w>0,\text{ and }\pi\in\Pi^{e}_{d}.$
  3. (3)
    $\pi^{(\varphi(0))}$ is uniformly strictly optimal, that is,
    $J(\pi^{(\varphi(0))};t,w)>J(\pi;t,w)\text{ for all }t\in[0,T),w>0,\text{ and }\pi\in\Pi^{e}_{d}\setminus\{\pi^{(\varphi(0))}\},$
    if and only if
[…]
    If it is the case, then $\varphi(0)=T$.
Proof.
Combining (4.1) with Lemma 4.2, we see that $J_{0}(0,\pi^{(T_{0})})$ attains its maximum precisely when $v^{(T_{0})}(0)$ is maximal. Moreover, by Lemma 4.3, $v^{(T_{0})}(0)$ is strictly increasing in $T_{0}\in\mathcal{T}$. Thus, finding the optimal equilibrium strategy is equivalent to identifying the maximal parameter $T_{0}\in\mathcal{T}$. By Lemma 3.3, $\mathcal{T}=\{\varphi(\eta)\mid\eta\in\mathcal{A}\}$. Because $\varphi$ is strictly decreasing on $\mathcal{A}$, the existence of a maximal element in $\mathcal{T}$ is equivalent to the existence of a minimal element in $\mathcal{A}$.
If $H(\infty)>\Lambda(0)$, then $\mathcal{A}=[0,\Lambda(0)]$ includes its minimum at 0, and hence a maximal strategy exists. Conversely, if $H(\infty)\leq\Lambda(0)$, then $\mathcal{A}=(\Lambda(0)-H(\infty),\Lambda(0)]$ is left-open and admits no minimum. Thus, $\mathcal{T}$ has a maximal element if and only if $H(\infty)>\Lambda(0)$. In conclusion, the optimal strategy exists if and only if $H(\infty)>\Lambda(0)$, and when it exists, it is given by $\pi^{(\varphi(0))}$.
We now turn to the uniformly optimal equilibrium. In view of the conclusion in assertion (1), we restrict attention to the case $H(\infty)>\Lambda(0)$. Once this condition holds, and following the same line of reasoning as in assertion (1), we immediately obtain that $\pi^{(\varphi(0))}$ is also uniformly optimal, because (4.1) remains valid for every $t\in[0,T]$.
Finally, we consider the uniformly strictly optimal equilibrium. If $\Lambda(t)=0$ for some $t\in[0,T)$, then $v^{(T_{0})}(t)\equiv 0$ for any $T_{0}\in\mathcal{T}$, which implies $J(t,\pi^{(T_{0})})\equiv\exp(\mathcal{L}(0))=1$ for all $\pi^{(T_{0})}\in\Pi_{d}^{e}$. Hence, no uniformly strictly optimal equilibrium exists in this case. Conversely, assume $\Lambda(t)>0$ for all $t\in[0,T)$ and $H(\infty)>\Lambda(0)$. Because $\Lambda(t)>0$ for all $t<T$, we have $\varphi(0)=T$. Let $T_{0}\in\mathcal{T}\setminus\{T\}$, by Lemma 4.3, for any $t\in[0,T)$, it holds that
$v^{(T)}(t)>v^{(T_{0})}(t).$
Because $\mathcal{L}(z)$ is strictly increasing in $z$, we obtain $J_{0}(t,\pi^{(T)})>J_{0}(t,\pi^{(T_{0})})$. Therefore, $\pi^{(\varphi(0))}=\pi^{(T)}$ is the unique uniformly strictly optimal equilibrium. ∎ Analogous to Proposition 3.8, we now investigate the impact of future market opportunities on the risk exposure vector in the case of $\tilde{\mathbb{E}}[\bm{R}]=\infty$. Given the multiplicity of equilibria in this setting, we focus specifically on the optimal equilibrium characterized in Theorem 4.4. The following proposition presents the corresponding comparative statics results. We omit the proof as it follows a similar line of reasoning to that of Proposition 3.8.
Proposition 4.5.
Let $\bm{R}$ be a non-negative random variable with $\tilde{\mathbb{E}}[\bm{R}]=\infty$. Fix a vector $\lambda_{0}\in\mathbb{R}^{d}$. Consider a sequence of market settings indexed by $n\in\mathbb{N}$, defined by the pairs $\{(\lambda_{n}(\cdot),T_{n})\}_{n\geq 1}$, where $T_{n}>0$ and $\lambda_{n}:[0,T_{n}]\to\mathbb{R}^{d}$ satisfy $\lambda_{n}(0)=\lambda_{0}$. Let $\Lambda_{n}\triangleq\int_{0}^{T_{n}}|\lambda_{n}(s)|^{2}ds$ and assume $H(\infty)>\Lambda_{n}$ for all $n$. Let $a_{n}(0)$ denote the risk exposure vector at time 0 corresponding to the optimal equilibrium of the $n$-th market setting.
  1. (1)
    If the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly increasing and $\Lambda_{n}\nearrow H(\infty)$, then
    $\lim_{n\to\infty}a_{n}(0)=\frac{\lambda_{0}}{r_{0}},$
    where $r_{0}\triangleq\operatorname{essinf}\bm{R}$. Moreover, $|a_{n}(0)|$ converges monotonically increasing, i.e.,
    $|a_{n}(0)|\nearrow\frac{|\lambda_{0}|}{r_{0}}.$
  2. (2)
    If the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly decreasing and $\Lambda_{n}\searrow 0$, then
    $\lim_{n\to\infty}a_{n}(0)=\mathbf{0}.$
    Moreover, $|a_{n}(0)|$ converges monotonically decreasing, i.e.,
    $|a_{n}(0)|\searrow 0.$
Here, we adopt the convention that $\frac{1}{0}=\infty$.
5 Comparative Statics
In the previous sections, we fully characterized the equilibrium strategies associated with a given distribution of risk aversion $\bm{R}$ and discussed how to select the optimal equilibrium. We now turn to comparative statics and ask a natural economic question: how does a shift in the distribution of risk aversion affect the resulting equilibrium investment behavior? In this section, we only consider the case where the equilibrium exists.
To formalize such comparisons, we begin by recalling the notion of first-order stochastic dominance, which provides a standard criterion for ranking risk aversion distributions.
Definition 5.1.
Let $\bm{R_{1}}$ and $\bm{R_{2}}$ be two random variables. We say that $\bm{R_{1}}$ dominates $\bm{R_{2}}$ in the sense of first-order stochastic dominance, denoted as $\bm{R_{1}}\succeq_{1}\bm{R_{2}}$, if
$\tilde{\mathbb{P}}(\bm{R_{1}}\geq x)\geq\tilde{\mathbb{P}}(\bm{R_{2}}\geq x)\quad\text{for all }x\in\mathbb{R}.$
Intuitively, one might expect that if Investor 1 is “more risk-averse” than Investor 2 in the sense of first-order stochastic dominance, then Investor 1 should take a smaller position in the risky asset at all times. Surprisingly, this intuition turns out to be false, as the following numerical example shows.
Example 5.2.
Assume that there is one stock whose market price of risk is constant, $\lambda(\cdot)\equiv\lambda=0.4$. The time horizon $T=20$. The risk-aversion distributions of the two investors are:
Investor 1: $\tilde{\mathbb{P}}(\bm{R}_{1}=1)=0.9$, $\tilde{\mathbb{P}}(\bm{R}_{1}=3)=0.1$.
Investor 2: $\tilde{\mathbb{P}}(\bm{R}_{2}=1)=0.9$, $\tilde{\mathbb{P}}(\bm{R}_{2}=2)=0.1$.
Obviously, $\bm{R}_{1}\succeq_{1}\bm{R}_{2}$. Figure 1 displays the comparative evolution of $|a_{i}|$, i=1,2i=1,2.555Because $\bar{\pi}=(\sigma^{\top})^{-1}a$, in the one-dimensional case comparing $|a_{i}(t)|$ is equivalent to comparing $|\pi|_{i}$. More generally, $|a|$ represents the volatility of the investor’s wealth process and thus captures the level of risk borne by the investor. Therefore, in what follows, we focus conceptually on comparing the risk exposure magnitude. The orange-shaded Reversal Region highlights that the more risk-averse Investor 1 adopts a strictly more aggressive position than Investor 2 when the time to maturity is sufficiently long. This counterexample shows that under first-order stochastic dominance, the magnitudes of equilibrium investment strategies need not be monotone.
Figure 1: Comparative evolution of $a_{i}(t)$ under first-order stochastic dominance. The red solid line corresponds to Investor 1 (more risk-averse), with distribution $\tilde{\mathbb{P}}(\bm{R}_{1}=1)=0.9$, $\tilde{\mathbb{P}}(\bm{R}_{1}=3)=0.1$; the blue dashed line corresponds to Investor 2 (less risk-averse), with distribution $\tilde{\mathbb{P}}(\bm{R}_{2}=1)=0.9$, $\tilde{\mathbb{P}}(\bm{R}_{2}=2)=0.1$. The orange shaded region indicates where $|a_{1}(t)|>|a_{2}(t)|$. Parameters: $\lambda=0.4$, horizon $T=20$.
The appearance of this counterexample (Figure 1) may suggest that the search for a general comparative result under first-order stochastic dominance is fundamentally obstructed. In fact, the observed reversal is not an artifact of a particular parameter choice. We show below that such a reversal in investment magnitudes is unavoidable when the investment horizon $T$ is sufficiently large, provided that the risk-aversion distributions satisfy the conditions stated in the following proposition.
In the following proposition, for convenience, we extend the market to an infinite horizon and assume that the market price of risk $\lambda$ is defined on $0,\infty)$. For every fixed $T>0$, we require $\Lambda(0,T)\triangleq\int_{0}^{T}|\lambda(s)|^{2}\mathrm{d}s<\infty$. In addition, we assume that $\lim_{T\to\infty}\Lambda(0,T)=\infty$ and $\lambda(0)\neq 0$. As discussed in the preceding sections, for each fixed $T$, we can define the equilibrium strategy $\{\pi_{i}(t,T):t\in[0,T)\}$ associated with the distribution $\bm{R}_{i}$ (as used in the proposition below). The proposition shows that, as long as the distributions $\bm{R}_{i}$ satisfy certain conditions, the magnitudes $|\pi_{i}(0,T)|$ must exhibit a strict and deterministic ordering when the horizon $T$ is sufficiently large.
Proposition 5.3.
Let $\bm{R}_{1}$ and $\bm{R}_{2}$ be two non-negative random variables with finite expectations i.e., $\tilde{\mathbb{E}}[\bm{R}_{i}]<\infty$, i=1, 2i=1,\ 2. Assume that the following conditions are satisfied:
  1. (1)
    Let $r_{0}=\operatorname{essinf}\bm{R}_{1}=\operatorname{essinf}\bm{R}_{2}>0$, and assume that there is a positive probability mass at $r_{0}$, i.e., $p_{i}=\tilde{\mathbb{P}}(\bm{R}_{i}=r_{0})\in(0,1)$ for $i=1,2$.
  2. (2)
    Define the infimum of the first support point of $\bm{R}_{i}$ strictly greater than $r_{0}$ as $\tilde{r}_{i}=\inf\{r\in supp(\bm{R}_{i}):r>r_{0}\}$, and let $\delta_{i}=\tilde{r}_{i}-r_{0}$ with $\delta_{1}>\delta_{2}>0$.
Under these conditions, for a sufficiently large $T$, we have
$|\pi_{1}(0,T)|>|\pi_{2}(0,T)|.$
Proof.
See Appendix B.3. ∎
As investigated by Desmettre2023, we now focus on the two-point distribution, which represents the closest deviation from the standard expected utility model. This setting allows us to derive a single crossing property, rigorously characterizing the transition from short-term dominance to the long-term reversal observed in Proposition 5.3.
Proposition 5.4.
Under the conditions of Proposition 5.3, suppose further that $\lambda(t)\neq 0$ for all $t\in[0,T)$, and that $\bm{R}_{1}$ and $\bm{R}_{2}$ follow two-point distributions with $\tilde{\mathbb{E}}[\bm{R}_{1}]>\tilde{\mathbb{E}}[\bm{R}_{2}]$. Moreover, assume that $T$ is sufficiently large so that $|a_{1}(0)|>|a_{2}(0)|$. Then $|a_{1}(\cdot)|$ and $|a_{2}(\cdot)|$ exhibit a single-crossing property: there exists a unique $t^{*}\in[0,T]$ such that $|a_{1}(t^{*})|=|a_{2}(t^{*})|$.
Proof.
See Appendix B.4. ∎
Having established the single crossing property, we investigate how the probability distribution drives the timing of this reversal. Intuitively, a larger probability assignment accelerates the strategy’s convergence to the log-utility benchmark, leading to a faster intersection as time recedes from the horizon.
Proposition 5.5.
Under the conditions of Proposition 5.3 and Proposition 5.4, assume additionally that $p_{1}=p_{2}=p$. Let $t^{*}$ be the intersection time where $|a_{1}(t^{*})|=|a_{2}(t^{*})|$. Then, $t^{*}$ is strictly increasing with respect to $p$.
Proof.
Recalling that $Q_{i}(x)=\dfrac{\delta_{i}(1-p)e^{-\delta_{i}\frac{x}{2}}}{p+(1-p)e^{-\delta_{i}\frac{x}{2}}}$, we know that $Q_{i}$ is strictly decreasing and satisfies $U_{i}=U_{i}(p)\triangleq Q_{i}(0)=\delta_{i}(1-p)$. Moreover, $\Lambda(t)=\int_{0}^{v_{i}(t)}(r_{0}+Q_{i}(x))^{2}\,\mathrm{d}x=2\int_{Q_{i}(v_{i}(t))}^{U_{i}(p)}\dfrac{(r_{0}+q)^{2}}{q(\delta_{i}-q)}\,\mathrm{d}q\triangleq 2\int_{Q_{i}(v_{i}(t))}^{U_{i}(p)}G(q,\delta)\,\mathrm{d}q$ where the second equality arises from $q=Q_{i}(x)$, $\mathrm{d}x=\dfrac{2}{q(q-\delta_{i})}\,\mathrm{d}q$ and $G(q,\delta)\triangleq\dfrac{(r_{0}+q)^{2}}{q(\delta-q)}$. At time $t^{*}$, let $Q=Q(p)\triangleq Q_{1}(v_{1}(t^{*}))=Q_{2}(v_{2}(t^{*}))$ and $\Lambda^{*}=\Lambda^{*}(p)\triangleq\Lambda(t^{*})$. The condition $|a_{1}(t^{*})|=|a_{2}(t^{*})|$ implies
$\Lambda^{*}=2\int_{Q}^{U_{1}(p)}G(q,\delta_{1})\,dq=2\int_{Q}^{U_{2}(p)}G(q,\delta_{2})\,dq.$ (5.1)
Differentiating (5.1) with respect to $p$ and we have
$\dfrac{1}{2}\dfrac{\mathrm{d}\Lambda^{*}}{\mathrm{d}p}$ $=G(U_{1},\delta_{1})\dfrac{\mathrm{d}U_{1}}{\mathrm{d}p}-G(Q,\delta_{1})\dfrac{\mathrm{d}Q}{\mathrm{d}p}$
$=-\delta_{1}G(U_{1},\delta_{1})-G(Q,\delta_{1})\dfrac{\mathrm{d}Q}{\mathrm{d}p}.$
Similarly, for the second distribution, we have
$\dfrac{1}{2}\dfrac{\mathrm{d}\Lambda^{*}}{\mathrm{d}p}=-\delta_{2}G(U_{2},\delta_{2})-G(Q,\delta_{2})\dfrac{\mathrm{d}Q}{\mathrm{d}p}.$
Eliminating $\dfrac{\mathrm{d}Q}{\mathrm{d}p}$ from the system, we obtain
$\dfrac{1}{2}\dfrac{\mathrm{d}\Lambda^{*}}{\mathrm{d}p}=\dfrac{\delta_{2}G(U_{2},\delta_{2})G(Q,\delta_{1})-\delta_{1}G(U_{1},\delta_{1})G(Q,\delta_{2})}{G(Q,\delta_{2})-G(Q,\delta_{1})}.$ (5.2)
Substituting $U_{i}=\delta_{i}(1-p)$ into $G$, we have
$\delta_{i}G(U_{i},\delta_{i})=\dfrac{(r_{0}+\delta_{i}(1-p))^{2}}{p(1-p)\delta_{i}}.$
Recalling the form of $Q_{i}$ and the fact that $Q_{i}$ is strictly decreasing, we have $\delta_{1}>\delta_{2}>U_{2}>Q$, which implies $\delta_{1}-Q>\delta_{2}-Q>0$ and $G(Q,\delta_{1})<G(Q,\delta_{2})$. Thus, the denominator of (5.2) is positive. For the numerator, we have
$\dfrac{\delta_{1}G(U_{1},\delta_{1})}{G(Q,\delta_{1})}$ $=\dfrac{(r_{0}+\delta_{1}(1-p))^{2}}{p(1-p)\delta_{1}}\cdot\dfrac{Q(\delta_{1}-Q)}{(r_{0}+Q)^{2}}=\dfrac{Q}{p(1-p)(r_{0}+Q)^{2}}(r_{0}+\delta_{1}(1-p))^{2}\left(1-\dfrac{Q}{\delta_{1}}\right),$
$\dfrac{\delta_{2}G(U_{2},\delta_{2})}{G(Q,\delta_{2})}$ $=\dfrac{Q}{p(1-p)(r_{0}+Q)^{2}}(r_{0}+\delta_{2}(1-p))^{2}\left(1-\dfrac{Q}{\delta_{2}}\right).$
Because $\delta_{1}>\delta_{2}$, it holds that $(r_{0}+\delta_{1}(1-p))^{2}>(r_{0}+\delta_{2}(1-p))^{2}$ and $1-\dfrac{Q}{\delta_{1}}>1-\dfrac{Q}{\delta_{2}}$. Therefore, $\dfrac{\delta_{1}G(U_{1},\delta_{1})}{G(Q,\delta_{1})}>\dfrac{\delta_{2}G(U_{2},\delta_{2})}{G(Q,\delta_{2})}$, which implies $\delta_{2}G(U_{2},\delta_{2})G(Q,\delta_{1})-\delta_{1}G(U_{1},\delta_{1})G(Q,\delta_{2})<0$. Consequently, $\dfrac{\mathrm{d}\Lambda^{*}}{\mathrm{d}p}<0$. Because $\Lambda(t)$ is a strictly decreasing function of $t$, a decrease in the required energy $\Lambda^{*}$ corresponds to an increase in the intersection time $t^{*}$. ∎
Building on this analysis, we now relax the symmetry assumption to examine the individual effects. The following corollary presents the comparative statics of the intersection time with respect to the probability parameter of each investor.
Corollary 5.6.
Under the conditions of Proposition 5.3 and Proposition 5.4, we have the following assertions.
  1. (1)
    If $p_{2}$ is fixed, $t^{*}$ is strictly increasing with respect $p_{1}$.
  2. (2)
    If $p_{1}$ is fixed, $t^{*}$ is strictly decreasing with respect $p_{2}$
Proof.
We retain the notations from the proof of Proposition 5.4.
Differentiating (5.1) with respect to $p_{1}$ yields
$\dfrac{1}{2}\dfrac{\partial\Lambda^{*}}{\partial p_{1}}$ $=-\delta_{1}G(U_{1},\delta_{1})-G(Q,\delta_{1})\dfrac{\partial Q}{\partial p_{1}},$
$\dfrac{1}{2}\dfrac{\partial\Lambda^{*}}{\partial p_{1}}$ $=-G(Q,\delta_{2})\dfrac{\partial Q}{\partial p_{1}}.$
Eliminating $\dfrac{\partial Q}{\partial p_{1}}$, we obtain
$\dfrac{1}{2}\dfrac{\partial\Lambda^{*}}{\partial p_{1}}=\dfrac{-\delta_{1}G(U_{1},\delta_{1})G(Q,\delta_{2})}{G(Q,\delta_{2})-G(Q,\delta_{1})}.$
Because the denominator is positive and the numerator is negative, we have $\dfrac{\partial\Lambda^{*}}{\partial p_{1}}<0$. Given that $\Lambda(t)$ is strictly decreasing, $t^{*}$ must increase.
Differentiating (5.1) with respect to $p_{2}$ yields
$\dfrac{1}{2}\dfrac{\partial\Lambda^{*}}{\partial p_{2}}$ $=-G(Q,\delta_{1})\dfrac{\partial Q}{\partial p_{2}},$
$\dfrac{1}{2}\dfrac{\partial\Lambda^{*}}{\partial p_{2}}$ $=-\delta_{2}G(U_{2},\delta_{2})-G(Q,\delta_{2})\dfrac{\partial Q}{\partial p_{2}}.$
Eliminating $\dfrac{\partial Q}{\partial p_{2}}$, we have
$\dfrac{1}{2}\dfrac{\partial\Lambda^{*}}{\partial p_{2}}=\dfrac{-\delta_{2}G(U_{2},\delta_{2})G(Q,\delta_{1})}{G(Q,\delta_{1})-G(Q,\delta_{2})}.$
Because the denominator is negative while the numerator remains negative, we have $\dfrac{\partial\Lambda^{*}}{\partial p_{2}}>0$, which implies that $t^{*}$ must decrease. ∎
It is important to note that the conditions in Proposition 5.3 and Proposition 5.4 are compatible with first-order stochastic dominance. One can construct risk aversion distributions such that $\bm{R}_{1}\succeq_{1}\bm{R}_{2}$ holds while the conditions in Proposition 5.3 and 5.4 are simultaneously satisfied. For instance, the parameters in Figure 1 satisfy these conditions with $r_{0}=1$, $\delta_{1}=4$, $\delta_{2}=1$, $\tilde{\mathbb{E}}[\bm{R}_{1}]=3>1.5=\tilde{\mathbb{E}}[\bm{R}_{2}]$ and $\lambda\equiv 0.4$. Consequently, first-order stochastic dominance fails to rule out the reversal of investment magnitudes characterized by $|\pi_{1}(0,T)|>|\pi_{2}(0,T)|$ for a sufficiently large $T$ and we can also observe the single crossing property.
However, we demonstrate that the desired monotonicity can be restored under a stronger stochastic order. Specifically, we introduce the concept of reverse hazard rate order and show that if $\bm{R_{1}}$ dominates $\bm{R_{2}}$ in this order, the intuition holds, i.e., $|\pi_{1}(t)|\leq|\pi_{2}(t)|$.
Definition 5.7.
Let $\bm{R_{1}}$ and $\bm{R_{2}}$ be two random variables with cumulative distribution functions $F_{\bm{R_{1}}}$ and $F_{\bm{R_{2}}}$, respectively. We say that $\bm{R_{1}}$ dominates $\bm{R_{2}}$ in the reverse hazard rate order, denoted as $\bm{R_{1}}\succeq_{rh}\bm{R_{2}}$, if the ratio
$\frac{F_{\bm{R_{2}}}(\gamma)}{F_{\bm{R_{1}}}(\gamma)}$
is decreasing in $\gamma$ for all $\gamma$ in the union of the supports of $\bm{R_{1}}$ and $\bm{R_{2}}$ where $F_{\bm{R_{1}}}(\gamma)>0$.
The following proposition compares the equilibrium strategies of two investors ranked by the reverse hazard rate order.
Proposition 5.8.
Let $\bm{R_{1}}$ and $\bm{R_{2}}$} be two non-negative random variables with $\tilde{\mathbb{E}}\left[\bm{R_{i}}\right]<+\infty$ ($i=1,2$), such that $\bm{R_{1}}\succeq_{rh}\bm{R_{2}}$ and their distributions are distinct. Let $\pi_{i}(t)=(\sigma^{\top})^{-1}a_{i}(t)$ be the corresponding deterministic equilibrium strategies. For any $t\in[0,T)$, the magnitude of $a_{i}(t)$ satisfies
$|a_{1}(t)|\leq|a_{2}(t)|.$
Moreover, if $\lambda(t)\neq 0$, the inequality is strict. As a result, the investor with higher risk aversion adopts a less aggressive equilibrium strategy. i.e., $|\pi_{1}(t)|\leq|\pi_{2}(t)|$.
Proof.
Because $h_{1}(x)<h_{2}(x)$ for all positive $s$, it follows that $H_{1}(y)>H_{2}(y)$ for all $y>0$. This implies that the inverse functions satisfy $H_{2}^{-1}(z)<H_{2}^{-1}(z)$ for all $z>0$. From Theorem 3.2 (1), we have $v_{i}(t)=H_{i}^{-1}(\Lambda(t))$, where $\Lambda(t)=\int_{t}^{T}|\lambda(s)|^{2}ds$ and $v_{i}(t)=\int_{t}^{T}|a_{i}(s)|^{2}\mathrm{d}s$. Consequently, monotonicity of the inverse functions implies $v_{1}(t)\leq v_{2}(t)$. By (2.5), we have $|a_{i}(t)|=h(v_{i}(t),\bm{R_{i}})|\lambda(t)|$. From Proposition A.6, we know that $h_{1}(v_{1}(t))<h_{2}(v_{1}(t))$. Because $v_{1}(t)\leq v_{2}(t)$, it holds that $h_{2}(v_{1}(t))\leq h_{2}(v_{2}(t))$. Combining these relations yields $h_{1}(v_{1}(t))<h_{2}(v_{2}(t))$. Multiplying by $|\lambda(t)|$, we obtain $|a_{1}(t)|\leq|a_{2}(t)|$. If $\lambda(t)\neq 0$, the strict inequality $|a_{1}(t)|<|a_{2}(t)|$ is preserved. ∎
For the case where the expected risk aversion is infinite, assume that the optimal equilibria exist for both $\bm{R}_{1}$ and $\bm{R}_{2}$. Let $a_{i}(t)$ denote the deterministic coefficient corresponding to the optimal equilibrium strategy for $\bm{R}_{i}$ ($i=1,2$). Note that $\varphi(0)=\min\{t\in[0,T]\mid\Lambda(t)=0\}$ is determined solely by the market parameters and is therefore identical for both investors. Consequently, we can establish the comparative result by examining the magnitudes of $a_{i}(t)$ on the intervals $0,\varphi(0))$ and $\varphi(0),T]$ separately. Specifically, on the interval $0,\varphi(0))$, the analysis parallels the finite expectation case derived in Proposition 5.8, and thus the monotonicity of the equilibrium strategies is preserved. On the interval $\varphi(0),T]$, the definition of $\varphi(0)$ implies that $\lambda(t)=0$ almost everywhere. Consequently, the equilibrium strategies for both investors vanish, trivially satisfying the comparison inequality. Based on this decomposition, we have the following corollary.
Corollary 5.9.
Let $\bm{R}_{1}$ and $\bm{R}_{2}$ be two non-negative random variables with $\tilde{\mathbb{E}}[\bm{R}_{i}]=\infty$ ($i=1,2$), such that $\bm{R}_{1}\succeq_{rh}\bm{R}_{2}$ and their distributions are distinct. Assume that the optimal equilibria exist for both investors (i.e., $H_{i}(\infty)>\Lambda(0)$, $i=1,2$). Let $\pi_{i}(t)$ be the corresponding deterministic optimal equilibrium strategies. Then, for any $t\in[0,T)$, the magnitude of $\pi_{i}(t)$ satisfies
$|\pi_{1}(t)|\leq|\pi_{2}(t)|.$
Moreover, if $\lambda(t)\neq 0$, the inequality is strict.
While the previous results focus on comparing distinct distributions ranked by stochastic orders, it is also of theoretical interest to understand the effect of preference aggregation. We now investigate the equilibrium strategy when the risk aversion parameter is formed by a convex combination (i.e., a weighted average) of independent and identically distributed components.
Proposition 5.10.
Let $\bm{R}_{1},\cdots,\bm{R}_{n}$ be i.i.d. non-negative random variables with $\tilde{\mathbb{E}}\left[\bm{R_{i}}\right]<+\infty$ ($i=1,2,\cdots,n$) and $\bm{R}=\sum_{i=1}^{n}w_{i}\bm{R}_{i}$ be a convex combination with $w_{i}\in(0,1)$ and $\sum_{i=1}^{n}w_{i}=1$. Let $a(t)$ and $a_{i}(t)$ denote the risk exposure vector associated to the risk aversion $\bm{R}$ and $\bm{R}_{i}$, respectively. Then we have
$|a(t)|\leq|a_{1}(t)|\quad\text{for all }t\in[0,T).$
If $\bm{R}_{i}$ is non-degenerate and $\lambda(t)\neq 0$, the inequality is strict.
Proof.
Let $h(x)$ and $v(t)$ (resp. $h_{i}(x)$ and $v_{i}(t)$) be the functions associated to $\bm{R}$ (resp. $\bm{R_{i}}$ ). From the definition of $\bm{R}$, we have
$\dfrac{1}{h(x)}=\sum_{j=1}^{n}w_{j}\dfrac{\tilde{\mathbb{E}}[\bm{R}_{j}e^{-\frac{1}{2}w_{j}\bm{R}_{j}x}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}w_{j}\bm{R}_{j}x}]}$ (5.3)
Because $\bm{R}_{j}$ are identically distributed to $\bm{R}_{1}$, each fraction in the sum is exactly $\dfrac{1}{h_{1}(w_{j}x)}$. Thus (5.3) is actually
$\dfrac{1}{h(x)}=\sum_{j=1}^{n}w_{j}\dfrac{1}{h_{1}(w_{j}x)}$
According to Proposition A.6, we know that $\dfrac{1}{h_{1}(x)}$ is decreasing in $x$. Because $w_{j}\in(0,1)$ and $x>0$, we have $w_{j}x<x$ and $\dfrac{1}{h_{1}(w_{j}x)}\geq\dfrac{1}{h_{1}(x)}$ and
$\dfrac{1}{h(x)}=\sum_{j=1}^{n}w_{j}\dfrac{1}{h_{1}(w_{j}x)}\geq\sum_{j=1}^{n}w_{j}\dfrac{1}{h_{1}(x)}=\dfrac{1}{h_{1}(x)}$
This implies $h(x)\leq h_{1}(x)$. Analogous to the proof in Proposition 5.8, we have $v(t)\leq v_{1}(t)$. Finally, because $h(x)$ is increasing in $x$, we have
$|a(t)|=h(v(t))|\lambda(t)|\leq h(v_{1}(t))|\lambda(t)|\leq h_{1}(v_{1}(t))|\lambda(t)|=|a_{1}(t)|$
If $\bm{R}_{i}$ is non-degenerate and $\lambda(t)\neq 0$, $h$ is strictly increasing, and all inequalities in the above proof become strict. Thus we have $|a(t)|<|a_{1}(t)|$
It is important to note that the inequality established above relies critically on the i.i.d. assumption. The following example demonstrates that if we retain independence but relax the identical distribution assumption, the equilibrium strategy for the convex combination need not be bounded by the strategies of the individual components.
Example 5.11.
Assume that there is one stock whose market price of risk is constant, $\lambda(\cdot)\equiv\lambda=0.5$. The time horizon $T=50$. The risk-aversion distributions of the two investors are:
Investor 1: $\tilde{\mathbb{P}}(\bm{R}_{1}=0.1)=0.2$, $\tilde{\mathbb{P}}(\bm{R}_{1}=8)=0.8$.
Investor 2: $\tilde{\mathbb{P}}(\bm{R}_{2}=1.5)=1$.
Assume that $\bm{R}_{1}$ and $\bm{R}_{2}$ are independent. We consider a composite investor with risk aversion $\bm{R}=0.5\bm{R}_{1}+0.5\bm{R}_{2}$.
Figure 2 displays the comparative evolution of $|a(t)|$, $|a_{1}(t)|$ and $|a_{2}(t)|$. The red-shaded Reversal Region highlights that the investor with the mixed risk aversion $\bm{R}$ adopts a strictly more aggressive position than both Investor 1 and Investor 2 when the time to maturity is sufficiently long. This counterexample shows that without the i.i.d. assumption, the magnitudes of equilibrium investment strategies for a convex combination of risk aversions need not be bounded by the strategies of the individual components.
Figure 2: Comparative evolution of $a(t)$ for a convex combination of risk aversions. The black dashed line corresponds to Investor 1, with distribution $\tilde{\mathbb{P}}(\bm{R}_{1}=0.1)=0.2,\tilde{\mathbb{P}}(\bm{R}_{1}=8.0)=0.8$; the blue dash-dotted line corresponds to Investor 2, with degenerate distribution $\tilde{\mathbb{P}}(\bm{R}_{2}=1.5)=1.0$. The red solid line corresponds to the mixed investor with risk aversion $\bm{R}=0.5\bm{R}_{1}+0.5\bm{R}_{2}$. The red-shaded Reversal Region indicates where $|a(t)|>\max\{|a_{1}(t)|,|a_{2}(t)|\}$. Parameters: $\lambda=0.5$, horizon $T=50$.
We now specifically investigate the aggregation of i.i.d. risk aversions. Distinct from the general result above, the following proposition utilizes the i.i.d. property to prove that the equilibrium strategy converges monotonically to its limit.
Proposition 5.12.
Suppose that $\{\bm{R}_{i}\}_{i=1}^{\infty}$ is a sequence of independent and identically distributed non-negative random variables with $\tilde{\mathbb{E}}\left[\bm{R_{i}}\right]=\mu\in(0,+\infty)$ ($i=1,2,\cdots$,). Let $S_{n}=\frac{1}{n}\sum_{i=1}^{n}\bm{R}_{i}$ denote the sample mean risk aversion and $a_{n}(t)$ denote the risk exposure vector associated to the risk aversion $S_{n}$. Then, as $n\to\infty$, we have
$|a_{n}(t)|\searrow\frac{|\lambda(t)|}{\mu}\quad\text{for all }t\in[0,T).$ (5.4)
Proof.
Let $h_{n}(x)$ and $v_{n}(t)$ be the functions associated to $S_{n}$. First, substituting $S_{n}=\frac{1}{n}\sum_{i=1}^{n}\bm{R}_{i}$ into the definition of $h_{n}(x)$ and utilizing the i.i.d. property of $\{\bm{R}_{i}\}$, we obtain
$h_{n}(x)=\frac{\tilde{\mathbb{E}}\left[\prod_{i=1}^{n}e^{-\frac{x}{2n}\bm{R}_{i}}\right]}{\displaystyle\frac{1}{n}\sum_{j=1}^{n}\tilde{\mathbb{E}}\left[\bm{R}_{j}e^{-\frac{x}{2n}\bm{R}_{j}}\prod_{k\neq j}e^{-\frac{x}{2n}\bm{R}_{k}}\right]}$ $=\frac{(\tilde{\mathbb{E}}\left[e^{-\frac{x}{2n}\bm{R}_{1}}\right])^{n}}{\tilde{\mathbb{E}}\left[\bm{R}_{1}e^{-\frac{x}{2n}\bm{R}_{1}}\right](\tilde{\mathbb{E}}\left[e^{-\frac{x}{2n}\bm{R}_{1}}\right])^{n-1}}$
$=\frac{\tilde{\mathbb{E}}\left[e^{-\frac{1}{2}\bm{R}_{1}(\frac{x}{n})}\right]}{\tilde{\mathbb{E}}\left[\bm{R}_{1}e^{-\frac{1}{2}\bm{R}_{1}(\frac{x}{n})}\right]}=h_{1}\left(\frac{x}{n}\right).$
According to Proposition A.6, $h_{1}(x)$ is increasing. For any fixed $x\geq 0$, the sequence $\frac{x}{n}$ is decreasing in $n$, implying that $h_{n}(x)$ is decreasing in $n$. Furthermore, because $H_{n}(y)=\int_{0}^{y}h_{n}^{-2}(z)dz$ and $v_{n}(t)=H_{n}^{-1}(\Lambda(t))$, the monotonicity of $h_{n}$ implies that $v_{n}(t)$ is strictly decreasing in $n$. Then we have
$h_{n+1}(v_{n+1}(t))\leq h_{n}(v_{n+1}(t))\leq h_{n}(v_{n}(t)),$
equivalently, we have
$|a_{n+1}(t)|\leq|a_{n}(t)|,\quad\forall t\in[0,T).$
Now we determine the limit of $|a_{n}(t)|$. By the Strong Law of Large Numbers, $S_{n}\to\mu$ almost surely as $n\to\infty$. This implies that $S_{n}$ converges in distribution to the degenerate random variable $R\equiv\mu$. Moreover, because $\tilde{\mathbb{E}}[S_{n}]=\mu$ for all $n$, the assumptions of Proposition 3.8 are satisfied. Thus we conclude that
$\lim_{n\to\infty}|a_{n}(t)|=h(v(t))|\lambda(t)|=\frac{|\lambda(t)|}{\mu}.$
Combining this with the monotonicity established above, we obtain (5.4). ∎
Acknowledgement
Zongxia Liang acknowledges the financial support from the National Natural Science Foundation of China (grant 12271290). Sheng Wang acknowledges Professor Ka Chun Cheung and the financial supports as a postdoctoral fellow from the Department of Statistics and Actuarial Science, School of Computing and Data Science, the University of Hong Kong. Jianming Xia acknowledges the financial support from the National Natural Science Foundation of China (grants 12431017 and 12471447).
Appendix A Properties of the function $h$
In this appendix, we present some properties of the function $h$ defined in (2.6).
Lemma A.1 (Monotonicity).
The function $h$ is increasing in $0,\infty)$. The monotonicity is strict unless $\bm{R}$ is a constant.
Proof.
By direct differentiation, we have
$h^{\prime}(x)=\frac{1}{2}\frac{\tilde{\mathbb{E}}\left[e^{-\frac{1}{2}\bm{R}x}\right]\tilde{\mathbb{E}}\left[\bm{R}^{2}e^{-\frac{1}{2}\bm{R}x}\right]-\left(\tilde{\mathbb{E}}\left[\bm{R}e^{-\frac{1}{2}\bm{R}x}\right]\right)^{2}}{\left(\tilde{\mathbb{E}}\left[\bm{R}e^{-\frac{1}{2}\bm{R}x}\right]\right)^{2}}\quad\text{for all }x>0.$
By Hölder’s inequality,
$\left(\tilde{\mathbb{E}}\left[\bm{R}e^{-\frac{1}{2}\bm{R}x}\right]\right)^{2}=\left(\tilde{\mathbb{E}}\left[\bm{R}e^{-\frac{1}{2}\bm{R}x}\right]\cdot\tilde{\mathbb{E}}\left[\bm{R}e^{-\frac{1}{2}\bm{R}x}\right]\right)\leq\tilde{\mathbb{E}}\left[e^{-\frac{1}{2}\bm{R}x}\right]\tilde{\mathbb{E}}\left[\bm{R}^{2}e^{-\frac{1}{2}\bm{R}x}\right].$ (A.1)
Hence, $h^{\prime}\geq 0$ and $h$ is increasing. The inequality in (A.1) becomes an equality if and only if $\tilde{\mathbb{E}}\left[e^{-\frac{1}{2}\bm{R}x}\right]\bm{R}^{2}e^{-\frac{1}{2}\bm{R}x}=e^{-\frac{1}{2}\bm{R}x}\tilde{\mathbb{E}}\left[\bm{R}^{2}e^{-\frac{1}{2}\bm{R}x}\right]$, which is equivalent to $\bm{R}$ following a single-point distribution. Therefore, $h$ is strictly increasing unless $\bm{R}$ is a constant. ∎
Lemma A.2 (Asymptotic Analysis).
Let $\bm{R}$ be a non-negative random variable, and let $r_{0}\triangleq\operatorname{essinf}\bm{R}$. Then,
$\lim_{x\to\infty}h(x)=\frac{1}{r_{0}},$
where we adopt the convention that $\frac{1}{0}=\infty$.
Proof.
Define $\bm{X}\triangleq\bm{R}-r_{0}$, then
$\frac{1}{h(x)}=\frac{\tilde{\mathbb{E}}[\bm{R}e^{-\frac{1}{2}\bm{R}x}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}x}]}=\frac{\tilde{\mathbb{E}}[(r_{0}+\bm{X})e^{-\frac{1}{2}(r_{0}+\bm{X})x}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}(r_{0}+\bm{X})x}]}=r_{0}+Q(x),$ (A.2)
where $Q(x)\triangleq\frac{\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}\bm{X}x}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}x}]}$. For any arbitrary $\varepsilon>0$, because $\operatorname{essinf}\bm{X}=0$, there exists a probability mass $p_{\varepsilon}\triangleq\tilde{\mathbb{P}}(0\leq\bm{X}\leq\varepsilon)>0$. Therefore,
$Q(x)\leq\frac{\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}\bm{X}x}\mathbb{I}_{\{\bm{X}\leq\varepsilon\}}+\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}\bm{X}x}\mathbb{I}_{\{\bm{X}>\varepsilon\}}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}x}\mathbb{I}_{\{\bm{X}\leq\varepsilon\}}]}\leq\varepsilon+\frac{\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}\bm{X}x}\mathbb{I}_{\{\bm{X}>\varepsilon\}}]}{p_{\varepsilon}e^{-\frac{1}{2}\varepsilon x}}=\varepsilon+\frac{\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}(\bm{X}-\varepsilon)x}\mathbb{I}_{\{\bm{X}>\varepsilon\}}]}{p_{\varepsilon}}.$
Note that for any $x\geq 1$, $\bm{X}e^{-\frac{1}{2}(\bm{X}-\varepsilon)x}\mathbb{I}_{\{\bm{X}>\varepsilon\}}$ is dominated by $\bm{X}e^{-\frac{1}{2}(\bm{X}-\varepsilon)}\mathbb{I}_{\{\bm{X}>\varepsilon\}}$, which is integrable regardless of whether $\tilde{\mathbb{E}}[\bm{X}]$ is finite. Hence, by the dominated convergence theorem, $\limsup\limits_{x\to\infty}Q(x)\leq\varepsilon$. Because $\varepsilon$ is arbitrary, we have $\lim\limits_{x\to\infty}Q(x)=0$. By (A.2), the convergence result for $h$ follows immediately. ∎
Lemma A.3 (Uniform Convergence).
Let $\{\bm{R}_{n}\}_{n\geq 1}$ be a sequence of nonnegative random variables that converges in distribution to $\bm{R}$ and $\lim\limits_{n\to\infty}\tilde{\mathbb{E}}[\bm{R}_{n}]=\tilde{\mathbb{E}}[\bm{R}]\in(0,\infty)$. Then, on every compact subset of $0,\infty)$, the sequence $\{h_{\bm{R}_{n}}\}_{n\geq 1}$ converges uniformly to $h_{\bm{R}}$.
Proof.
Fix a compact subset $\mathcal{K}\subset[0,\infty)$.
We first show that $\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}_{n}x}]\to\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}x}]$ uniformly on $\mathcal{K}$. Observe that
$\left|\dfrac{\mathrm{d}\left(\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}_{n}x}]\right)}{\mathrm{d}x}\right|=\left|-\frac{1}{2}\tilde{\mathbb{E}}[\bm{R}_{n}e^{-\frac{1}{2}\bm{R}_{n}x}]\right|\leq\frac{1}{2}\tilde{\mathbb{E}}[\bm{R}_{n}]\quad\text{for all}\,\,x\geq 0.$
Because $\tilde{\mathbb{E}}[\bm{R}_{n}]\to\tilde{\mathbb{E}}[\bm{R}]$, the sequence $\{\tilde{\mathbb{E}}[\bm{R}_{n}]\}_{n\geq 1}$ is bounded. Hence, the family of functions $\{\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}_{n}x}]\}_{n\geq 1}$ is equicontinuous on $\mathcal{K}$. Moreover, because $\bm{R}_{n}\xrightarrow{d}\bm{R}$ and the function $r\mapsto e^{-\frac{1}{2}rx}$ is bounded and continuous, we have $\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}_{n}x}]\to\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}x}]$ for any $x\geq 0$. It follows that $\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}_{n}x}]\to\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}x}]$ uniformly on $\mathcal{K}$; see, e.g., rudin1976principles.
Now we show the uniform convergence of $\{\tilde{\mathbb{E}}[\bm{R}_{n}e^{-\frac{1}{2}\bm{R}_{n}x}]\}_{n\geq 1}$ on $\mathcal{K}$. By $\bm{R}_{n}\xrightarrow{d}\bm{R}$ and $\tilde{\mathbb{E}}[\bm{R}_{n}]\to\tilde{\mathbb{E}}[\bm{R}]<\infty$, we know that the family $\{\bm{R}_{n}\}_{n\geq 1}$ is uniformly integrable. Consequently, for any $\varepsilon>0$, there exists $M>0$ such that $F_{\bm{R}}$ is continuous at $M$, $\sup_{n}\tilde{\mathbb{E}}[\bm{R}_{n}\mathbb{I}_{\{\bm{R}_{n}>M\}}]<\frac{\varepsilon}{3}$, and $\tilde{\mathbb{E}}[\bm{R}\mathbb{I}_{\{\bm{R}>M\}}]<\frac{\varepsilon}{3}$. Let $f_{n}^{M}(x)\triangleq\tilde{\mathbb{E}}[\bm{R}_{n}e^{-\frac{1}{2}\bm{R}_{n}x}\mathbb{I}_{\{\bm{R}_{n}\leq M\}}]$. Then $f_{n}^{M}$ is continuously differentiable and its derivative is uniformly bounded by $\frac{1}{2}M^{2}$, which implies that the family $\{f_{n}^{M}\}_{n\geq 1}$ is equicontinuous on $\mathcal{K}$. Furthermore, because $\bm{R}_{n}\xrightarrow{d}\bm{R}$, we have the pointwise convergence $f_{n}^{M}(x)\to f^{M}(x)\triangleq\tilde{\mathbb{E}}[\bm{R}e^{-\frac{1}{2}\bm{R}x}\mathbb{I}_{\{\bm{R}\leq M\}}]$; see, e.g., follmer2016stochastic. As a result, $f_{n}^{M}$ converges uniformly to $f^{M}$ on $\mathcal{K}$. Hence, for all sufficiently large $n$,
$\sup_{x\in\mathcal{K}}|f_{n}^{M}(x)-f^{M}(x)|<\frac{\varepsilon}{3}.$
Observe that
$|\tilde{\mathbb{E}}[\bm{R}_{n}e^{-\frac{1}{2}\bm{R}_{n}x}]-\tilde{\mathbb{E}}[\bm{R}e^{-\frac{1}{2}\bm{R}x}]|\leq|f_{n}^{M}(x)-f^{M}(x)|+\tilde{\mathbb{E}}[\bm{R}_{n}\mathbb{I}_{\{\bm{R}_{n}>M\}}]+\tilde{\mathbb{E}}[\bm{R}\mathbb{I}_{\{\bm{R}>M\}}].$
Therefore, for $n$ large enough,
$\sup_{x\in\mathcal{K}}\left|\tilde{\mathbb{E}}[\bm{R}_{n}e^{-\frac{1}{2}\bm{R}_{n}x}]-\tilde{\mathbb{E}}[\bm{R}e^{-\frac{1}{2}\bm{R}x}]\right|<\varepsilon,$
which implies the uniform convergence of $\{\tilde{\mathbb{E}}[\bm{R}_{n}e^{-\frac{1}{2}\bm{R}_{n}x}]\}_{n\geq 1}$ on $\mathcal{K}$.
Finally, because $\tilde{\mathbb{E}}[\bm{R}]>0$, the function $\tilde{\mathbb{E}}[\bm{R}e^{-\frac{1}{2}\bm{R}x}]$ is bounded away from zero on $\mathcal{K}$. Therefore, $\{h_{\bm{R}_{n}}\}_{n\geq 1}$ converges uniformly to $h_{\bm{R}}$ on $\mathcal{K}$. ∎
The reverse hazard rate order can be characterized via expectation ratios and connected to first-order stochastic dominance under weighted measures, as established by caperaa1988tail (see also SS2007) and wang2024weighted.
Lemma A.4 (caperaa1988tail).
Let $\bm{R_{1}}$ and $\bm{R_{2}}$ be two random variables. Then $R_{1}\succeq_{rh}R_{2}$ if and only if
$\frac{\tilde{\mathbb{E}}\left[u(\bm{R_{1}})w(\bm{R_{1}})\right]}{\tilde{\mathbb{E}}[w(\bm{R_{1}})]}\geq\frac{\tilde{\mathbb{E}}[u(\bm{R_{2}})w(\bm{R_{2}})]}{\tilde{\mathbb{E}}[w(\bm{R_{2}})]},$
for all functions $u$ and $w$ for which the expectations exist, such that $w$ is positive, continuous, and decreasing, and $u$ is continuous and strictly increasing.
Lemma A.5 (wang2024weighted).
Let $\bm{R_{1}}$ and $\bm{R_{2}}$ be two random variables with cumulative distribution functions $F_{\bm{R_{1}}}$ and $F_{\bm{R_{2}}}$, respectively. Assume that $\bm{R_{1}}\succeq_{rh}\bm{R_{2}}$. Let $w(\cdot)$ be a positive, continuous, and strictly decreasing function. Define the weighted probability measures $\tilde{F}_{i}$ associated with $\bm{R}_{i}$ for $i=1,2$ by
$\mathrm{d}\tilde{F}_{i}(x)=\frac{w(x)\mathrm{d}F_{\bm{R_{i}}}}{\tilde{\mathbb{E}}\left[w(\bm{R_{i}})\right]}.$ (A.3)
Then, $\bm{R}_{1}$ dominates $\bm{R}_{2}$ under the weighted measures in the sense of first-order stochastic dominance, i.e.,
$\tilde{F}_{1}\succeq_{1}\tilde{F}_{2}.$
[…]
Lemma A.6.
For fixed $x>0$, $h(x)$ is strictly decreasing with respect to $\bm{R}$ in the sense of the reverse-hazard-rate order among distinct distributions.
Proof.
Fix $x>0$. Let $\bm{R_{1}}$ and $\bm{R_{2}}$} be two non-negative random variables such that $\bm{R_{1}}\succeq_{rh}\bm{R_{2}}$ and their distributions are distinct. For notational simplicity, in the sequel, let $h_{i}(\cdot)$ and $H_{i}(\cdot)$ denote the function $h(\cdot)$ and the function defined in Theorem 3.2 corresponding to the random variable $\bm{R}_{i}$ for $i=1,2$, respectively. According to Lemma A.4, let $u(\gamma)=\gamma$ and $w(\gamma)=\exp(-\frac{1}{2}\gamma x)$. Note that $u(\gamma)$ is a strictly increasing function, and for $x>0$, $w(\gamma)$ is a positive, continuous, and strictly decreasing function of $\gamma$. Substituting them into this inequality, we have
$\frac{\tilde{\mathbb{E}}[\bm{R_{1}}\exp(-\frac{1}{2}\bm{R_{1}}x)]}{\tilde{\mathbb{E}}[\exp(-\frac{1}{2}\bm{R_{1}}x)]}\geq\frac{\tilde{\mathbb{E}}[\bm{R_{2}}\exp(-\frac{1}{2}\bm{R_{2}}x)]}{\tilde{\mathbb{E}}[\exp(-\frac{1}{2}\bm{R_{2}}x)]}.$
Moreover, assume that the equality holds. Define the weighted probability measures $\tilde{F}_{i}$ for $i=1,2$ by (A.3). The given condition is equivalent to $\tilde{\mathbb{E}}_{\tilde{F}_{1}}[\bm{R_{1}}]=\tilde{\mathbb{E}}_{\tilde{F}_{2}}[\bm{R_{1}}]$. From Lemma A.5, we have $\tilde{F}_{1}\succeq_{1}\tilde{F}_{2}$ and consequently, $\mathrm{d}\tilde{F}_{1}(\gamma)=\mathrm{d}\tilde{F}_{2}(\gamma)$. Given that $w(\gamma)$ is positive, this implies $\frac{\mathrm{d}\tilde{F}_{1}}{\tilde{\mathbb{E}}\left[w(\bm{R_{1}})\right]}=\frac{\mathrm{d}\tilde{F}_{2}}{\tilde{\mathbb{E}}\left[w(\bm{R_{2}})\right]}$. Integrating both sides over the entire support and using the property $\int\mathrm{d}\tilde{F}_{i}(\gamma)=1$, we obtain $1=\frac{\tilde{\mathbb{E}}\left[w(\bm{R_{1}})\right]}{\tilde{\mathbb{E}}\left[w(\bm{R_{2}})\right]}$, which implies $\tilde{\mathbb{E}}\left[w(\bm{R_{1}})\right]=\tilde{\mathbb{E}}\left[w(\bm{R_{2}})\right]$. Substituting this back yields $\mathrm{d}F_{\bm{R_{1}}}(\gamma)=\mathrm{d}F_{\bm{R_{2}}}(\gamma)$, confirming that $F_{\bm{R_{1}}}(\gamma)=F_{\bm{R_{2}}}(\gamma)$. This contradicts the different distributions of $\bm{R_{1}}$ and $\bm{R_{2}}$. Thus, we have
$\frac{\tilde{\mathbb{E}}[\bm{R_{1}}\exp(-\frac{1}{2}\bm{R_{1}}x)]}{\tilde{\mathbb{E}}[\exp(-\frac{1}{2}\bm{R_{1}}x)]}>\frac{\tilde{\mathbb{E}}[\bm{R_{2}}\exp(-\frac{1}{2}\bm{R_{2}}x)]}{\tilde{\mathbb{E}}[\exp(-\frac{1}{2}\bm{R_{2}}x)]},$
and equivalently, $h_{1}(x)<h_{2}(x)$. ∎
Appendix B Proofs
B.1 Proof of Proposition 3.8
Assume that the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly increasing and $\Lambda_{n}\nearrow H(\infty)$. Because the convergence is strictly increasing toward $H(\infty)$, we have $\Lambda_{n}<H(\infty)$ for all $n\geq 1$, ensuring the existence of the equilibrium. Thus, for each $n$, $v_{n}(0)$ is well-defined and uniquely determined by $v_{n}(0)=H^{-1}(\Lambda_{n})$. Consequently, the convergence $\Lambda_{n}\to H(\infty)$ implies that $\lim_{n\to\infty}v_{n}(0)=\infty$. By Theorem 3.2, we have $a_{n}(0)=h(v_{n}(0))\lambda_{0}$. Using the asymptotic property of $h$ established in Lemma A.2, namely $\lim_{x\to\infty}h(x)=\frac{1}{r_{0}}$, we obtain
$\lim_{n\to\infty}a_{n}(0)=\lambda_{0}\lim_{v_{n}(0)\to\infty}h(v_{n}(0))=\frac{\lambda_{0}}{r_{0}}.$
Furthermore, because $\{\Lambda_{n}\}$ is strictly increasing, $\{v_{n}(0)\}$ is strictly increasing. Combined with the monotonicity of $h$ (see Lemma A.1), it follows that $|a_{n}(0)|=|\lambda_{0}|h(v_{n}(0))$ converges monotonically increasing to $\frac{|\lambda_{0}|}{r_{0}}$.
Assume that the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly decreasing and $\Lambda_{n}\searrow 0$. Because $H(\infty)>0$, and $\Lambda_{n}\searrow 0$, there exists an integer $N$ such that for all $n>N$, $\Lambda_{n}<H(\infty)$. Thus, the equilibrium exists for all sufficiently large $n$. For such $n$, we have $v_{n}(0)=H^{-1}(\Lambda_{n})$. Because $H(0)=0$ and $H^{-1}$ is continuous, we have
$\lim_{n\to\infty}v_{n}(0)=H^{-1}(0)=0.$
Recall that $h(0)=\frac{1}{\tilde{\mathbb{E}}[\bm{R}]}$. Taking the limit of $a_{n}(0)$, we have
$\lim_{n\to\infty}a_{n}(0)=\lambda_{0}\lim_{v_{n}(0)\to 0}h(v_{n}(0))=\frac{\lambda_{0}}{\tilde{\mathbb{E}}[\bm{R}]}.$
Finally, because $\{\Lambda_{n}\}$ is strictly decreasing, $\{v_{n}(0)\}$ is strictly decreasing. The strict monotonicity of $h$ implies that $|a_{n}(0)|$ converges monotonically decreasing to $\frac{|\lambda_{0}|}{\tilde{\mathbb{E}}[\bm{R}]}$.
B.2 Proof of Theorem 3.11
For notational simplicity, let $h_{n}=h_{\bm{R}_{n}}$, $H_{n}=H_{\bm{R}_{n}}$, and $v_{n}=v_{\bm{R}_{n}}$ ($n\geq 1$). By Lemma A.3, the sequence $\{h_{n}\}_{n\geq 1}$ converges uniformly to $h$ on every compact subset of $0,\infty)$. Because the limit function $h$ is strictly positive, it follows that $\{H_{n}\}_{n\geq 1}$ converges uniformly to $H$ on every compact subset of $0,\infty)$ as well. By Lemma B.1 below, the sequence of the inverse functions $\{H_{n}^{-1}\}_{n\geq 1}$ converge uniformly to $H^{-1}$ on $0,\Lambda(0)]$.
We now see the uniform convergence of $\{v_{n}\}_{n\geq 1}$ to $v$ on $[0,T]$:
$\sup_{t\in[0,T]}|v_{n}(t)-v(t)|=\sup_{t\in[0,T]}|H_{n}^{-1}(\Lambda(t))-H^{-1}(\Lambda(t))|\leq\sup_{y\in[0,\Lambda(0)]}|H_{n}^{-1}(y)-H^{-1}(y)|\to 0.$
Next, we show the uniform convergence of the composite sequence $h_{n}(v_{n}(\cdot))$. Indeed, by the triangle inequality,
$\sup_{t\in[0,T]}|h_{n}(v_{n}(t))-h(v(t))|\leq\sup_{t\in[0,T]}|h_{n}(v_{n}(t))-h(v_{n}(t))|+\sup_{t\in[0,T]}|h(v_{n}(t))-h(v(t))|.$ (B.1)
Let $\mathcal{K}\triangleq\left[0,\max_{t\in[0,T]}|v(t)|+1\right]$. Then $v_{n}(t)\in\mathcal{K}$ for all $t\in[0,T]$ and all sufficiently large $n$. Therefore, the uniform convergence of $\{h_{n}\}_{n\geq 1}$ on $\mathcal{K}$ implies $\sup_{t\in[0,T]}|h_{n}(v_{n}(t))-h(v_{n}(t))|\to 0$. Moreover, $h$ is continuous and therefore uniformly continuous on the compact set $\mathcal{K}$, which combined with the fact that $\sup_{t\in[0,T]}|v_{n}(t)-v(t)|\to 0$ implies $\sup_{t\in[0,T]}|h(v_{n}(t))-h(v(t))|\to 0$. Consequently, by (B.1), we have
$\lim_{n\to\infty}\sup_{t\in[0,T]}|h_{n}(v_{n}(t))-h(v(t))|=0.$ (B.2)
Finally, recalling that $a_{n}(t)=h_{n}(v_{n}(t))\lambda(t)$ and $a(t)=h(v(t))\lambda(t)$, we get from (B.2) that $\lim_{n\to\infty}a_{n}(t)=a(t)$ for every $t\in[0,T)$. Furthermore, if $\lambda\in L^{\infty}([0,T])$, then the above convergence is uniform on $[0,T)$.
Lemma B.1.
Let $H$ and $H_{n}$ ($n\geq 1$) be strictly increasing and continuous functions defined on $0,\infty)$ with $H(0)=H_{n}(0)=0$ for all $n\geq 1$. Suppose that $\{H_{n}\}_{n\geq 1}$ converges uniformly to $H$ on every compact subset of $0,\infty)$. Then, for any bounded interval $0,z]\subset[0,H(\infty))$, there exists an integer $N$ such that for all $n\geq N_{0}$, the inverse function $H_{n}^{-1}$ is well-defined on $[0,z]$, and the sequence $\{H_{n}^{-1}\}_{n\geq N_{0}}$ converges uniformly to $H^{-1}$ on $[0,z]$.
Proof.
Fix $0,z]\subset[0,H(\infty))$ and a constant $M>H^{-1}(z)$. By the uniform convergence of $\{H_{n}\}_{n\geq 1}$ on $[0,M]$, there exists $N_{0}\geq 1$ such that $H_{n}(M)>z$ for all $n\geq N_{0}$. Because each $H_{n}$ is strictly increasing and continuous with $H_{n}(0)=0$, it follows that $0,z]\subset[0,H_{n}(M)]=H_{n}([0,M])$ for all $n\geq N_{0}$. Thus, for all $n\geq N_{0}$, the inverse function $H_{n}^{-1}$ is well-defined on $[0,z]$, and moreover, for any $x\in[0,z]$, we have $H_{n}^{-1}(x)\in[0,M]$ and $H^{-1}(x)\in[0,M]$.
Next, because $H^{-1}$ is continuous on the compact interval $[0,H(M)]$, it is uniformly continuous. Thus, for any $\varepsilon>0$, there exists $\delta>0$ such that for any $x^{\prime},x^{\prime\prime}\in[0,H(M)]$ with $|x^{\prime}-x^{\prime\prime}|<\delta$, we have
$|H^{-1}(x^{\prime})-H^{-1}(x^{\prime\prime})|<\varepsilon.$ (B.3)
By uniform convergence of $H_{n}\to H$ on $[0,M]$, there exists $N\geq N_{0}$ such that for all $n\geq N$ and all $y\in[0,M]$,
$|H_{n}(y)-H(y)|<\delta.$
For any $x\in[0,z]$, let $x_{n}=H_{n}^{-1}(x)$. Then, $x_{n}\in[0,M]$ and
$|H(x_{n})-x|=|H(x_{n})-H_{n}(x_{n})|<\delta.$
Because $x,H(x_{n})\in[0,H(M)]$ and $x_{n}=H^{-1}(H(x_{n}))$, inequality (B.3) yields
$|H_{n}^{-1}(x)-H^{-1}(x)|=|x_{n}-H^{-1}(x)|=|H^{-1}(H(x_{n}))-H^{-1}(x)|<\varepsilon.$
As this holds uniformly for all $x\in[0,z]$, the convergence is uniform. ∎
B.3 Proof of Proposition 5.3
We first establish a technical lemma.
Lemma B.2.
Let $\bm{R}_{1}$ and $\bm{R}_{2}$ be two non-negative random variables satisfying the conditions (1) and (2) in Proposition 5.3. Recall the definition in Proposition 3.8 that $\bm{X}_{i}=\bm{R}_{i}-r_{0}$ and $Q_{i}(x)=\frac{\tilde{\mathbb{E}}[\bm{X}_{i}e^{-\frac{1}{2}\bm{X}_{i}x}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}_{i}x}]}$, $x\geq 0$. Then, the following assertions hold:
  1. (1)
    For each $i\in\{1,2\}$, there exist constant $M_{i}>0$, such that $Q_{i}(x)\leq M_{i}e^{-\frac{1}{2}\delta_{i}x}$ for any $x\geq 0$. In particular, $Q_{i}(x)\to 0$ as $x\to\infty$.
  2. (2)
    For each $i\in\{1,2\}$, the integral $\int_{0}^{\infty}(h_{i}^{-2}(s)-r_{0}^{2})ds$ converges to a finite constant $K_{i}$.
  3. (3)
    For any $x_{1},x_{2}\geq 0$ and any $\varepsilon>0$, it holds that $\dfrac{Q_{1}(x_{1})}{Q_{2}(x_{2})}\leq C\dfrac{e^{-\frac{1}{2}\delta_{1}x_{1}}}{e^{-\frac{1}{2}(\delta_{2}+\varepsilon)x_{2}}}$, where $C$ is a constant depending on $\varepsilon$.
Proof.
For $i\in\{1,2\}$. By condition (1) in Proposition 5.3, we have $\mathbb{P}\mathbb{(}\bm{X}_{i}=0)=p_{i}$. The Dominated Convergence Theorem then implies $\lim_{x\to\infty}\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}_{i}x}]=p_{i}\in(0,1)$. Hence, there exists $m_{i}>0$ such that $\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}_{i}x}]\geq m_{0}$ for any $x\geq 0$. Moreover, by condition (2) in Proposition 5.3, the support of $\bm{X}_{i}$ on $(0,\infty)$ is contained in $\delta_{i},\infty)$. Therefore,
$\tilde{\mathbb{E}}[\bm{X}_{i}e^{-\frac{1}{2}\bm{X}_{i}x}]=\int_{\delta_{i}}^{\infty}\gamma e^{-\frac{1}{2}\gamma x}\mathrm{d}F_{\bm{X}_{i}}(\gamma)\leq e^{-\frac{1}{2}\delta_{i}x}\int_{\delta_{i}}^{\infty}\gamma\mathrm{d}F_{\bm{X}_{i}}(\gamma)=e^{-\frac{1}{2}\delta_{i}x}\tilde{\mathbb{E}}[\bm{X}_{i}\mathbb{I}_{\{\bm{X}_{i}\geq\delta_{i}\}}].$
The desired conclusion then follows by taking $M_{i}\triangleq\frac{\tilde{\mathbb{E}}[\bm{X}_{i}\mathbb{I}_{\{\bm{X}_{i}\geq\delta_{i}\}}]}{m_{i}}$.
For $i\in\{1,2\}$, based on (A.2), the conclusion follows immediately from assertion (1) and the fact that $\delta_{i}>0$.
For any $\varepsilon>0$. By the definition of the support, we have $p_{\varepsilon,2}\triangleq P(\delta_{2}\leq\bm{X_{2}}<\delta_{2}+\varepsilon)>0$. Restricting the integration to this interval yields
$\tilde{\mathbb{E}}[\bm{X_{2}}e^{-\frac{1}{2}\bm{X_{2}}x}]\geq\int_{\delta_{2}}^{\delta_{2}+\varepsilon}xe^{-\frac{1}{2}x\gamma}\mathrm{d}F_{\bm{X_{2}}}(\gamma)\geq\delta_{2}e^{-\frac{1}{2}(\delta_{2}+\varepsilon)x}p_{\varepsilon,2}.$
Combining this with assertion (1) gives the desired bound with $C\triangleq\frac{M_{1}}{\delta_{2}p_{\varepsilon,2}}$. ∎
Now we return to prove Proposition B.3.
By assertion (1) of Lemma B.2 and (A.2), we have $\lim\limits_{x\to\infty}h_{i}(x)=\frac{1}{r_{0}}$. Because $r_{0}>0$, it follows that the integral $\int_{0}^{\infty}h_{i}^{-2}(x)\mathrm{d}x$ diverges, which implies $H_{i}(\infty)=\infty$. Consequently, Theorem 3.2 (1) ensures the existence of a unique equilibrium $\pi_{i}$ for $\bm{R}_{i}$ for any $T>0$. Based on the relation ${\pi}_{i}=(\sigma^{\top})^{-1}a_{i}$, it suffices to compare $a_{i}$. By (2.5), we have $|a_{i}(0,T)|=h_{i}(v_{i}(0,T))|\lambda(0)|$.
First, we establish the asymptotic relationship between $v_{1}(0,T)$ and $v_{2}(0,T)$ as $T\to\infty$. By Theorem 3.2 (1), we have $v_{i}(0,T)=H_{i}^{-1}(\Lambda(0,T))$. It then follows that $\lim_{T\to\infty}v_{i}(0,T)=\infty$, because $H_{i}(\infty)=\infty$ and $\lim_{T\to\infty}\Lambda(0,T)=\infty$. Moreover, the relation $v_{i}(0,T)=H_{i}^{-1}(\Lambda(0,T))$ yields
$\Lambda(0,T)=\int_{0}^{v_{i}(0,T)}(r_{0}^{2}+(h_{i}^{-2}(x)-r_{0}^{2}))\mathrm{d}x=r_{0}^{2}v_{i}(0,T)+\int_{0}^{v_{i}(0,T)}(h_{i}^{-2}(x)-r_{0}^{2})\mathrm{d}x.$ (B.4)
Define the tail integral $R_{i}(y)\triangleq\int_{y}^{\infty}(h_{i}^{-2}(x)-r_{0}^{2})\mathrm{d}x$, which satisfies $\lim_{y\to\infty}R_{i}(y)=0$ by Lemma B.2 (2). Then (B.4) can be rewritten as
$\Lambda(0,T)=r_{0}^{2}v_{i}(0,T)+K_{i}-R_{i}(v_{i}(0,T)).$ (B.5)
Equating (B.5) for $i=1$ and $i=2$ gives
$v_{1}(0,T)=v_{2}(0,T)+\dfrac{K_{2}-K_{1}}{r_{0}^{2}}+\dfrac{R_{2}(v_{2}(0,T))-R_{1}(v_{1}(0,T))}{r_{0}^{2}}.$
Let $\Delta K=\dfrac{K_{2}-K_{1}}{r_{0}^{2}}$ and $\xi(T)=\dfrac{R_{2}(v_{2}(0,T))-R_{1}(v_{1}(0,T))}{r_{0}^{2}}$. Because $v_{i}(0,T)\to\infty$ as $T\to\infty$, the tail difference term $\xi(T)$ tends to 0. Consequently, we obtain the asymptotic relation $v_{1}(0,T)=v_{2}(0,T)+\Delta K+\xi(T)$.
Second, we compare $h_{1}(v_{1}(0,T))$ and $h_{2}(v_{2}(0,T))$. Because $h_{i}(s)=(r_{0}+Q_{i}(s))^{-1}$, this reduces to comparing $Q_{1}(v_{1}(0,T))$ and $Q_{2}(v_{2}(0,T))$. Let $\varepsilon>0$ be a constant satisfying $\delta_{1}-\delta_{2}-\varepsilon>0$. Using the bound established in Lemma B.2 (3), we obtain
$\dfrac{Q_{1}(v_{1}(0,T))}{Q_{2}(v_{2}(0,T))}\leq C\dfrac{e^{-\frac{1}{2}\delta_{1}v_{1}(0,T)}}{e^{-\frac{1}{2}(\delta_{2}+\varepsilon)v_{2}(0,T)}}.$
Substituting $v_{1}(0,T)=v_{2}(0,T)+\Delta K+\xi(T)$ into the exponent yields
$\dfrac{Q_{1}(v_{1}(0,T))}{Q_{2}(v_{2}(0,T))}\leq Ce^{-\frac{1}{2}\delta_{1}(\Delta K+\xi(T))}\cdot e^{-\frac{1}{2}(\delta_{1}-\delta_{2}-\varepsilon)v_{2}(0,T)}.$
Because $\lim_{T\to\infty}\xi(T)=0$, the first exponential term converges to a finite positive constant. Moreover, as $v_{2}(0,T)\to\infty$, the second term tends to 0, and hence $\frac{Q_{1}(v_{1}(0,T))}{Q_{2}(v_{2}(0,T))}\to 0$ as $T\to\infty$.
This implies for sufficiently large $T$, $Q_{1}(v_{1}(0,T))<Q_{2}(v_{2}(0,T))$. Because $h_{i}(x)=(r_{0}+Q_{i}(x))^{-1}$ is strictly decreasing in $Q_{i}$, we have $h_{1}(v_{1}(0,T))>h_{2}(v_{2}(0,T))$. Consequently, $|a_{1}(0,T)|>|a_{2}(0,T)|$ because $\lambda(0)\neq 0$, and equivalently, $|\pi_{1}(0,T)|>|\pi_{2}(0,T)|$.
B.4 Proof of Proposition 5.4
For $i=1,2$, let $\mathbb{P}(\bm{R}_{i}=r_{0})=p_{i}$. The function $Q_{i}(\cdot)$ is explicitly given by
$Q_{i}(x)=\frac{\tilde{\mathbb{E}}[\bm{X}_{i}e^{-\bm{X}_{i}\frac{x}{2}}]}{\tilde{\mathbb{E}}[e^{-\bm{X}_{i}\frac{x}{2}}]}=\frac{\delta_{i}(1-p_{i})e^{-\delta_{i}\frac{x}{2}}}{p_{i}+(1-p_{i})e^{-\delta_{i}\frac{x}{2}}},\quad x\in[0,\infty).$
We first verify the existence of a solution to $|a_{1}(t)|=|a_{2}(t)|$. This equality is equivalent to $Q_{1}(v_{1}(t))=Q_{2}(v_{2}(t))$, because $|a_{i}(t)|=|\lambda(t)|(r_{0}+Q_{i}(v_{i}(t)))^{-1}$. Define $\mathcal{D}(t)\triangleq Q_{1}(v_{1}(t))-Q_{2}(v_{2}(t))$. At $t=0$, by assumption, $\mathcal{D}(0)<0$. At $t=T$, because $v_{i}(T)=0$ and $Q_{i}(0)=\tilde{\mathbb{E}}[\bm{R}_{i}]-r_{0}$, the condition $\tilde{\mathbb{E}}[\bm{R}_{1}]>\tilde{\mathbb{E}}[\bm{R}_{2}]$ implies $Q_{1}(0)>Q_{2}(0)$, and therefore $\mathcal{D}(T)>0$. Because $\mathcal{D}(t)$ is continuous, there exists at least one $t^{*}\in(0,T)$ such that $\mathcal{D}(t^{*})=0$.
Next, we show uniqueness by differentiating $\mathcal{D}(\cdot)$. To this end, we compute $Q_{i}^{\prime}(x)$:
$Q_{i}^{\prime}(x)$ $=\frac{-\frac{\delta_{i}^{2}}{2}(1-p_{i})e^{-\delta_{i}\frac{x}{2}}\left(p_{i}+(1-p_{i})e^{-\delta_{i}\frac{x}{2}}\right)-\delta_{i}(1-p_{i})e^{-\delta_{i}\frac{x}{2}}\left(-\frac{\delta_{i}}{2}(1-p_{i})e^{-\delta_{i}\frac{x}{2}}\right)}{\left(p_{i}+(1-p_{i})e^{-\delta_{i}\frac{x}{2}}\right)^{2}}$
$=-\frac{\delta_{i}}{2}\left(\frac{\delta_{i}(1-p_{i})e^{-\delta_{i}\frac{x}{2}}}{p_{i}+(1-p_{i})e^{-\delta_{i}\frac{x}{2}}}\right)+\frac{1}{2}\left(\frac{\delta_{i}(1-p_{i})e^{-\delta_{i}\frac{x}{2}}}{p_{i}+(1-p_{i})e^{-\delta_{i}\frac{x}{2}}}\right)^{2}$
$=-\frac{\delta_{i}}{2}Q_{i}(x)+\frac{1}{2}Q_{i}(x)^{2}$
$=\frac{1}{2}Q_{i}(x)(Q_{i}(x)-\delta_{i}).$
Differentiating $\mathcal{D}(t)$ and using $v_{i}^{\prime}(t)=-|\lambda(t)|^{2}(r_{0}+Q_{i}(v_{i}(t)))^{-2}$ yields
$\mathcal{D}^{\prime}(t)$ $=Q_{1}^{\prime}(v_{1})v_{1}^{\prime}-Q_{2}^{\prime}(v_{2})v_{2}^{\prime}$
$=\frac{1}{2}Q_{1}(Q_{1}-\delta_{1})\left(\frac{-|\lambda(t)|^{2}}{(r_{0}+Q_{1})^{2}}\right)-\frac{1}{2}Q_{2}(Q_{2}-\delta_{2})\left(\frac{-|\lambda(t)|^{2}}{(r_{0}+Q_{2})^{2}}\right)$
$=\frac{|\lambda(t)|^{2}}{2}\left[\frac{Q_{2}(Q_{2}-\delta_{2})}{(r_{0}+Q_{2})^{2}}-\frac{Q_{1}(Q_{1}-\delta_{1})}{(r_{0}+Q_{1})^{2}}\right].$
At any solution $t^{*}$ where $\mathcal{D}(t^{*})=0$, we have $Q=Q_{1}(v_{1}(t^{*}))=Q_{2}(v_{2}(t^{*}))>0$. Substituting $Q_{1}=Q_{2}=Q$ yields
$\mathcal{D}^{\prime}(t^{*})=\frac{|\lambda(t^{*})|^{2}Q}{2(r_{0}+Q)^{2}}(\delta_{1}-\delta_{2}).$
Because $\delta_{1}>\delta_{2}$ and $\lambda(t^{*})\neq 0$, we have $\mathcal{D}^{\prime}(t^{*})>0$. Therefore, the solution $t^{*}$ is unique.

### v3

@@
 Abstract
-This paper investigates infinite-dimensional portfolio selection problem under a general distribution of the risk aversion parameter.
-We provide a complete characterization of all deterministic equilibrium investment strategies.
-Our results reveal that the solution structure depends critically on the distribution of risk aversion: the equilibrium is unique whenever it exists in the case of finite expected risk aversion, whereas an infinite expectation can lead to infinitely many equilibria or to a unique trivial one ($\bar{\pi}\equiv\mathbf{0}$).
-To address this multiplicity, we introduce three optimality criteria—optimal, uniformly optimal, and uniformly strictly optimal—and explicitly characterize the existence and uniqueness of the corresponding equilibria.
-Under the same necessary and sufficient condition, the optimal and uniformly optimal equilibria exist uniquely and coincide.
-Furthermore, assuming that the market price of risk is non-zero near the terminal time, we show that the optimal (and hence uniformly optimal) equilibrium is also uniformly strictly optimal.
-Finally, we perform comparative statics to demonstrate that a risk aversion distribution dominating another in the reverse hazard rate order leads to a less aggressive equilibrium strategy.
+This paper studies a continuous-time portfolio selection problem under a general distribution of random risk aversion (RRA).
+We provide a complete characterization of all deterministic equilibrium strategies in closed form.
+Our results show that the structure of the solution depends crucially on the distribution of RRA: the equilibrium is unique (if exits) when the expectation of RRA is finite, whereas an infinite expectation leads either to infinitely many equilibria or to a unique trivial one (i.e.
+risk-free investment).
+To resolve this multiplicity of equilibria, we select, among all deterministic equilibria, the one that maximizes the objective functional at the initial time.
+We establish a necessary and sufficient condition for the existence of such an optimal equilibrium, which is then shown to be unique and uniformly optimal.
+Finally, we conduct a comparative statics.
+Using counterexamples based on two-point distributed RRA, we demonstrate that a larger risk aversion in the sense of first-order stochastic dominance does not necessarily lead to less risky investment.
+Within the two-point distribution framework, we further examine the single-crossing property of equilibrium strategies and the monotonicity of the crossing time.
+We show that a larger risk aversion under a stronger stochastic order—the reverse hazard rate order—always leads to less risky investment.
+In addition, we analyze how the convex combination of independent and identically distributed RRAs influences investment.
 Keywords: Random Risk Aversion; Time-Inconsistency; Portfolio Selection; Intra-Personal Equilibrium; Multiple Equilibria; Comparative Statics
 1 Introduction
-The classical framework for portfolio selection, pioneered by merton1969lifetime; merton1971 and samuelson1969lifetime, is based on the assumption that an investor’s risk aversion can be captured by a single known constant.
-In practice, however, this coefficient is very difficult for an agent to specify with certainty, and a growing body of empirical evidence suggests that it varies dynamically with market conditions, often increasing during financial crises and bear markets, as shown in studies by gordon2000preference; chetty2006new and guiso2018time.
-These challenges have motivated a move towards models that incorporate random risk aversion.
-Within this paradigm, a more economically coherent approach, as advanced by Desmettre2023, is to aggregate certainty equivalents over the distribution of the risk aversion parameter.
-While the random risk aversion in Desmettre2023 provides a consistent performance measure in monetary units, it introduces a nonlinearity in expectation that naturally leads to the time inconsistency of the preference.
-The first systematic treatment of time inconsistency was given by Strotz1955, who proposed the notion of consistent planning, laying the foundation for the intra-personal game-theoretic framework.
-Building on this idea, ekeland2006being later developed a rigorous continuous-time formalization, which in turn inspired extensive research on continuous-time control with time-inconsistent preferences.
-For example, bjork2010general (published version Bjork2017) established a general theoretical framework by deriving an extended HJB equation and hu2012time; Hu2017 investigated a time-inconsistent stochastic linear–quadratic control problem.
-For further discussions and related developments, see, e.g., yan2019time, he2021equilibrium, Hernandez2023, and the references therein.
-Within the Black-Scholes market model, Desmettre2023 provide a verification theorem on the intra-personal equilibrium strategies for a CRRA utility function with a general random risk aversion and, in the case of a binomial random risk aversion, characterize the equilibrium by a three-dimensional ODE system without establishing the existence and uniqueness of the solution.
+The classical framework for portfolio selection, pioneered by Merton (1969, 1971) and Samuelson (1969), is based on the assumption that an investor’s risk aversion can be captured by a single known constant.
+In practice, however, this coefficient is very difficult for an agent to specify with certainty, and a growing body of empirical evidence suggests that it varies dynamically with market conditions, often increasing during financial crises and bear markets, as shown in studies by Gordon and St-Amour (2000), Chetty (2006), and Guiso et al.
+(2018).
+These challenges have motivated a move towards models that incorporate random risk aversion (RRA).
+Within this paradigm, a more economically coherent approach, as advanced by Desmettre and Steffensen (2023), is to aggregate the certainty equivalents over the distribution of the risk aversion parameter.
+While the aggregation in Desmettre and Steffensen (2023) provides a consistent performance measure in monetary units, it introduces a nonlinearity in expectation, which naturally leads to the time inconsistency of the preference.
+The first systematic treatment of time inconsistency was given by Strotz (1955), who proposed the notion of consistent planning, laying the foundation for the intra-personal game-theoretic framework.
+Building on this idea, Ekeland and Lazrak (2006) later developed a rigorous continuous-time formalization, which in turn inspired extensive research on continuous-time control with time-inconsistent preferences.
+For example, Bjork and Murgoci (2010) (published version Björk et al.
+(2017)) established a general theoretical framework by deriving an extended HJB equation and Hu, Jin, and Zhou (2012, 2017) investigated a time-inconsistent stochastic linear–quadratic control problem.
+For further discussions and related developments, see, e.g., Yan and Yong (2019), He and Jiang (2021), Hernández and Possamaï (2023), and the references therein.
+Within the Black-Scholes market model, Desmettre and Steffensen (2023) provide a verification theorem on the intra-personal equilibrium strategies for a general RRA and, in the case of a two-point distributed RRA, characterize the equilibrium by a three-dimensional ODE system without establishing the existence and uniqueness of the solution.
 [^1] This work has spurred further research extending the model to more complex settings.
-For example, chen2025equilibrium incorporate a regime-switching framework in which market dynamics and preferences co-evolve; WANG2025103140 study the management of the DC pension plan, providing semi-explicit solutions for a CRRA utility with a binomial random risk aversion; he2025dynamic establish the existence and uniqueness of the solution to an infinite-dimensional Riccati equation which characterizes the equilibrium strategies under Heston’s SV model for a CRRA or CARA utility function with a general bounded random risk aversion.
-For an investigation on the corresponding pre-commitment problem, see, e.g., xia2024optimal.
-This technical gap is filled by liang2025short.
-For general random risk aversion, liang2025short characterize intra-personal equilibrium strategies using an integral equation.
-They establish the existence and uniqueness of the solution to the integral equation under the assumptions that the variance of the random risk aversion is finite and that the function $h$ defined in (2.6) below is bounded.
-This paper aims to offer more insight into the portfolio selection problem with random risk aversion within the simplest model, in line with Desmettre2023 and liang2025short.
-Our main contributions are summarized as follows:
-  • Complete characterization of all deterministic equilibria.
-We relax the assumptions in the existing literature and derive a full characterization of all deterministic equilibrium strategies.
-An important finding is that there may be infinitely many equilibria when the expectation of the random risk aversion is infinite.
-The uniqueness of the equilibrium strategy for the continuous-time time-inconsistent control problem is a very challenging problem (see, e.g., hjz2021).
+For example, Chen, Guan, and Liang (2025) incorporate a regime-switching framework in which market dynamics and preferences co-evolve; Aquino et al.
+(2025) study a continuous-time portfolio choice problem where state-dependent preferences are determined by an exogenous factor evolving as an Itô diffusion; Wang and Jia (2025) study the management of the DC pension plan, providing semi-explicit solutions for a two-point distributed RRA; He, Jiang, and Xia (2025) establish the existence and uniqueness of the solution to an infinite-dimensional Riccati equation which characterizes the equilibrium strategies under Heston’s SV model for a general bounded RRA.
+For an investigation on the corresponding pre-commitment problem, see, e.g., Xia (2024).
+This technical gap is filled by Liang, Wang, and Xia (2025a).
+For a general RRA, Liang, Wang, and Xia (2025a) characterize intra-personal equilibrium strategies using an integral equation.
+They establish the existence and uniqueness of the solution to the integral equation under the assumptions that the variance of the RRA is finite and that the function $h$ defined in (2.6) below is bounded.
+This paper aims to provide deeper insight into the portfolio selection problem with RRA within the simplest modeling framework, in line with Desmettre and Steffensen (2023) and Liang, Wang, and Xia (2025a).
+Our main contributions are summarized below.
+  • Complete characterization of all deterministic equilibria in closed form.
+We relax the assumptions in the existing literature and derive, in closed form, all deterministic equilibrium strategies.
+An important finding is that there may be infinitely many equilibria when the expectation of the RRA is infinite.
+The uniqueness of the equilibrium strategy for the continuous-time time-inconsistent control problem is a very challenging problem (see, e.g., Hu, Jin, and Zhou (2021, Section 7)).
 This paper sheds light on this issue by demonstrating that equilibrium strategies, even within the class of deterministic strategies, may lack uniqueness.
@@
 We suggest that the agent should maximize the objective functional at the initial time over all available equilibria.
-This leads to the discussion of the so-called optimal equilibria.
-To our best knowledge, there is only one existing paper, wei2024time, on optimal equilibria for a time-inconsistent control problem, and there are three papers, hz2019; hz2020mf and hw2021, on optimal equilibria for time-inconsistent stopping problems.
-We give a necessary and sufficient condition for the existence of optimal equilibria and observe that the optimal equilibrium, if exists, is unique and is also uniformly optimal.
+This leads to a discussion of the so-called optimal equilibria.
+To our best knowledge, there is only one existing paper, Wei, Xia, and Zhao (2026), on optimal equilibria for a time-inconsistent control problem, and there are three papers, Huang and Zhou (2019, 2020) and Huang and Wang (2021), on optimal equilibria for time-inconsistent stopping problems.
+We give a necessary and sufficient condition for the existence of optimal equilibria and observe that the optimal equilibrium, if it exists, is unique and is also uniformly optimal.
 Moreover, if the market price of risk remains non-zero near the terminal time, the optimal equilibrium is also uniformly strictly optimal.
   • Comparative statics on equilibrium strategies.
-We examine how the distribution of the random risk aversion influences the equilibrium strategies.
-It is natural to conjecture that a “larger” random risk aversion would lead to less risky investments.
-We observe by counterexamples that the “largeness” in the sense of the first-order stochastic dominance is insufficient to validate the conjecture.
-However, a stronger stochastic order, the reverse hazard rate dominance, is sufficient.
-Thus we get an analogy with Borell2007 and xia2011risk, where the comparative statics on the optimal portfolio with respect to risk aversion are carried out for the standard expected utility maximization with a general utility function and within the Black-Scholes market model (The corresponding comparative statics for a static model with a risk-free asset and a risky asset goes back to Arrow1970 and Pratt1964:RiskAversion).
-The paper is organized as follows.
-Section 2 presents the model and derives the integral equation characterizing the equilibrium.
-Section 3 analyzes both the finite and infinite expectation cases, providing a complete characterization of all equilibria.
-Section 4 introduces three optimality criteria for equilibrium selection and establishes the necessary and sufficient conditions for their existence and uniqueness.
-Section 5 performs comparative statics to investigate the impact of risk aversion distributions on equilibrium strategies, establishing monotonicity under the reverse hazard rate order, and provides numerical illustrations that verify the theoretical results and present a counterexample under first-order stochastic dominance.
-[^2] Herein, “increasing” means “non-decreasing” and “decreasing” means “non-increasing.”
+We investigate how the distribution of RRA affects the equilibrium strategies.
+Intuitively, one would expect a “comparative property”: a “larger” RRA should lead to less risky investment.
+However, through counterexamples involving two-point distributed RRA, we demonstrate that “largeness” in the sense of first-order stochastic dominance is insufficient to guarantee this property.
+Within the same two-point distribution framework, we further examine the single-crossing property of equilibrium strategies and the monotonicity of the crossing time with respect to (w.r.t.) the probability parameter.
+In contrast, a stronger stochastic order—reverse hazard rate dominance—is shown to be sufficient for the comparative property to hold.
+This yields an analogy with Borell (2007) and Xia (2011), who conducted comparative statics on optimal portfolios w.r.t.
+risk aversion in the context of expected utility maximization with general utility functions under the Black–Scholes market model.
+(The corresponding analysis for a static model with one risk-free asset and one risky asset dates back to Arrow (1970) and Pratt (1964).) Finally, we show that a convex combination of independent and identically distributed (i.i.d.) RRAs results in less risky investment.
+The rest of the paper is organized as follows.
+Section 2 introduces the model and derives the integral equation that characterizes the equilibrium.
+Section 3 analyzes both finite- and infinite-expectation cases, offering closed-form solutions for all equilibria.
+Section 4 resolves equilibrium multiplicity by proposing an optimality-based selection criterion, and establishes necessary and sufficient conditions for the existence, uniqueness, and uniform (strict) optimality of optimal equilibria.
+Section 5 performs a comparative statics of equilibrium strategies w.r.t.
+the distribution and the aggregation of RRA, identifying conditions under which “larger” risk aversion leads to less risky investment.
 2 Problem Formulation
-Let $T>0$ be a finite time horizon and $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ a filtered complete probability space, where $\mathbb{F}=\{\mathcal{F}_{t},\ 0\leq t\leq T\}$ is the augmented natural filtration generated by a $d$-dimensional standard Brownian motion $\{B(t),\ 0\leq t\leq T\}$ and $\mathcal{F}=\mathcal{F}_{T}$.
+Let $T>0$ be a finite time horizon and $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ a filtered complete probability space, where $\mathbb{F}=\{\mathcal{F}_{t}:\ 0\leq t\leq T\}$ is the augmented natural filtration generated by a $d$-dimensional standard Brownian motion $\{B(t):\ 0\leq t\leq T\}$ and $\mathcal{F}=\mathcal{F}_{T}$.
 The market consists of one risk-free asset (bond) and $d$ risky assets (stocks).
 For simplicity, we assume that the interest rate of the bond is zero.
-The dynamics of the stock price processes $S_{i}$, $i=1,\cdots,d$, are given by
+The dynamics of the stock price processes $S_{i}$, $i=1,\cdots,d$, is given by
 $dS_{i}(t)=S_{i}(t)[\mu_{i}(t)dt+\sigma_{i}(t)dB(t)],\quad t\in[0,T],$
@@
 $\int_{0}^{T}|\mu(t)|dt+\sum_{i=1}^{d}\int_{0}^{T}|\sigma_{i}(t)|^{2}dt<\infty$
-and $\sigma(t)$ is invertible for every $t\in[0,T]$.
+and that $\sigma(t)$ is invertible for every $t\in[0,T]$.
 The market price of risk is $\lambda(t)\triangleq(\sigma(t))^{-1}\mu(t)$, $t\in[0,T]$.
 We also assume that $\int_{0}^{T}|\lambda(t)|^{2}\mathrm{d}t<\infty$.
-For any $m\geq 1$ and $\mathbb{S}\subset\mathbb{R}^{m}$, $L^{0}(\mathbb{F},\mathbb{S})$ is the space of $\mathbb{S}$-valued, $\mathbb{F}$-progressively measurable processes.
+For any $m\geq 1$ and $\mathbb{S}\subset\mathbb{R}^{m}$, $L^{0}(\mathbb{F},\mathbb{S})$ is the space consisting of $\mathbb{S}$-valued, $\mathbb{F}$-progressively measurable processes.
 For each $t\in[0,T]$, $p\in[1,\infty]$, $L^{p}(\mathcal{F}_{t},\mathbb{S})$ is the set of all $L^{p}$-integrable, $\mathbb{S}$-valued, and $\mathcal{F}_{t}$-measurable random variables.
-A trading strategy is a process $\pi=\{\pi_{t},\ t\in[0,T)\}\in L^{0}(\mathbb{F},\mathbb{R}^{d})$ such that
+A trading strategy is a process $\pi=\{\pi_{t}:\ t\in[0,T)\}\in L^{0}(\mathbb{F},\mathbb{R}^{d})$ such that
 $\int_{0}^{T}|\pi^{\top}_{t}\mu(t)|\mathrm{d}t+\int_{0}^{T}|\sigma^{\top}(t)\pi_{t}|^{2}\mathrm{d}t<\infty\text{ a.s.,}$
-where $\pi_{t}$ represents the vector of portfolio weights determining the investment of wealth in the stocks at time $t$.
-For any $t\in[0,T)$ and any initial wealth level $w>0$, let $W^{t,w,\pi}=\{W^{t,w,\pi}_{s},\ t\leq s\leq T\}$ denote the self-financing wealth process starting from $w$ at time $t$ and driven by a trading strategy $\pi$.
+where $\pi_{t}$ is the vector of portfolio weights at time $t$.
+For any $t\in[0,T)$ and any initial wealth level $w>0$, let $W^{t,w,\pi}=\{W^{t,w,\pi}_{s}:\ t\leq s\leq T\}$ denote the self-financing wealth process starting from $w$ at time $t$ and driven by a trading strategy $\pi$.
 It evolves according to the stochastic differential equation (SDE):
-$\left\{\begin{aligned} &\mathrm{d}W^{t,w,\pi}_{s}=W^{t,w,\pi}_{s}\pi^{\top}_{s}\mu(s)\mathrm{d}s+W^{t,w,\pi}_{s}\pi^{\top}_{s}\sigma(s)\mathrm{d}B(s),\quad s\geq t,\\ &W^{t,w,\pi}_{t}=w.\end{aligned}\right$.
+$\left\{\begin{aligned} &\mathrm{d}W^{t,w,\pi}_{s}=W^{t,w,\pi}_{s}\pi^{\top}_{s}\mu(s)\mathrm{d}s+W^{t,w,\pi}_{s}\pi^{\top}_{s}\sigma(s)\mathrm{d}B(s),\quad s\in[t,T],\\ &W^{t,w,\pi}_{t}=w.\end{aligned}\right$.
 (2.1)
 Consider the constant relative risk aversion (CRRA) utility functions:
-$u^{\gamma}(x)=\begin{cases}\frac{x^{1-\gamma}}{1-\gamma},&\gamma>0,\gamma\neq 1,x>0,\\ \log x,&\gamma=1,x>0.\end{cases}$
-Under the expected utility theory, for each CRRA utility function $u^{\gamma}$, at time $t$, the certainty equivalent of $W^{t,w,\pi}_{T}$ is $\left(u^{\gamma}\right)^{-1}\left(\mathbb{E}_{t}\left[u^{\gamma}\left(W^{t,w,\pi}_{T}\right)\right]\right)$.
-Following Desmettre2023, we assume that the objective functional of the agent at time $t$ is the weighted average of the certainty equivalents:
+$u^{\gamma}(x)=\begin{cases}\frac{x^{1-\gamma}}{1-\gamma},&\gamma\geq 0,\gamma\neq 1,x>0,\\ \log x,&\gamma=1,x>0.\end{cases}$
+Under the expected utility theory, for each CRRA utility function $u^{\gamma}$, at time $t$, the certainty equivalent of $W^{t,w,\pi}_{T}$ is $\left(u^{\gamma}\right)^{-1}\left(\mathbb{E}_{t}\left[u^{\gamma}\left(W^{t,w,\pi}_{T}\right)\right]\right)$, where $\mathbb{E}_{t}$ denotes the conditional expectation w.r.t.
+$\mathcal{F}_{t}$.
+Following Desmettre and Steffensen (2023), we assume that the objective functional of the agent at time $t$ is the weighted average of the certainty equivalents:
 $J\left(\pi;t,w\right)=\int_{[0,\infty)}\left(u^{\gamma}\right)^{-1}\left(\mathbb{E}_{t}\left[u^{\gamma}\left(W^{t,w,\pi}_{T}\right)\right]\right)\mathrm{d}F_{\bm{R}}(\gamma),$ (2.2)
@@
 A trading strategy $\pi$ is called admissible if for all $t\in[0,T)$ and $w>0$,
-$\begin{cases}\mathbb{E}_{t}\left[\left|u^{\gamma}\left(W^{t,w,\pi}_{T}\right)\right|\right]<\infty\,\ a.s.\quad\text{ for all }\gamma\in\mathrm{supp}(F_{\bm{R}}),\\ \int_{[0,\infty)}\left|\left(u^{\gamma}\right)^{-1}\left(\mathbb{E}_{t}\left[u^{\gamma}\left(W^{t,w,\pi}_{T}\right)\right]\right)\right|\mathrm{d}F_{\bm{R}}(\gamma)<\infty\,\ \ a.s..\end{cases}$
+$\begin{cases}\mathbb{E}_{t}\left[\left|u^{\gamma}\left(W^{t,w,\pi}_{T}\right)\right|\right]<\infty\,\ a.s.\quad\text{ for all }\gamma\in\operatorname*{supp}(F_{\bm{R}}),\\ \displaystyle\int_{[0,\infty)}\left|\left(u^{\gamma}\right)^{-1}\left(\mathbb{E}_{t}\left[u^{\gamma}\left(W^{t,w,\pi}_{T}\right)\right]\right)\right|\mathrm{d}F_{\bm{R}}(\gamma)<\infty\,\ \ a.s.\end{cases}$
 Let $\Pi$ denote the set of all admissible strategies.
@@
 $\pi^{t,\varepsilon,k}_{s}\triangleq\left\{\begin{aligned} &{\bar{\pi}_{s}+k},&s\in[t,t+\varepsilon),\\ &\bar{\pi}_{s},&s\notin[t,t+\varepsilon).\end{aligned}\right$.
-The strategy $\pi^{t,\varepsilon,k}$ serves as a perturbation of $\bar{\pi}$. Following hu2012time; Hu2017 and yan2019time, we introduce the definition of equilibrium strategies as follows.
+The strategy $\pi^{t,\varepsilon,k}$ serves as a perturbation of $\bar{\pi}$.
+Following Hu, Jin, and Zhou (2012, 2017) and Yan and Yong (2019), we introduce the definition of equilibrium strategies as follows.
 Definition 2.1.
-A strategy $\bar{\pi}$ is called an equilibrium strategy if for any $t\in[0,T)$, $w>0$, and $k\in L^{\infty}(\mathcal{F}_{t},\mathbb{R}^{d})$, whenever $\pi^{t,\varepsilon,k}\in\Pi$ for all sufficiently small $\varepsilon>0$, and for any positive sequence $\{\varepsilon_{n},n\geq 1\}$ satisfying $\lim\limits_{n\to\infty}\varepsilon_{n}=0$ we have
-$\limsup_{n\to\infty}\frac{J(\pi^{t,\varepsilon_{n},k};t,w)-J(\bar{\pi};t,w)}{\varepsilon_{n}}\leq 0\quad\text{a.s.}.$ (2.3)
+A strategy $\bar{\pi}$ is called an equilibrium strategy if, for any $t\in[0,T)$, $w>0$, $k\in L^{\infty}(\mathcal{F}_{t},\mathbb{R}^{d})$ with $\pi^{t,\varepsilon,k}\in\Pi$ for all sufficiently small $\varepsilon>0$, and any positive sequence $\{\varepsilon_{n},n\geq 1\}$ with $\lim\limits_{n\to\infty}\varepsilon_{n}=0$, we have
+$\limsup_{n\to\infty}\frac{J(\pi^{t,\varepsilon_{n},k};t,w)-J(\bar{\pi};t,w)}{\varepsilon_{n}}\leq 0\quad\text{a.s.}$ (2.3)
 Remark 2.2.
-In literature, the inequality is usually
-$\limsup_{\varepsilon\downarrow 0}\frac{J(\pi^{t,\varepsilon,k};t,w)-J(\bar{\pi};t,w)}{\varepsilon}\leq 0\quad\text{a.s.}.$
-However, the left-hand side of the inequality might be unmeasurable. To guarantee the measurability, we modify the definition of equilibrium accordingly.
+In literature, the following inequality
+$\limsup_{\varepsilon\downarrow 0}\frac{J(\pi^{t,\varepsilon,k};t,w)-J(\bar{\pi};t,w)}{\varepsilon}\leq 0\quad\text{a.s.}$
+is usually used to define equilibrium strategies.
+However, the left-hand side of the inequality might be unmeasurable.
+To guarantee the measurability, we use the upper limit of a sequence instead of the upper limit of an uncountable net.
 Remark 2.3.
@@
 where
-$J_{0}(t,\pi)\triangleq\int_{[0,\infty)}\left(u^{\gamma}\right)^{-1}\left(\tilde{\mathbb{E}}_{t}\left[u^{\gamma}\left({W^{\pi}_{T}/W^{\pi}_{t}}\right)\right]\right)\mathrm{d}F_{\bm{R}}(\gamma)$
-is the objective functional used in liang2025short and $W^{\pi}$ evolves according to
-$\mathrm{d}W^{\pi}_{t}=W^{\pi}_{t}\pi^{\top}_{t}\mu(t)\mathrm{d}t+W^{\pi}_{t}\pi^{\top}_{t}\sigma(t)\mathrm{d}B(t),\quad W^{\pi}_{0}=w_{0}>0.$
+$J_{0}(t,\pi)\triangleq\int_{[0,\infty)}\left(u^{\gamma}\right)^{-1}\left(\mathbb{E}_{t}\left[u^{\gamma}\left({W^{\pi}_{T}/W^{\pi}_{t}}\right)\right]\right)\mathrm{d}F_{\bm{R}}(\gamma)$
+is the objective functional used in Liang, Wang, and Xia (2025a) and $W^{\pi}$ evolves according to
+$\left\{\begin{aligned} &\mathrm{d}W^{\pi}_{t}=W^{\pi}_{t}\pi^{\top}_{t}\mu(t)\mathrm{d}t+W^{\pi}_{t}\pi^{\top}_{t}\sigma(t)\mathrm{d}B(t),\quad t\in[0,T],\\ &W^{\pi}_{0}=w_{0}>0.\end{aligned}\right$.
 Thus, under Definition 2.1, an equilibrium strategy for $J$ is also an equilibrium strategy for $J_{0}$, and vice versa. We will frequently use $J_{0}$ in the subsequent analysis.
-In our analysis, we follow the approach of liang2025short and focus on the equilibrium strategies given by $\bar{\pi}=(\sigma^{\top})^{-1}a$, where $a$, referred to as the risk exposure vector, is a deterministic and right-continuous funcion in the $L^{2}$ space.
+In our analysis, we follow the approach of Liang, Wang, and Xia (2025a) and focus on the equilibrium strategies given by $\bar{\pi}=(\sigma^{\top})^{-1}a$, where $a$, referred to as the risk exposure vector, is a deterministic and right-continuous function in the $L^{2}$-space.
 Let $\Pi_{d}\subset\Pi$ denote the set of all such strategies.
@@
 The wealth process $W^{t,w,\bar{\pi}}$ of the portfolio $\bar{\pi}$ satisfies the following SDE
-$\mathrm{d}W^{t,w,\bar{\pi}}_{s}=W^{t,w,\bar{\pi}}_{s}[a(s)^{\top}\lambda(s)\mathrm{d}s+a(s)^{\top}\mathrm{d}B(s)],\quad s\geq t,$
+$\left\{\begin{aligned} &\mathrm{d}W^{t,w,\bar{\pi}}_{s}=W^{t,w,\bar{\pi}}_{s}[a(s)^{\top}\lambda(s)\mathrm{d}s+a(s)^{\top}\mathrm{d}B(s)],\quad s\in[t,T],\\ &W^{t,w,\bar{\pi}}_{t}=w,\end{aligned}\right$.
 and $|a|$ is referred to as the risk exposure magnitude (or local volatility).
@@
 $J_{0}(t,\bar{\pi})=\tilde{\mathbb{E}}\left[\mathrm{e}^{-\frac{1}{2}\left(\bm{R}v_{a}(t)-2y_{a}(t)\right)}\right].$ (2.4)
-The following theorem shows that $\pi\equiv\mathbf{0}$ is an equilibrium if $\tilde{\mathbb{E}}[\bm{R}]=\infty$.
-(This result is natural: if the random risk aversion is very large, then the investor puts no money into the risky asset.
-However, in Theorem 3.2(2) we will see that $\pi\equiv\mathbf{0}$ is not necessarily the unique equilibrium.)
+The following theorem shows that $\bar{\pi}\equiv\mathbf{0}$ is an equilibrium if $\tilde{\mathbb{E}}[\bm{R}]=\infty$.
+(This result is natural: if the RRA is very large, then the investor puts no money into the risky asset.
+However, in Theorem 3.2(2) we will see that $\bar{\pi}\equiv\mathbf{0}$ is not necessarily the unique equilibrium.)
 Theorem 2.4.
-$\bar{\pi}\equiv\mathbf{0}$ is an equilibrium if $\tilde{\mathbb{E}}[\bm{R}]=\infty$.
+The risk-free investment ($\bar{\pi}\equiv\mathbf{0}$) is an equilibrium if $\tilde{\mathbb{E}}[\bm{R}]=\infty$.
 Proof.
@@
 $\lim_{\varepsilon\to 0}\frac{\psi(\varepsilon)-1}{\varepsilon}$ $=-\frac{1}{2}|\sigma^{\top}(t)k|^{2}\lim_{\varepsilon\to 0}\int_{[0,\infty)}\gamma\mathrm{e}^{-\frac{\gamma}{2}\int_{t}^{t+\varepsilon}|\sigma^{\top}(s)k|^{2}\mathrm{d}s}\mathrm{d}F_{\bm{R}}(\gamma)$
-$=-\frac{1}{2}|\sigma^{\top}(t)k|^{2}\int_{[0,\infty)}\gamma\mathrm{d}F_{\bm{R}}(\gamma),$
+$=-\frac{1}{2}|\sigma^{\top}(t)k|^{2}\int_{[0,\infty)}\gamma\mathrm{d}F_{\bm{R}}(\gamma)=-\frac{1}{2}|\sigma^{\top}(t)k|^{2}\tilde{\mathbb{E}}[\bm{R}],$
 where the first equality uses the mean value theorem and the second the monotone convergence theorem.
@@
 $\lim_{n\to\infty}\frac{J_{0}(t,\pi^{t,\varepsilon_{n},k})-J_{0}(t,\bar{\pi})}{\varepsilon_{n}}$ $=\lim_{n\to\infty}\frac{\phi(\varepsilon_{n})\psi(\varepsilon_{n})-1}{\varepsilon_{n}}=\lim_{n\to\infty}\frac{(\phi(\varepsilon_{n})-1)\psi(\varepsilon_{n})+\psi(\varepsilon_{n})-1}{\varepsilon_{n}}$
-$=\left(\mu^{\top}(t)k-\frac{1}{2}|\sigma^{\top}(t)k|^{2}\int_{[0,\infty)}\gamma\mathrm{d}F_{\bm{R}}(\gamma)\right).$
+$=\left(\mu^{\top}(t)k-\frac{1}{2}|\sigma^{\top}(t)k|^{2}\tilde{\mathbb{E}}[\bm{R}]\right).$
 Therefore, $\mathbf{0}$ is an equilibrium if $\tilde{\mathbb{E}}[\bm{R}]=\infty$. ∎
@@
 Theorem 2.5.
-$\bar{\pi}=(\sigma^{\top})^{-1}a\in\Pi_{d}$ is an equilibrium if and only if
+A strategy $\bar{\pi}=(\sigma^{\top})^{-1}a\in\Pi_{d}$ is an equilibrium if and only if
 $a(t)=h(v_{a}(t))\lambda(t)$ (2.5)
-holds for any $t\in[0,T)$, where [^3] This function $h$ is slightly different from that in liang2025short: $h(x)$ there amounts to $h(x^{2})$ here.
+holds for any $t\in[0,T)$, where [^3] This function $h$ is slightly different from that one in Liang, Wang, and Xia (2025a): $h(x)$ there amounts to $h(x^{2})$ here.
 Nevertheless, the relevant properties, such as the boundedness of $h$, remain equivalent under the two formulations.
@@
 When $\tilde{\mathbb{E}}[\bm{R}]<\infty$, we have $h(0)={1}/{\tilde{\mathbb{E}}[\bm{R}]}\in(0,\infty)$ and $h\in C([0,\infty))$.
-In this case, an argument similar to the proof of liang2025short yields the equivalence.
+In this case, an argument similar to the proof of Liang, Wang, and Xia (2025a, Theorem 3.4) yields the equivalence.
 When $\tilde{\mathbb{E}}[\bm{R}]=\infty$, we have $h(0)=0$.
-In this case, similar to the case $\tilde{\mathbb{E}}[\bm{R}]<\infty$, (2.3) and (2.5) are equivalent for $t\in[0,T_{0})$, where $T_{0}=\inf\{t\in[0,T]:v_{a}(t)=0\}$.
+In this case, similar to the case $\tilde{\mathbb{E}}[\bm{R}]<\infty$, (2.3) and (2.5) are equivalent for $t\in[0,T_{0})$, where $T_{0}\triangleq\inf\{t\in[0,T]:\ v_{a}(t)=0\}$.
 Moreover, they are automatically satisfied for $t\in[T_{0},T)$ based on Theorem 2.4 and $h(0)=0$.
@@
 $l(y)\triangleq\tilde{\mathbb{E}}[e^{-\bm{R}y}],\quad y\geq 0.$
-The function $h$ defined in (2.6) plays a central role in our analysis. Here, we present two examples for which $h$ can be explicitly worked out.
+Its derivative $l^{\prime}(0)$ exists and is finite for all $y>0$.
+In the case $\tilde{\mathbb{E}}[\bm{R}]<\infty$, we have $l^{\prime}(0)=-\tilde{\mathbb{E}}[\bm{R}]$.
+In the case $\tilde{\mathbb{E}}[\bm{R}]=\infty$, we extend the definition by setting $l^{\prime}(0)=$-$\infty$.
+Then, for any $x\geq 0$, the function $h$ is related to $l$ through $h(x)=-{l(x/2)}/{l^{\prime}(x/2)}$.
+The function $h$ defined in (2.6) plays a central role in our analysis.
+Here, we present two examples for which $h$ can be explicitly worked out.
 Example 2.6.
-Consider the following two distributions for the random risk aversion $\bm{R}$.
+Consider the following two distributions for the RRA $\bm{R}$.
   1. (1)
@@
     $h(x)=-\frac{l(\frac{x}{2})}{l^{\prime}(\frac{x}{2})}=\frac{1}{\alpha\beta}\left(1+\frac{\beta}{2}x\right),\quad x\geq 0.$
-In the two cases of the above example, the functions $h$ are unbounded, violating the boundedness assumption on $h$ made by liang2025short.
+In the two cases of the above example, the functions $h$ are unbounded, violating the boundedness assumption on $h$ made by Liang, Wang, and Xia (2025a).
 Moreover, as we will see in Example 3.5 below, the equilibrium strategies can be explicitly obtained for both cases.
@@
 We first transform Condition (2.5) into an ordinary differential equation (ODE) for $v_{a}$.
-Differentiating $v_{a}(t)=\int_{t}^{T}|a(s)|^{2}ds$ with respect to $t$, we get from (2.5) that $v_{a}$ satisfies the following ODE:
+Differentiating $v_{a}(t)=\int_{t}^{T}|a(s)|^{2}\mathrm{d}s$ w.r.t.
+$t$, we get from (2.5) that $v_{a}$ satisfies the following ODE:
 $\begin{cases}v^{\prime}(t)=-h^{2}(v(t))|\lambda(t)|^{2},&t\in\left[0,T\right),\\ v(T)=0.\end{cases}$ (3.1)
-Conversely, assume that $v$ solves the ODE (3.1) and let $a(\cdot)=h(v(\cdot))\lambda(\cdot)$. Then $a$ satisfies (2.5).
-The ODE (3.1) is an equation with separated variables. We can obtain its solutions in closed form.
+Conversely, assume that $v$ solves ODE (3.1) and let $a(\cdot)\triangleq h(v(\cdot))\lambda(\cdot)$. Then $a$ satisfies (2.5).
+ODE (3.1) is an equation with separated variables. We can obtain its solutions in closed form.
 Now we introduce two functions. The first function $\Lambda$ is defined by
@@
 The second function $H$ is defined by
-$H(y)\triangleq\int_{0}^{y}\frac{1}{h^{2}(x)}\mathrm{d}x=\int_{0}^{y}\left({\tilde{\mathbb{E}}\left[\bm{R}\mathrm{e}^{-\frac{1}{2}\bm{R}x}\right]}\left/{\tilde{\mathbb{E}}\left[\mathrm{e}^{-\frac{1}{2}\bm{R}x}\right]}\right.\right)^{2}\mathrm{d}x,\quad y\in[0,\infty].$
+$\displaystyle H(y)\triangleq\int_{0}^{y}\frac{1}{h^{2}(x)}\mathrm{d}x=\int_{0}^{y}\left({\tilde{\mathbb{E}}\left[\bm{R}\mathrm{e}^{-\frac{1}{2}\bm{R}x}\right]}\left/{\tilde{\mathbb{E}}\left[\mathrm{e}^{-\frac{1}{2}\bm{R}x}\right]}\right.\right)^{2}\mathrm{d}x,\quad y\in[0,\infty].$ (3.2)
 Occasionally, we write $H_{\bm{R}}$ for $H$ to emphasize its dependence on $\bm{R}$.
-If $\Lambda(0)=0$, then $\lambda(\cdot)\equiv 0$ on $[0,T)$ and, in view of (2.5), there is a unique equilibrium $\pi\equiv\mathbf{0}$.
+If $\Lambda(0)=0$, then $\lambda(\cdot)\equiv\mathbf{0}$ on $[0,T)$ and, in view of (2.5), there is a unique equilibrium $\bar{\pi}\equiv\mathbf{0}$.
 This trivial case is excluded by making the following standing assumption.
@@
 $\Lambda(0)\in(0,\infty)$.
-The next theorem provides, in closed form, the solutions to the ODE (3.1) and the equilibrium strategies in $\Pi_{d}$ .
+The next theorem provides, in closed form, the solutions to ODE (3.1) and the equilibrium strategies in $\Pi_{d}$. We will use $\Pi_{d}^{e}$ to denote all equilibrium strategies in $\Pi_{d}$, that is,
+$\Pi_{d}^{e}\triangleq\{\pi\in\Pi_{d}\,:\ \pi\text{ is an equilibrium strategy}\}.$
 Theorem 3.2.
@@
   2. (2)
-    If $\tilde{\mathbb{E}}[\bm{R}]=\infty$ and $H(\varepsilon)<\infty$ for all $\varepsilon>0$, the ODE (3.1) admits a one-parameter family of solutions, indexed by $T_{0}\in\mathcal{T}$:
+    If $\tilde{\mathbb{E}}[\bm{R}]=\infty$ and $H(\varepsilon)<\infty$ for all $\varepsilon>0$, then ODE (3.1) admits a one-parameter family of solutions, indexed by $T_{0}\in\mathcal{T}$:
     $v^{(T_{0})}(t)\triangleq\begin{cases}H^{-1}\left(\int_{t}^{T_{0}}|\lambda(s)|^{2}\mathrm{d}s\right),&0\leq t<T_{0},\\ 0,&T_{0}\leq t\leq T,\end{cases}$ (3.3)
     where
-    $\mathcal{T}\triangleq\left\{T_{0}\in[0,T]\;\bigg|\;H(\infty)>\Lambda(0)-\Lambda(T_{0})\text{ and }\Lambda(t)>\Lambda(T_{0})\text{ for all }t\in[0,T_{0})\right\}.$
-    Then the set of all equilibrium strategies in $\Pi_{d}$ is :
+    $\mathcal{T}\triangleq\left\{T_{0}\in[0,T]:H(\infty)>\Lambda(0)-\Lambda(T_{0})\text{ and }\Lambda(t)>\Lambda(T_{0})\text{ for all }t\in[0,T_{0})\right\}.$
+    In this case,
     $\Pi^{e}_{d}=\left\{\pi^{(T_{0})}(\cdot)\triangleq(\sigma^{\top}(\cdot))^{-1}h(v^{(T_{0})}(\cdot))\lambda(\cdot):\ T_{0}\in\mathcal{T}\right\}.$
@@
 Proof.
-Assume $\tilde{\mathbb{E}}[\bm{R}]<\infty$. In this case, $h(x)>0$ for all $x\geq 0$. Suppose that $v(\cdot)$ is a solution to the ODE (3.1). Separating variables in (3.1) yields
-$-|\lambda(t)|^{2}=\frac{v’(t)}{h^{2}(v(t))},\quad t\in[0,T).$ (3.3)
+Assume $\tilde{\mathbb{E}}[\bm{R}]<\infty$. In this case, $h(x)>0$ for all $x\geq 0$. Suppose that $v(\cdot)$ is a solution to ODE (3.1). Separating variables in (3.1) yields
+$-|\lambda(t)|^{2}=\frac{v^{\prime}(t)}{h^{2}(v(t))},\quad t\in[0,T).$ (3.4)
 Integrating (3.4) from $t$ to $T$ yields
-$-\Lambda(t)=\int_{t}^{T}\frac{v’(s)}{h^{2}(v(s))}\mathrm{d}s=\int_{v(t)}^{0}\frac{\mathrm{d}y}{h^{2}(y)}=-\int_{0}^{v(t)}\frac{\mathrm{d}y}{h^{2}(y)}=-H(v(t)),$
+$-\Lambda(t)=\int_{t}^{T}\frac{v^{\prime}(s)}{h^{2}(v(s))}\mathrm{d}s=\int_{v(t)}^{0}\frac{\mathrm{d}y}{h^{2}(y)}=-\int_{0}^{v(t)}\frac{\mathrm{d}y}{h^{2}(y)}=-H(v(t)),$
 where the second equality follows from the change of variables $y=v(s)$.
 Thus, $H(v(t))=\Lambda(t)$ for all $t\in[0,T]$.
-As $h(0)={1}/{\tilde{\mathbb{E}}[\bm{R}]}>0$, it follows that $H(0)=0$ and $H(\cdot)\in{C}^{1}\left([0,+\infty)\right)$ is strictly increasing.
+Because $h(0)={1}/{\tilde{\mathbb{E}}[\bm{R}]}>0$, it follows that $H(0)=0$ and $H(\cdot)\in{C}^{1}\left([0,\infty)\right)$ is strictly increasing.
 Thus, the inverse function $H^{-1}$ is well-defined, continuous, and strictly increasing in $\left[0,H(\infty)\right)$.
 Thus, $H(\infty)>H(v(0))=\Lambda(0)$ and $v(t)=H^{-1}(\Lambda(t))$ for all $t\in[0,T]$, which yields the uniqueness of the solution.
-Conversely, if $H(\infty)>\Lambda(0)$, it is straightforward to verify that $v(\cdot)=H^{-1}(\Lambda(\cdot))$ indeed solves uniquely the ODE.
+Conversely, if $H(\infty)>\Lambda(0)$, it is straightforward to verify that $v(\cdot)\triangleq H^{-1}(\Lambda(\cdot))$ indeed solves uniquely the ODE.
 Assume that $\tilde{\mathbb{E}}[\bm{R}]=\infty$ and $H(\varepsilon)<\infty$ for all $\varepsilon>0$.
-We can see that $H(0)=0$, $H\in C^{0}([0,+\infty])\cap C^{1}((0,+\infty))$, and $H$ is strictly increasing.
-Suppose that $v(\cdot)$ is a solution to the ODE (3.1).
-Let $T_{0}=\inf\{t\in[0,T]:v(t)=0\}$.
+We can see that $H(0)=0$, $H\in C^{0}([0,\infty])\cap C^{1}((0,\infty))$, and $H$ is strictly increasing.
+Suppose that $v(\cdot)$ is a solution to ODE (3.1).
+Let $T_{0}\triangleq\inf\{t\in[0,T]:\ v(t)=0\}$.
 Then $v(t)=0$ for any $t\in[T_{0},T]$ and $v(t)>0$ for any $t\in[0,T_{0})$.
@@
 By continuity, we have $v(t^{\prime})=0$ and $v(s)>0$ for any $s\in[t,t^{\prime})$.
-Integrating the ODE (3.3) from $t$ to $t^{\prime}$ yields
+Integrating ODE (3.4) from $t$ to $t^{\prime}$ yields
 $\Lambda(t^{\prime})-\Lambda(t)=\int_{t}^{t^{\prime}}-|\lambda(s)|^{2}\mathrm{d}s=\int_{v(t)}^{0}\frac{\mathrm{d}y}{h^{2}(y)}=-\infty,$
-which is impossible. Therefore, the ODE (3.1) admits the unique solution $v\equiv 0$. ∎
-Assertion (2) of Theorem 3.2 characterizes the solutions to the ODE (3.1) via the parameter $T_{0}\in\mathcal{T}$.
+which is impossible. Therefore, ODE (3.1) admits the unique solution $v\equiv 0$. ∎
+Assertion (2) of Theorem 3.2 provides, in closed form, the solutions to ODE (3.1) via the parameter $T_{0}\in\mathcal{T}$.
 The following lemma provides an alternative explicit characterization of the set $\mathcal{T}$ via an auxiliary mapping.
@@
 Let $\mathcal{A}\triangleq[0,\Lambda(0)]\cap(\Lambda(0)-H(\infty),\Lambda(0)]$ and
-$\quad\varphi(\eta)\triangleq\min\{t\in[0,T]\mid\Lambda(t)=\eta\},\quad\eta\in[0,\Lambda(0)].$
-Then we have $\mathcal{T}=\{\varphi(\eta)\mid\eta\in\mathcal{A}\}$.
+$\quad\varphi(\eta)\triangleq\min\{t\in[0,T]:\,\Lambda(t)=\eta\},\quad\eta\in[0,\Lambda(0)].$
+Then we have $\mathcal{T}=\{\varphi(\eta):\,\eta\in\mathcal{A}\}$.
 Proof.
@@
 $T_{0}=\inf\{t\in[0,T]:\Lambda(t)=\Lambda(T_{0})\}=\varphi(\Lambda(T_{0})).$
-Thus, $\mathcal{T}\subset\{\varphi(\eta)\mid\eta\in\mathcal{A}\}$.
+Thus, $\mathcal{T}\subset\{\varphi(\eta):\,\eta\in\mathcal{A}\}$.
 Conversely, for any $\eta\in\mathcal{A}$, the continuity of $\Lambda$ implies $\Lambda(\varphi(\eta))=\eta\in\mathcal{A}$.
 Moreover, for any $t\in[0,\varphi(\eta))$, we have $\Lambda(t)>\Lambda(\varphi(\eta))$ by the definition of $\varphi$.
-Therefore, $\varphi(\eta)\in\mathcal{T}$ and hence $\{\varphi(\eta)\mid\eta\in\mathcal{A}\}\subset\mathcal{T}$.
+Therefore, $\varphi(\eta)\in\mathcal{T}$ and hence $\{\varphi(\eta):\,\eta\in\mathcal{A}\}\subset\mathcal{T}$.
 ∎
 Remark 3.4.
-Because $\varphi$ is strictly decreasing, the sets $\mathcal{T}$ and $\mathcal{A}$ are in one-to-one correspondence.
+The sets $\mathcal{T}$ and $\mathcal{A}$ are in one-to-one correspondence because $\varphi$ is strictly decreasing.
 It is easily seen that $\mathcal{A}$ contains infinitely many elements.
 Consequently, under the conditions $\tilde{\mathbb{E}}[\bm{R}]=\infty$ and $H(\varepsilon)<\infty$ for all $\varepsilon>0$, we obtain infinitely many equilibria.
-It should be noted that the existing studies, such as Desmettre2023 and liang2025short, focus mainly on the specific cases covered by Theorem 3.2(1), under the additional assumptions that $h(\cdot)$ is bounded and $\bm{R}$ is square-integrable.
-Our result covers any general random risk aversion $\bm{R}$.
+It should be noted that the existing studies, such as Desmettre and Steffensen (2023) and Liang, Wang, and Xia (2025a), focus mainly on the specific cases covered by Theorem 3.2(1), under the additional assumptions that $h(\cdot)$ is bounded and $\bm{R}$ is square-integrable.
+Our result covers any general RRA $\bm{R}$.
 In the following two examples, the functions $h$ are unbounded and the equilibrium strategies are explicitly worked out.
@@
 If $\Lambda(0)<\theta^{2}$, then $v(t)=H^{-1}(\Lambda(t))=-\ln\left(1-\dfrac{\Lambda(t)}{\theta^{2}}\right)$.
-In this case, the uinque equilibrium strategy is explicitly given by
-    $\bar{\pi}(t)=(\sigma^{\top})(t)^{-1}a(t)=(\sigma^{\top}(t))^{-1}h(v(t))\lambda(t)=\dfrac{(\sigma^{\top}(t))^{-1}\lambda(t)}{\sqrt{\theta^{2}-\Lambda(t)}},\quad t\in[0,T).$
+In this case, the unique equilibrium strategy is explicitly given by
+    $\bar{\pi}(t)=(\sigma^{\top}(t))^{-1}a(t)=(\sigma^{\top}(t))^{-1}h(v(t))\lambda(t)=\dfrac{(\sigma^{\top}(t))^{-1}\lambda(t)}{\sqrt{\theta^{2}-\Lambda(t)}},\quad t\in[0,T).$
   2. (2)
@@
 This example is provided for Assertions (2) and (3) in Theorem 3.2.
-Let $\bm{R}_{\alpha}\sim G_{\alpha}$, where $\alpha\in(0,1)$ is a fixed parameter and $G_{\alpha}$ denotes the distribution function whose Laplace transform satisfies [^4] See Pollard1946 and Feller1971.
+Let $\bm{R}_{\alpha}\sim G_{\alpha}$, where $\alpha\in(0,1)$ is a fixed parameter and $G_{\alpha}$ denotes the distribution function whose Laplace transform satisfies (see Pollard (1946) and Feller (1971, Section XIII.6))
 $\int_{0}^{\infty}\mathrm{e}^{-vx}\mathrm{d}G_{\alpha}(x)=\mathrm{e}^{-v^{\alpha}},\quad v\geq 0.$
@@
 Obviously, $h$ is unbounded.
-For simplicity, we assume that $d=1$ and $\sigma(\cdot)\equiv\sigma\neq 0$ and $\lambda(\cdot)\equiv\lambda\neq 0$ are constants in $\mathbb{R}$.
-  • Let $\alpha\in(0.5,1)$. By direct calculation, the solutions to the ODE (3.1) take the form
+For simplicity, we assume that $d=1$, $\sigma(\cdot)\equiv\sigma\neq 0$ and $\lambda(\cdot)\equiv\lambda\neq 0$ are constants in $\mathbb{R}$.
+  • Let $\alpha\in(0.5,1)$. By direct calculation, the solutions to ODE (3.1) take the form
     $v^{(T_{0})}(t)=\begin{cases}\left(2^{2\alpha-2}\dfrac{1}{\alpha^{2}}(2\alpha-1)\lambda^{2}(T_{0}-t)\right)^{\frac{1}{2\alpha-1}},&t\in[0,T_{0}),\\ 0,&t\in[T_{0},T],\end{cases}$
@@
     where $T_{0}\in[0,T]$.
-  • Let $\alpha\in(0,0.5]$. Direct calculation shows that the ODE (3.1) admits only the trivial solution $v\equiv 0$ and therefore $\pi=0$ is the unique equilibrium.
+  • Let $\alpha\in(0,0.5]$. Direct calculation shows that ODE (3.1) admits only the trivial solution $v\equiv 0$ and therefore $\bar{\pi}=0$ is the unique equilibrium.
 Assume $\tilde{\mathbb{E}}[\bm{R}]<\infty$ and $H(\infty)>\Lambda(0)$.
-Assertion (1) of Theorem 3.2 shows that the ODE (3.1) admits a unique solution $v$ and therefore there exists a unique function $a$ such that (2.5) holds for all $t\in[0,T)$.
+Assertion (1) of Theorem 3.2 shows that ODE (3.1) admits a unique solution $v$ and therefore the deterministic equilibrium strategy exists and is unique.
 To emphasize the dependence on $\bm{R}$, we occasionally write $v_{\bm{R}}$ and $a_{\bm{R}}$ for $v$ and $a$, respectively.
 Remark 3.7.
-The explicit characterization of the equilibrium strategy in Theorem 3.2(1) yields a clear structural decomposition of the risk exposure vector, given by
+The closed form of equilibrium strategy in Theorem 3.2(1) yields a clear structural decomposition of the risk exposure vector $a$, given by
 $a(t)=h(H^{-1}(\Lambda(t)))\lambda(t),\quad t\in[0,T).$ (3.5)
-Note that the definitions of the functions $h(\cdot)$ and $H(\cdot)$ depend solely on the distribution of the random risk aversion $\bm{R}$}.
-The representation (3.4) shows that, at time $t$, the effects of three distinct factors on the risk exposure vector are cleanly separated: the instantaneous market opportunity $\lambda(t)$, the aggregate future market opportunities $\Lambda(t)$ over the remaining interval $(t,T)$, and the random risk aversion preference itself.
+Note that the functions $h(\cdot)$ and $H(\cdot)$ depend solely on the distribution of the RRA $\bm{R}$}.
+The representation (3.5) shows that, at time $t$, the effects of three distinct factors on the risk exposure vector are totally separated: the instantaneous market opportunity $\lambda(t)$, the aggregate future market opportunities $\Lambda(t)$ over the remaining interval $(t,T)$, and the RRA itself (represented by $h$ and $H$).
 More specifically, at time $t$, if $\Lambda(t)$ and $\bm{R}$} are fixed, a larger current risk premium $|\lambda(t)|$ leads to a larger $|a(t)|$.
-If $|\lambda(t)|$ and $\bm{R}$} are fixed, then a larger future market opportunity $\Lambda(t)$ implies a larger value of $H^{-1}(\Lambda(t))$, and hence a larger value of $h\big(H^{-1}(\Lambda(t))\big)$ (see Proposition 3.8 for a detailed discussion; see alos Lemma A.1 for the monotonicity of $h$).
+If $|\lambda(t)|$ and $\bm{R}$} are fixed, then a larger future market opportunity $\Lambda(t)$ implies a larger value of $H^{-1}(\Lambda(t))$, and hence a larger value of $h\big(H^{-1}(\Lambda(t))\big)$ (see Proposition 3.8 for a detailed discussion; see also Lemma A.1 for the monotonicity of $h$).
 Finally, if both $\lambda(t)$ and $\Lambda(t)$ are fixed, one can study how $\bm{R}$} affects $h$ and $H$, and thereby its impact on $a(t)$; see the comparative statics analysis in Section 5.
-We emphasize that, in contrast to the classical CRRA utility, which generates a “myopic” strategy reacting primarily to the instantaneous market price of risk $\lambda(t)$, the equilibrium under random risk aversion is inherently forward-looking due to the dependence on $\Lambda(t)$.
-Finally, we refer to liang2024dynamic and liang2025dynamic for similar structural representations, although monotonicity properties and comparative statics are not investigated therein.
-Next, we study the impact of changes in future market opportunities on the risk exposure vector.
+We emphasize that, in contrast to the classical CRRA utility, which generates a “myopic” strategy reacting primarily to the instantaneous market price of risk $\lambda(t)$, the equilibrium under RRA is inherently forward-looking due to the dependence on $\Lambda(t)$.
+Finally, we refer to Liang et al.
+(2024) and Liang, Xia, and Yuan (2025b) for similar structures of the equilibrium strategies under other preferences, although monotonicity properties and comparative statics are not investigated therein.
+Next, we study the effect of changes in future market opportunities on the risk exposure vector.
 Proposition 3.8.
-Let $\bm{R}$ be a non-negative random variable with $\tilde{\mathbb{E}}[\bm{R}]\in(0,+\infty)$.
+Let $\bm{R}$ be a non-negative random variable with $\tilde{\mathbb{E}}[\bm{R}]\in(0,\infty)$.
 Fix a vector $\lambda_{0}\in\mathbb{R}^{d}$.
-Consider a sequence of market settings indexed by $n\in\mathbb{N}$, defined by the pairs $\{(\lambda_{n}(\cdot),T_{n})\}_{n\geq 1}$, where $T_{n}>0$ and $\lambda_{n}:[0,T_{n}]\to\mathbb{R}^{d}$ satisfying $\lambda_{n}(0)=\lambda_{0}$.
+Consider a sequence of market settings indexed by $n\in\mathbb{N}$, defined by the pairs $\{(\lambda_{n}(\cdot),T_{n})\}_{n\geq 1}$, where $T_{n}>0$ and $\lambda_{n}:[0,T_{n}]\to\mathbb{R}^{d}$ satisfy $\lambda_{n}(0)=\lambda_{0}$.
 Let $\Lambda_{n}\triangleq\int_{0}^{T_{n}}|\lambda_{n}(s)|^{2}ds$.
@@
   1. (1)
-    If the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly increasing and $\Lambda_{n}\nearrow H(\infty)$, then
-    $\lim_{n\to\infty}a_{n}(0)=\frac{\lambda_{0}}{r_{0}}\quad\text{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(if $r_{0}>0$)},$
-    where $r_{0}\triangleq\operatorname{essinf}\bm{R}$. Moreover, $|a_{n}(0)|$ converges monotonically increasing, i.e.,
-    $|a_{n}(0)|\nearrow\frac{|\lambda_{0}|}{r_{0}}.$
+    If the sequence $\{\Lambda_{n}\}_{n\geq 1}$ strictly increases to $H(\infty)$, then $\lim\limits_{n\to\infty}a_{n}(0)=\frac{\lambda_{0}}{r_{0}}$, where $r_{0}\triangleq\operatorname{essinf}\bm{R}$.
+Moreover, $|a_{n}(0)|\nearrow\frac{|\lambda_{0}|}{r_{0}}$ as $n\to\infty$.
   2. (2)
-    If the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly decreasing and $\Lambda_{n}\searrow 0$, then
-    $\lim_{n\to\infty}a_{n}(0)=\frac{\lambda_{0}}{\tilde{\mathbb{E}}[\bm{R}]}.$
-    Moreover, $|a_{n}(0)|$ converges monotonically decreasing, i.e.,
-    $|a_{n}(0)|\searrow\frac{|\lambda_{0}|}{\tilde{\mathbb{E}}[\bm{R}]}.$
-Here, we adopt the convention that $\frac{1}{0}=\infty$.
+    If the sequence $\{\Lambda_{n}\}_{n\geq 1}$ strictly decreases to 0, then $\lim\limits_{n\to\infty}a_{n}(0)=\frac{\lambda_{0}}{\tilde{\mathbb{E}}[\bm{R}]}$.
+Moreover, $|a_{n}(0)|\searrow\frac{|\lambda_{0}|}{\tilde{\mathbb{E}}[\bm{R}]}$ as $n\to\infty$.
+Here, we adopt the conventions $\frac{1}{0}=\infty$ and $0\cdot\infty=0$.
 Proof.
@@
 The convergence results in Proposition 3.8 depend solely on the sequence of integral values $\{\Lambda_{n}\}_{n\geq 1}$, and no convergence of the parameter sequence $\{(\lambda_{n}(\cdot),T_{n})\}_{n\geq 1}$ itself is required.
-Moreover, focusing on time 0 entails no loss of generality.
-For any arbitrary time $t$, analogous conclusions follow by a simple time translation, with $t$ treated as the initial time.
+Moreover, focusing on time 0 does not cause a loss of generality.
+For any time $t$, analogous conclusions follow by a simple time translation, with $t$ treated as the initial time.
 Remark 3.10.
-In the single-stock case, assuming that $\lambda>0$ and $\sigma>0$ are constants, one can use $\hat{\pi}(\cdot)=\frac{1}{\sigma}h(v(\cdot))\lambda$, together with the monotonicity of $h$ given in Lemma A.1, to establish a favorable property: $\pi(\cdot)$ is decreasing.
-See Section 4.3 of Desmettre2023 for a detailed discussion.
-However, the results in Desmettre2023 do not yield Proposition 3.8.
-The proof of Proposition 3.8 crucially relies on our explicit characterization of the risk exposure vector, given in (3.4), which is not available in their framework.
-As the final part of this section, we use the explicit characterization of the equilibria in Theorem 3.2 to investigate the continuity properties of $a(\cdot)$ with respect to the distribution of the random risk aversion.
-More specifically, the following theorem shows that if a sequence of random risk aversion $\{R_{n}\}_{n\geq 1}$ converges in distribution to $R$, then the corresponding sequence $\{a_{n}\}_{n\geq 1}$ of volatility converges to $a$.
+In the single-stock case, assuming that $\lambda>0$ and $\sigma>0$ are constants, one can use $\widehat{\pi}(\cdot)=\frac{1}{\sigma}h(v(\cdot))\lambda$, together with the monotonicity of $h$ given in Lemma A.1, to establish a favorable property: $\pi(\cdot)$ is decreasing.
+See Section 4.3 of Desmettre and Steffensen (2023) for a detailed discussion.
+However, the results in Desmettre and Steffensen (2023) do not yield Proposition 3.8.
+The proof of Proposition 3.8 is crucially based on our closed-form expression for the risk exposure vector, given in (3.5), which is not available in their work.
+As the final part of this section, we use the closed form of the equilibria in Theorem 3.2 to investigate the continuity properties of $a(\cdot)$ w.r.t.
+the distribution of the random risk aversion.
+More specifically, the following theorem shows that if a sequence of RRA $\{\bm{R}_{n}\}_{n\geq 1}$ converges in distribution to $\bm{R}$, then the corresponding sequence of risk exposure vectors $\{a_{n}\}_{n\geq 1}$ converges to $a$.
 Theorem 3.11.
@@
 Then $\lim\limits_{n\to\infty}a_{n}(t)=a(t)$ for all $t\in[0,T)$.
-Moreover, if $\lambda\in L^{\infty}([0,T])$, then the convergence is uniform on $[0,T)$.
+Moreover, if $\lambda\in L^{\infty}([0,T))$, then the convergence is uniform on $[0,T)$.
 Proof.
@@
 Definition 4.1.
-An equilibrium $\hat{\pi}\in\Pi^{e}_{d}$ is called optimal if $J(\hat{\pi};0,w)\geq J(\pi;0,w)$ for all $\pi\in\Pi^{e}_{d}$ and for all $w>0$.
-Recalling the relationship $J(\pi;t,w)=wJ_{0}(t,\pi)$ in Remark 2.3, finding an optimal equilibrium in $\Pi^{e}_{d}$ is equivalent to finding a $\hat{\pi}\in\Pi^{e}_{d}$ that maximizes $J_{0}(t,\pi)$.
+An equilibrium $\widehat{\pi}\in\Pi^{e}_{d}$ is called optimal if $J(\widehat{\pi};0,w)\geq J(\pi;0,w)$ for all $\pi\in\Pi^{e}_{d}$ and for all $w>0$.
+Recalling the relationship $J(\pi;t,w)=wJ_{0}(t,\pi)$ in Remark 2.3, finding an optimal equilibrium in $\Pi^{e}_{d}$ is equivalent to finding a $\widehat{\pi}\in\Pi^{e}_{d}$ that maximizes $J_{0}(t,\pi)$.
 Let $\pi\in\Pi^{e}_{d}$.
@@
 where the third equality arises from the change of variable $y=v^{(T_{0})}(s)$. Thus, we obtain
-$J_{0}(t,\pi^{(T_{0})})=\exp\left\{\int_{0}^{v^{(T_{0})}(t)}\frac{1}{h(y)}\mathrm{d}y\right\}l\left(\frac{v^{(T_{0})}(t)}{2}\right).$
+$J_{0}(t,\pi^{(T_{0})})=\exp\left(\int_{0}^{v^{(T_{0})}(t)}\frac{1}{h(y)}\mathrm{d}y\right)l\left(\frac{v^{(T_{0})}(t)}{2}\right).$
 Let $\mathcal{L}(z):[0,\infty)\to\mathbb{R}$ be defined by $\mathcal{L}(z)=\int_{0}^{z}\frac{1}{h(y)}\mathrm{d}y+\log l\left(\frac{z}{2}\right).$ Then we have
 $J_{0}(t,\pi^{(T_{0})})=\exp\left(\mathcal{L}(v^{(T_{0})}(t))\right).$ (4.1)
-According to (4.1), in order to study the maximization of $J_{0}(t,\pi^{(T_{0})})$ over $T_{0}\in\mathcal{T}$, we need to analyze the monotonicity of $\mathcal{L}$ and the monotonicity of $v^{(T_{0})}$ with respect to $T_{0}$.
-These properties are characterized in Lemma 4.2 and Lemma 4.3, respectively.
+According to (4.1), in order to study the maximization of $J_{0}(t,\pi^{(T_{0})})$ over $T_{0}\in\mathcal{T}$, we need to analyze the monotonicity of $\mathcal{L}$ and the monotonicity of $v^{(T_{0})}$ w.r.t.
+$T_{0}$.
+These properties are provided in Lemma 4.2 and Lemma 4.3, respectively.
 Lemma 4.2.
@@
 By direct differentiation, we have
-$\frac{d\mathcal{L}}{dz}=\frac{1}{h(z)}-\frac{1}{2}\frac{\tilde{\mathbb{E}}\left[\bm{R}e^{-\frac{1}{2}\bm{R}z}\right]}{\tilde{\mathbb{E}}\left[e^{-\frac{1}{2}\bm{R}z}\right]}=\frac{1}{2h(z)}>0\quad\text{for all }z>0.$
+$\frac{d\mathcal{L}}{dz}(z)=\frac{1}{h(z)}-\frac{1}{2}\frac{\tilde{\mathbb{E}}\left[\bm{R}e^{-\frac{1}{2}\bm{R}z}\right]}{\tilde{\mathbb{E}}\left[e^{-\frac{1}{2}\bm{R}z}\right]}=\frac{1}{2h(z)}>0\quad\text{for all }z>0.$
 As such, $\mathcal{L}(z)$ is strictly increasing in $z$. ∎
@@
 Let $T_{0,1},T_{0,2}\in\mathcal{T}$ with $T_{0,1}<T_{0,2}$.
-For $t\in[T_{0,2},T]$, we have $v^{(T_{0,1})}(t)=v^{(T_{0,2})}(t)=0$ based on Definition of $\mathcal{T}$.
+For $t\in[T_{0,2},T]$, we have $v^{(T_{0,1})}(t)=v^{(T_{0,2})}(t)=0$ based on the definition of $\mathcal{T}$.
 For $t\in[T_{0,1},T_{0,2})$, we have $v^{(T_{0,2})}(t)>0=v^{(T_{0,1})}(t)$.
@@
 By the definition of $\mathcal{T}$, $\Lambda(T_{0,1})>\Lambda(T_{0,2})$.
-Because $H^{-1}$ is strictly increasing, it follows that $v^{(T_{0,1})}(t)<v^{(T_{0,2})}(t)$ for $t\in[0,T_{0,1})$.
+Because $H^{-1}$ is strictly increasing, we have $v^{(T_{0,1})}(t)<v^{(T_{0,2})}(t)$ for $t\in[0,T_{0,1})$.
 Thus, (4.2) is proved.
@@
 Theorem 4.4.
-Under the conditions of Assertion (2) of Theorem 3.2, we have the followings:
+Under the conditions of Assertion (2) of Theorem 3.2, we have the following assertions.
   1. (1)
@@
 Thus, finding the optimal equilibrium strategy is equivalent to identifying the maximal parameter $T_{0}\in\mathcal{T}$.
-By Lemma 3.3, $\mathcal{T}=\{\varphi(\eta)\mid\eta\in\mathcal{A}\}$.
-Because $\varphi$ is strictly decreasing on $\mathcal{A}$, the existence of a maximal element in $\mathcal{T}$ is equivalent to the existence of a minimal element in $\mathcal{A}$.
-If $H(\infty)>\Lambda(0)$, then $\mathcal{A}=[0,\Lambda(0)]$ includes its minimum at 0, and hence a maximal strategy exists.
+By Lemma 3.3, $\mathcal{T}=\{\varphi(\eta):\,\eta\in\mathcal{A}\}$.
+The existence of a maximal element in $\mathcal{T}$ is equivalent to the existence of a minimal element in $\mathcal{A}$, because $\varphi$ is strictly decreasing on $\mathcal{A}$.
+If $H(\infty)>\Lambda(0)$, then $\mathcal{A}=[0,\Lambda(0)]$ includes its minimum at 0, and hence a optimal strategy exists.
 Conversely, if $H(\infty)\leq\Lambda(0)$, then $\mathcal{A}=(\Lambda(0)-H(\infty),\Lambda(0)]$ is left-open and admits no minimum.
@@
 In conclusion, the optimal strategy exists if and only if $H(\infty)>\Lambda(0)$, and when it exists, it is given by $\pi^{(\varphi(0))}$.
-We now turn to the uniformly optimal equilibrium.
-In view of the conclusion in assertion (1), we restrict attention to the case $H(\infty)>\Lambda(0)$.
-Once this condition holds, and following the same line of reasoning as in assertion (1), we immediately obtain that $\pi^{(\varphi(0))}$ is also uniformly optimal, because (4.1) remains valid for every $t\in[0,T]$.
-Finally, we consider the uniformly strictly optimal equilibrium.
-If $\Lambda(t)=0$ for some $t\in[0,T)$, then $v^{(T_{0})}(t)\equiv 0$ for any $T_{0}\in\mathcal{T}$, which implies $J(t,\pi^{(T_{0})})\equiv\exp(\mathcal{L}(0))=1$ for all $\pi^{(T_{0})}\in\Pi_{d}^{e}$.
-Hence, no uniformly strictly optimal equilibrium exists in this case.
-Conversely, assume $\Lambda(t)>0$ for all $t\in[0,T)$ and $H(\infty)>\Lambda(0)$.
-Because $\Lambda(t)>0$ for all $t<T$, we have $\varphi(0)=T$.
-Let $T_{0}\in\mathcal{T}\setminus\{T\}$, by Lemma 4.3, for any $t\in[0,T)$, it holds that
+Following the same line of reasoning as in Assertion (1), we immediately obtain that $\pi^{(\varphi(0))}$ is also uniformly optimal as (4.1) remains valid for every $t\in[0,T]$.
+Assume that $\Lambda(t)=0$ for some $t\in[0,T)$.
+Then $v^{(T_{0})}(t)\equiv 0$ for any $T_{0}\in\mathcal{T}$, which implies $J(t,\pi^{(T_{0})})\equiv\exp(\mathcal{L}(0))=1$ for all $\pi^{(T_{0})}\in\Pi_{d}^{e}$.
+Hence, there is no uniformly strictly optimal equilibrium in this case.
+Conversely, assume that $\Lambda(t)>0$ for all $t\in[0,T)$ and $H(\infty)>\Lambda(0)$.
+We have $\varphi(0)=T$ because $\Lambda(t)>0$ for all $t<T$.
+Let $T_{0}\in\mathcal{T}\setminus\{T\}$.
+By Lemma 4.3, for any $t\in[0,T)$, it holds that
 $v^{(T)}(t)>v^{(T_{0})}(t).$
-Because $\mathcal{L}(z)$ is strictly increasing in $z$, we obtain $J_{0}(t,\pi^{(T)})>J_{0}(t,\pi^{(T_{0})})$.
-Therefore, $\pi^{(\varphi(0))}=\pi^{(T)}$ is the unique uniformly strictly optimal equilibrium.
-∎ Analogous to Proposition 3.8, we now investigate the impact of future market opportunities on the risk exposure vector in the case of $\tilde{\mathbb{E}}[\bm{R}]=\infty$.
-Given the multiplicity of equilibria in this setting, we focus specifically on the optimal equilibrium characterized in Theorem 4.4.
-The following proposition presents the corresponding comparative statics results.
+We obtain $J_{0}(t,\pi^{(T)})>J_{0}(t,\pi^{(T_{0})})$ as $\mathcal{L}(z)$ is strictly increasing in $z$. Therefore, $\pi^{(\varphi(0))}=\pi^{(T)}$ is uniformly strictly optimal. ∎
+In analogy with Proposition 3.8, we now analyze the effect of future market opportunities on the risk exposure vector for the case $\tilde{\mathbb{E}}[\bm{R}]=\infty$.
+Given the multiplicity of equilibria in this setting, we focus on the optimal equilibrium characterized in Theorem 4.4.
+The following proposition presents the corresponding results.
 We omit the proof as it follows a similar line of reasoning to that of Proposition 3.8.
@@
   1. (1)
-    If the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly increasing and $\Lambda_{n}\nearrow H(\infty)$, then
-    $\lim_{n\to\infty}a_{n}(0)=\frac{\lambda_{0}}{r_{0}},$
-    where $r_{0}\triangleq\operatorname{essinf}\bm{R}$. Moreover, $|a_{n}(0)|$ converges monotonically increasing, i.e.,
-    $|a_{n}(0)|\nearrow\frac{|\lambda_{0}|}{r_{0}}.$
+    If the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly increasing and $\Lambda_{n}\nearrow H(\infty)$, then $\lim\limits_{n\to\infty}a_{n}(0)=\frac{\lambda_{0}}{r_{0}}$, where $r_{0}\triangleq\operatorname{essinf}\bm{R}$.
+Moreover, $|a_{n}(0)|\nearrow\frac{|\lambda_{0}|}{r_{0}}$ as $n\to\infty$.
   2. (2)
-    If the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly decreasing and $\Lambda_{n}\searrow 0$, then
-    $\lim_{n\to\infty}a_{n}(0)=\mathbf{0}.$
-    Moreover, $|a_{n}(0)|$ converges monotonically decreasing, i.e.,
-    $|a_{n}(0)|\searrow 0.$
-Here, we adopt the convention that $\frac{1}{0}=\infty$.
+    If the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly decreasing and $\Lambda_{n}\searrow 0$, then $\lim\limits_{n\to\infty}a_{n}(0)=\mathbf{0}$.
+Moreover, $|a_{n}(0)|\searrow 0$ as $n\to\infty$.
+Here we adopt the convention $\frac{1}{0}=\infty$.
 5 Comparative Statics
-In the previous sections, we fully characterized the equilibrium strategies associated with a given distribution of risk aversion $\bm{R}$ and discussed how to select the optimal equilibrium.
-We now turn to comparative statics and ask a natural economic question: how does a shift in the distribution of risk aversion affect the resulting equilibrium investment behavior?
-In this section, we only consider the case where the equilibrium exists.
-To formalize such comparisons, we begin by recalling the notion of first-order stochastic dominance, which provides a standard criterion for ranking risk aversion distributions.
+In the previous sections, we have derived, in closed form, the equilibrium strategies associated with a given distribution of the RRA $\bm{R}$ and discussed how to select the optimal equilibrium.
+We now turn to comparative statics and consider a natural question: how does a shift in the distribution of the RRA affect the resulting equilibrium investment behavior?
+Throughout this section, we write $H_{i}$ for $H_{{\bm{R}_{i}}}$ and $a_{i}$ for $a_{{\bm{R}_{i}}}$, $i=1,2$.
+5.1 Two-Point Distributed RRA and Single-Crossing
+We begin by considering the first-order stochastic dominance, which should be a natural way to rank the risk aversion distributions.
 Definition 5.1.
 Let $\bm{R}_{1}$} and $\bm{R}_{2}$} be two random variables.
-We say that $\bm{R_{1}}$ dominates $\bm{R_{2}}$ in the sense of first-order stochastic dominance, denoted as $\bm{R_{1}}\succeq_{1}\bm{R_{2}}$, if
-$\tilde{\mathbb{P}}(\bm{R_{1}}\geq x)\geq\tilde{\mathbb{P}}(\bm{R_{2}}\geq x)\quad\text{for all }x\in\mathbb{R}.$
-Intuitively, one might expect that if Investor 1 is “more risk-averse” than Investor 2 in the sense of first-order stochastic dominance, then Investor 1 should take a smaller position in the risky asset at all times.
-Surprisingly, this intuition turns out to be false, as the following numerical example shows.
+We say that $\bm{R}_{1}$} dominates $\bm{R}_{2}$} in the sense of first-order stochastic dominance, denoted ${\bm{R}_{1}}\succeq_{1}{\bm{R}_{2}}$, if $\tilde{\mathbb{P}}({\bm{R}_{1}}\geq x)\geq\tilde{\mathbb{P}}({\bm{R}_{2}}\geq x)$ for all $x\in\mathbb{R}$.
+Intuitively, one might expect that if ${\bm{R}_{1}}\succeq_{1}{\bm{R}_{2}}$, then the corresponding risk exposure magnitudes $|a_{i}(\cdot)|$ ($i=1,2$) should satisfy $|a_{1}(t)|\leq|a_{2}(t)|$ for any $t\in[0,T)$.
+Surprisingly, this intuition turns out to be false, as the next example with two-point distributed RRA shows.
 Example 5.2.
-Assume that there is one stock whose market price of risk is constant, $\lambda(\cdot)\equiv\lambda=0.4$. The time horizon $T=20$. The risk-aversion distributions of the two investors are:
-Investor 1: $\tilde{\mathbb{P}}(\bm{R}_{1}=1)=0.9$, $\tilde{\mathbb{P}}(\bm{R}_{1}=3)=0.1$.
-Investor 2: $\tilde{\mathbb{P}}(\bm{R}_{2}=1)=0.9$, $\tilde{\mathbb{P}}(\bm{R}_{2}=2)=0.1$.
+Assume that there is only one stock, whose market price of risk is constant, $\lambda(\cdot)\equiv\lambda=0.4$. The time horizon $T=20$. There are two investors with distributions of RRA:
+  • Investor 1: $\tilde{\mathbb{P}}(\bm{R}_{1}=1)=0.9$ and $\tilde{\mathbb{P}}(\bm{R}_{1}=3)=0.1$;
+  • Investor 2: $\tilde{\mathbb{P}}(\bm{R}_{2}=1)=0.9$ and $\tilde{\mathbb{P}}(\bm{R}_{2}=2)=0.1$.
 Obviously, $\bm{R}_{1}\succeq_{1}\bm{R}_{2}$.
-Figure 1 displays the comparative evolution of $|a_{i}|$, i=1,2i=1,2.555Because $\bar{\pi}=(\sigma^{\top})^{-1}a$, in the one-dimensional case comparing $|a_{i}(t)|$ is equivalent to comparing $|\pi|_{i}$.
-More generally, $|a|$ represents the volatility of the investor’s wealth process and thus captures the level of risk borne by the investor.
-Therefore, in what follows, we focus conceptually on comparing the risk exposure magnitude.
-The orange-shaded Reversal Region highlights that the more risk-averse Investor 1 adopts a strictly more aggressive position than Investor 2 when the time to maturity is sufficiently long.
-This counterexample shows that under first-order stochastic dominance, the magnitudes of equilibrium investment strategies need not be monotone.
-Figure 1: Comparative evolution of $a_{i}(t)$ under first-order stochastic dominance.
-The red solid line corresponds to Investor 1 (more risk-averse), with distribution $\tilde{\mathbb{P}}(\bm{R}_{1}=1)=0.9$, $\tilde{\mathbb{P}}(\bm{R}_{1}=3)=0.1$; the blue dashed line corresponds to Investor 2 (less risk-averse), with distribution $\tilde{\mathbb{P}}(\bm{R}_{2}=1)=0.9$, $\tilde{\mathbb{P}}(\bm{R}_{2}=2)=0.1$.
-The orange shaded region indicates where $|a_{1}(t)|>|a_{2}(t)|$.
-Parameters: $\lambda=0.4$, horizon $T=20$.
-The appearance of this counterexample (Figure 1) may suggest that the search for a general comparative result under first-order stochastic dominance is fundamentally obstructed.
-In fact, the observed reversal is not an artifact of a particular parameter choice.
-We show below that such a reversal in investment magnitudes is unavoidable when the investment horizon $T$ is sufficiently large, provided that the risk-aversion distributions satisfy the conditions stated in the following proposition.
-In the following proposition, for convenience, we extend the market to an infinite horizon and assume that the market price of risk $\lambda$ is defined on $0,\infty)$.
-For every fixed $T>0$, we require $\Lambda(0,T)\triangleq\int_{0}^{T}|\lambda(s)|^{2}\mathrm{d}s<\infty$.
-In addition, we assume that $\lim_{T\to\infty}\Lambda(0,T)=\infty$ and $\lambda(0)\neq 0$.
-As discussed in the preceding sections, for each fixed $T$, we can define the equilibrium strategy $\{\pi_{i}(t,T):t\in[0,T)\}$ associated with the distribution $\bm{R}_{i}$ (as used in the proposition below).
-The proposition shows that, as long as the distributions $\bm{R}_{i}$ satisfy certain conditions, the magnitudes $|\pi_{i}(0,T)|$ must exhibit a strict and deterministic ordering when the horizon $T$ is sufficiently large.
+Figure 1 displays the comparative evolution of $|a_{i}(\cdot)|$ (i=1,2i=1,2).
+[^4] Because $\bar{\pi}=(\sigma^{\top})^{-1}a$, in the one-dimensional case comparing $|a_{i}(\cdot)|$ is equivalent to comparing $|\pi_{i}(\cdot)|$.
+More generally, $|a|$ represents the local volatility of the investor’s wealth process and thus captures the level of risk borne by the investor.
+Therefore, in what follows, we focus conceptually on comparing the risk exposure magnitude (local volatility) instead of the trading strategy.
+The red-shaded Reversal Region highlights that Investor 1 adopts a strictly riskier position than Investor 2 when the time to maturity is sufficiently long.
+This counterexample shows that a “larger” RRA in the sense of first-order stochastic dominance does not necessarily yield a less risky investment.
+Figure 1: Comparative evolution of $|a_{i}(\cdot)|$ ($i=1,2$) under the first-order stochastic dominance.
+The red solid line corresponds to Investor 1 with $\tilde{\mathbb{P}}(\bm{R}_{1}=1)=0.9$ and $\tilde{\mathbb{P}}(\bm{R}_{1}=3)=0.1$; the blue dashed line corresponds to Investor 2 with $\tilde{\mathbb{P}}(\bm{R}_{2}=1)=0.9$ and $\tilde{\mathbb{P}}(\bm{R}_{2}=2)=0.1$.
+The red shaded region indicates where $|a_{1}(\cdot)|>|a_{2}(\cdot)|$.
+Parameters: $\lambda=0.4$, $T=20$.
+The appearance of Example 5.2 might suggest that establishing a general comparative result under the first-order stochastic dominance is fundamentally obstructed.
+In fact, the observed reversal is not merely an artifact of a particular parameter choice.
+In the following proposition, we focus on the case where the risk aversions $\bm{R}_{i}$ ($i=1,2$) satisfy $\operatorname{essinf}\bm{R}_{1}=\operatorname{essinf}\bm{R}_{2}=r_{0}>0$.
+In this case, $H_{i}(\infty)=\infty$ ($i=1,2$) (see Lemma A.2).
+Consequently, according to Theorem 3.2(1), for any deterministic and right-continuous market price of risk $\lambda(\cdot)$ in $L^{2}([0,T))$, the equilibrium strategy exists and is unique.
+The next proposition shows that, under suitable conditions on $\bm{R}_{i}$, in particular when $\bm{R}_{1}$ and $\bm{R}_{2}$ share the same “smallest” value but $\bm{R}_{1}$ has a larger “second-smallest” value than $\bm{R}_{2}$, there exists $\overline{\Lambda}>0$ such that, for any $\lambda(\cdot)$ with $\lambda(0)\neq 0$ and $\int_{0}^{T}|\lambda(t)|^{2}\,\mathrm{d}t>\overline{\Lambda}$, the corresponding risk exposure magnitudes satisfy $|a_{1}(0)|>|a_{2}(0)|$.
+Moreover, if $\tilde{\mathbb{E}}[\bm{R}_{1}]>\tilde{\mathbb{E}}[\bm{R}_{2}]$ and $\lambda(T-)\neq 0$, then $|a_{1}(T-)|<|a_{2}(T-)|$.
 Proposition 5.3.
-Let $\bm{R}_{1}$ and $\bm{R}_{2}$ be two non-negative random variables with finite expectations i.e., $\tilde{\mathbb{E}}[\bm{R}_{i}]<\infty$, i=1, 2i=1,\ 2.
-Assume that the following conditions are satisfied:
-  1. (1)
-    Let $r_{0}=\operatorname{essinf}\bm{R}_{1}=\operatorname{essinf}\bm{R}_{2}>0$, and assume that there is a positive probability mass at $r_{0}$, i.e., $p_{i}=\tilde{\mathbb{P}}(\bm{R}_{i}=r_{0})\in(0,1)$ for $i=1,2$.
-  2. (2)
-    Define the infimum of the first support point of $\bm{R}_{i}$ strictly greater than $r_{0}$ as $\tilde{r}_{i}=\inf\{r\in supp(\bm{R}_{i}):r>r_{0}\}$, and let $\delta_{i}=\tilde{r}_{i}-r_{0}$ with $\delta_{1}>\delta_{2}>0$.
-Under these conditions, for a sufficiently large $T$, we have
-$|\pi_{1}(0,T)|>|\pi_{2}(0,T)|.$
+Let $\bm{R}_{1}$ and $\bm{R}_{2}$ be two non-negative random variables with $\tilde{\mathbb{E}}[\bm{R}_{i}]<\infty$ $(i=1,2)$ satisfying the following two conditions.
+  1. (c1)
+    Let $r_{i}\triangleq\operatorname{essinf}\bm{R}_{i}$, $i=1,2$. Assume that $r_{0}\triangleq r_{1}=r_{2}>0$ and $p_{i}\triangleq\tilde{\mathbb{P}}(\bm{R}_{i}=r_{0})\in(0,1)$ for $i=1,2$.
+  2. (c2)
+    Let $\tilde{r}_{i}\triangleq\inf\{r\in\operatorname*{supp}(\bm{R}_{i}):r>r_{0}\}$ and $\delta_{i}\triangleq\tilde{r}_{i}-r_{0}$, $i=1,2$. Assume that $\delta_{1}>\delta_{2}>0$.
+Then, there exists $\overline{\Lambda}>0$ such that, for any $\lambda(\cdot)$ with $\lambda(0)\neq 0$ and $\int_{0}^{T}|\lambda(t)|^{2}\,\mathrm{d}t>\overline{\Lambda}$, the corresponding risk exposure magnitudes satisfy $|a_{1}(0)|>|a_{2}(0)|$.
+In addition, if $\tilde{\mathbb{E}}[\bm{R}_{1}]>\tilde{\mathbb{E}}[\bm{R}_{2}]$ and $\lambda(T-)\neq 0$, then $|a_{1}(T-)|<|a_{2}(T-)|$.
 Proof.
 See Appendix B.3. ∎
-As investigated by Desmettre2023, we now focus on the two-point distribution, which represents the closest deviation from the standard expected utility model.
-This setting allows us to derive a single crossing property, rigorously characterizing the transition from short-term dominance to the long-term reversal observed in Proposition 5.3.
+Next, under the conditions of Proposition 5.3, we further investigate the crossing properties of the risk exposure magnitudes $|a_{i}(\cdot)|$ ($i=1,2$) corresponding to the two distributions.
+This includes the single-crossing property (Proposition 5.4) as well as the monotonicity of the crossing time w.r.t.
+the distribution parameters (Propositions 5.6 and 5.7).
+To this end, we focus on the two-point distributed RRA, which, as investigated by Desmettre and Steffensen (2023), is the closest RRA model to the standard expected utility framework.
+Specifically, we assume that the risk aversion $\bm{R}_{i}$ ($i=1,2$) satisfies
+$\begin{split}&\tilde{\mathbb{P}}(\bm{R}_{i}=r_{0})=p_{i}=1-\tilde{\mathbb{P}}(\bm{R}_{i}=r_{0}+\delta_{i})\quad\text{with }r_{0}>0,\,p_{i}\in(0,1),\delta_{1}>\delta_{2}>0.\end{split}$ (5.1)
+Below, we begin by establishing the single-crossing property of the risk exposure magnitudes; that is, under certain conditions, there exists a unique time $t^{*}$ such that $|a_{1}(t^{*})|=|a_{2}(t^{*})|$.
 Proposition 5.4.
-Under the conditions of Proposition 5.3, suppose further that $\lambda(t)\neq 0$ for all $t\in[0,T)$, and that $\bm{R}_{1}$ and $\bm{R}_{2}$ follow two-point distributions with $\tilde{\mathbb{E}}[\bm{R}_{1}]>\tilde{\mathbb{E}}[\bm{R}_{2}]$.
-Moreover, assume that $T$ is sufficiently large so that $|a_{1}(0)|>|a_{2}(0)|$.
-Then $|a_{1}(\cdot)|$ and $|a_{2}(\cdot)|$ exhibit a single-crossing property: there exists a unique $t^{*}\in[0,T]$ such that $|a_{1}(t^{*})|=|a_{2}(t^{*})|$.
+Suppose that the risk aversions $\bm{R}_{1}$ and $\bm{R}_{2}$ satisfy (5.1) and $\tilde{\mathbb{E}}[\bm{R}_{1}]>\tilde{\mathbb{E}}[\bm{R}_{2}]$.
+[^5] The condition $\tilde{\mathbb{E}}[\bm{R}_{1}]>\tilde{\mathbb{E}}[\bm{R}_{2}]$ is equivalent to $(1-p_{1})\delta_{1}>(1-p_{2})\delta_{2}$.
+Let $\widehat{\Lambda}\in(0,\infty)$ be the unique solution to the equation
+[…]
+Then, for any $\lambda(\cdot)$ satisfying $\lambda(t)\neq 0$ for all $t\in[0,T)$ and $\lambda(T-)\neq 0$, we have the following assertions:
+  1. (1)
+    If $\int_{0}^{T}|\lambda(t)|^{2}\mathrm{d}t<\widehat{\Lambda}$, then $|a_{1}(t)|<|a_{2}(t)|$ for all $t\in[0,T)$.
+  2. (2)
+    If $\int_{0}^{T}|\lambda(t)|^{2}\mathrm{d}t>\widehat{\Lambda}$, then there exists a unique $t^{*}\in(0,T)$ such that $|a_{1}(t)|<|a_{2}(t)|$ for all $t\in(t^{*},T)$ and $|a_{1}(t)|>|a_{2}(t)|$ for all $t\in[0,t^{*})$.
+  3. (3)
+    If $\int_{0}^{T}|\lambda(t)|^{2}\mathrm{d}t=\widehat{\Lambda}$, then $|a_{1}(0)|=|a_{2}(0)|$ and $|a_{1}(t)|<|a_{2}(t)|$ for all $t\in(0,T)$.
 Proof.
 See Appendix B.4. ∎
-Having established the single crossing property, we investigate how the probability distribution drives the timing of this reversal.
-Intuitively, a larger probability assignment accelerates the strategy’s convergence to the log-utility benchmark, leading to a faster intersection as time recedes from the horizon.
+Remark 5.5.
+It is important to note that the conditions in Propositions 5.3 and 5.4 are compatible with first-order stochastic dominance.
+One can construct risk aversion distributions such that $\bm{R}_{1}\succeq_{1}\bm{R}_{2}$ holds while the conditions in Propositions 5.3 and 5.4 are simultaneously satisfied: letting $p_{1}=p_{2}$ in (5.1).
+For instance, the parameters in Example 5.2 satisfy these conditions with $r_{0}=1$, $\delta_{1}=2$, $\delta_{2}=1$, $\tilde{\mathbb{E}}[\bm{R}_{1}]=1.2>1.1=\tilde{\mathbb{E}}[\bm{R}_{2}]$ and $\lambda\equiv 0.4\neq 0$.
+In particular, Figure 1 illustrates both the reversal in risk exposure magnitude and the single-crossing property.
+Having established the single-crossing property, we next investigate the monotonicity of $t^{*}$ w.r.t. $p\triangleq p_{1}=p_{2}$ under the conditions of Proposition 5.4.
 Proposition 5.6.
-Under the conditions of Proposition 5.3 and Proposition 5.4, assume additionally that $p_{1}=p_{2}=p$.
-Let $t^{*}$ be the intersection time where $|a_{1}(t^{*})|=|a_{2}(t^{*})|$.
-Then, $t^{*}$ is strictly increasing with respect to $p$.
+Let $T>0$ be fixed.
+Assume that the risk aversions $\bm{R}_{1}$ and $\bm{R}_{2}$ satisfy (5.1) with $p_{1}=p_{2}=p\in(0,1)$ and that $\lambda(t)\neq 0$ for all $t\in[0,T)$ with $\lambda(T-)\neq 0$.
+Let $a_{i,p}$ denote the risk exposure vector corresponding to $\bm{R}_{i}$ with $p_{i}=p$ for $i=1,2$.
+Let $t^{*}(p)\in(0,T)$ be the crossing time, where $|a_{1,p}(t^{*}(p))|=|a_{2,p}(t^{*}(p))|$.
+Then the set
+$I\triangleq\left\{p\in(0,1):t^{*}(p)\in(0,T)\right\}$
+is open and the crossing time $t^{*}(\cdot)$ is continuously differentiable in II.
+[^6] The set $I$ is non-empty if $\Lambda(0)$ is sufficiently large.
+Indeed, for any fixed $p\in(0,1)$, Proposition 5.4 ensures that the crossing time $t^{*}(p)$ exists in $(0,T)$ provided that $\Lambda(0)$ exceeds the threshold $\widehat{\Lambda}$ defined therein, which depends on $p$.
+Moreover, its derivative satisfies
+$\frac{\mathrm{d}t^{*}}{\mathrm{d}p}(p)>0\quad\text{for all }p\in I.$
 Proof.
-Recalling that $Q_{i}(x)=\dfrac{\delta_{i}(1-p)e^{-\delta_{i}\frac{x}{2}}}{p+(1-p)e^{-\delta_{i}\frac{x}{2}}}$, we know that $Q_{i}$ is strictly decreasing and satisfies $U_{i}=U_{i}(p)\triangleq Q_{i}(0)=\delta_{i}(1-p)$.
-Moreover, $\Lambda(t)=\int_{0}^{v_{i}(t)}(r_{0}+Q_{i}(x))^{2}\,\mathrm{d}x=2\int_{Q_{i}(v_{i}(t))}^{U_{i}(p)}\dfrac{(r_{0}+q)^{2}}{q(\delta_{i}-q)}\,\mathrm{d}q\triangleq 2\int_{Q_{i}(v_{i}(t))}^{U_{i}(p)}G(q,\delta)\,\mathrm{d}q$ where the second equality arises from $q=Q_{i}(x)$, $\mathrm{d}x=\dfrac{2}{q(q-\delta_{i})}\,\mathrm{d}q$ and $G(q,\delta)\triangleq\dfrac{(r_{0}+q)^{2}}{q(\delta-q)}$.
-At time $t^{*}$, let $Q=Q(p)\triangleq Q_{1}(v_{1}(t^{*}))=Q_{2}(v_{2}(t^{*}))$ and $\Lambda^{*}=\Lambda^{*}(p)\triangleq\Lambda(t^{*})$.
-The condition $|a_{1}(t^{*})|=|a_{2}(t^{*})|$ implies
-$\Lambda^{*}=2\int_{Q}^{U_{1}(p)}G(q,\delta_{1})\,dq=2\int_{Q}^{U_{2}(p)}G(q,\delta_{2})\,dq.$ (5.1)
-Differentiating (5.1) with respect to $p$ and we have
-$\dfrac{1}{2}\dfrac{\mathrm{d}\Lambda^{*}}{\mathrm{d}p}$ $=G(U_{1},\delta_{1})\dfrac{\mathrm{d}U_{1}}{\mathrm{d}p}-G(Q,\delta_{1})\dfrac{\mathrm{d}Q}{\mathrm{d}p}$
-$=-\delta_{1}G(U_{1},\delta_{1})-G(Q,\delta_{1})\dfrac{\mathrm{d}Q}{\mathrm{d}p}.$
-Similarly, for the second distribution, we have
-$\dfrac{1}{2}\dfrac{\mathrm{d}\Lambda^{*}}{\mathrm{d}p}=-\delta_{2}G(U_{2},\delta_{2})-G(Q,\delta_{2})\dfrac{\mathrm{d}Q}{\mathrm{d}p}.$
-Eliminating $\dfrac{\mathrm{d}Q}{\mathrm{d}p}$ from the system, we obtain
-$\dfrac{1}{2}\dfrac{\mathrm{d}\Lambda^{*}}{\mathrm{d}p}=\dfrac{\delta_{2}G(U_{2},\delta_{2})G(Q,\delta_{1})-\delta_{1}G(U_{1},\delta_{1})G(Q,\delta_{2})}{G(Q,\delta_{2})-G(Q,\delta_{1})}.$ (5.2)
-Substituting $U_{i}=\delta_{i}(1-p)$ into $G$, we have
-$\delta_{i}G(U_{i},\delta_{i})=\dfrac{(r_{0}+\delta_{i}(1-p))^{2}}{p(1-p)\delta_{i}}.$
-Recalling the form of $Q_{i}$ and the fact that $Q_{i}$ is strictly decreasing, we have $\delta_{1}>\delta_{2}>U_{2}>Q$, which implies $\delta_{1}-Q>\delta_{2}-Q>0$ and $G(Q,\delta_{1})<G(Q,\delta_{2})$.
-Thus, the denominator of (5.2) is positive.
-For the numerator, we have
-$\dfrac{\delta_{1}G(U_{1},\delta_{1})}{G(Q,\delta_{1})}$ $=\dfrac{(r_{0}+\delta_{1}(1-p))^{2}}{p(1-p)\delta_{1}}\cdot\dfrac{Q(\delta_{1}-Q)}{(r_{0}+Q)^{2}}=\dfrac{Q}{p(1-p)(r_{0}+Q)^{2}}(r_{0}+\delta_{1}(1-p))^{2}\left(1-\dfrac{Q}{\delta_{1}}\right),$
-$\dfrac{\delta_{2}G(U_{2},\delta_{2})}{G(Q,\delta_{2})}$ $=\dfrac{Q}{p(1-p)(r_{0}+Q)^{2}}(r_{0}+\delta_{2}(1-p))^{2}\left(1-\dfrac{Q}{\delta_{2}}\right).$
-Because $\delta_{1}>\delta_{2}$, it holds that $(r_{0}+\delta_{1}(1-p))^{2}>(r_{0}+\delta_{2}(1-p))^{2}$ and $1-\dfrac{Q}{\delta_{1}}>1-\dfrac{Q}{\delta_{2}}$.
-Therefore, $\dfrac{\delta_{1}G(U_{1},\delta_{1})}{G(Q,\delta_{1})}>\dfrac{\delta_{2}G(U_{2},\delta_{2})}{G(Q,\delta_{2})}$, which implies $\delta_{2}G(U_{2},\delta_{2})G(Q,\delta_{1})-\delta_{1}G(U_{1},\delta_{1})G(Q,\delta_{2})<0$.
-Consequently, $\dfrac{\mathrm{d}\Lambda^{*}}{\mathrm{d}p}<0$.
-Because $\Lambda(t)$ is a strictly decreasing function of $t$, a decrease in the required energy $\Lambda^{*}$ corresponds to an increase in the intersection time $t^{*}$.
-∎
-Building on this analysis, we now relax the symmetry assumption to examine the individual effects.
-The following corollary presents the comparative statics of the intersection time with respect to the probability parameter of each investor.
-Corollary 5.6.
-Under the conditions of Proposition 5.3 and Proposition 5.4, we have the following assertions.
-  1. (1)
-    If $p_{2}$ is fixed, $t^{*}$ is strictly increasing with respect $p_{1}$.
-  2. (2)
-    If $p_{1}$ is fixed, $t^{*}$ is strictly decreasing with respect $p_{2}$
+See Appendix B.5. ∎
+The condition $\dfrac{\mathrm{d}t^{*}}{\mathrm{d}p}>0$ indicates that the crossing time to maturity is strictly decreasing with respect to the probability parameter $p$.
+This confirms that a higher common probability of the smallest risk aversion value brings the crossing closer to maturity.
+Building on the analysis in Proposition 5.6, we now relax the assumption $p_{1}=p_{2}$ and investigate the individual effects of the probability parameters $p_{i}$ ($i=1,2$).
+The following proposition establishes the comparative statics of the crossing time w.r.t.
+each investor’s probability parameter.
+Proposition 5.7.
+Let $T>0$ be fixed.
+Assume that the risk aversions $\bm{R}_{1}$ and $\bm{R}_{2}$ satisfy (5.1) with $\tilde{\mathbb{E}}[\bm{R}_{1}]>\tilde{\mathbb{E}}[\bm{R}_{2}]$ and that $\lambda(t)\neq 0$ for all $t\in[0,T)$ with $\lambda(T-)\neq 0$.
+Let $a_{i,p_{i}}$ denote the risk exposure vector corresponding to $\bm{R}_{i}$ with $\tilde{\mathbb{P}}(\bm{R}_{i}=r_{0})=p_{i}$ for $i=1,2$.
+Let $t^{*}(p_{1},p_{2})\in(0,T)$ denote the crossing time, where $|a_{1,p_{1}}(t^{*}(p_{1},p_{2}))|=|a_{2,p_{2}}(t^{*}(p_{1},p_{2}))|$.
+Let
+$O\triangleq\left\{(p_{1},p_{2})\in(0,1)^{2}:(1-p_{1})\delta_{1}>(1-p_{2})\delta_{2}\text{ and }t^{*}(p_{1},p_{2})\in(0,T)\right\}.$
+Then, the set $O$ is open [^7] The set $O$ is nonempty provided that $\Lambda(0)$ is sufficiently large.
+The argument is analogous to that in Proposition 5.6.
+and the function $t^{*}(\cdot,\cdot)$ is continuously differentiable in $O$.
+Moreover, for any $(p_{1},p_{2})\in O$, the partial derivatives satisfy
+$\frac{\partial t^{*}}{\partial p_{1}}(p_{1},p_{2})>0\quad\text{and}\quad\frac{\partial t^{*}}{\partial p_{2}}(p_{1},p_{2})<0.$
 Proof.
-We retain the notations from the proof of Proposition 5.4.
-Differentiating (5.1) with respect to $p_{1}$ yields
-$\dfrac{1}{2}\dfrac{\partial\Lambda^{*}}{\partial p_{1}}$ $=-\delta_{1}G(U_{1},\delta_{1})-G(Q,\delta_{1})\dfrac{\partial Q}{\partial p_{1}},$
-$\dfrac{1}{2}\dfrac{\partial\Lambda^{*}}{\partial p_{1}}$ $=-G(Q,\delta_{2})\dfrac{\partial Q}{\partial p_{1}}.$
-Eliminating $\dfrac{\partial Q}{\partial p_{1}}$, we obtain
-$\dfrac{1}{2}\dfrac{\partial\Lambda^{*}}{\partial p_{1}}=\dfrac{-\delta_{1}G(U_{1},\delta_{1})G(Q,\delta_{2})}{G(Q,\delta_{2})-G(Q,\delta_{1})}.$
-Because the denominator is positive and the numerator is negative, we have $\dfrac{\partial\Lambda^{*}}{\partial p_{1}}<0$. Given that $\Lambda(t)$ is strictly decreasing, $t^{*}$ must increase.
-Differentiating (5.1) with respect to $p_{2}$ yields
-$\dfrac{1}{2}\dfrac{\partial\Lambda^{*}}{\partial p_{2}}$ $=-G(Q,\delta_{1})\dfrac{\partial Q}{\partial p_{2}},$
-$\dfrac{1}{2}\dfrac{\partial\Lambda^{*}}{\partial p_{2}}$ $=-\delta_{2}G(U_{2},\delta_{2})-G(Q,\delta_{2})\dfrac{\partial Q}{\partial p_{2}}.$
-Eliminating $\dfrac{\partial Q}{\partial p_{2}}$, we have
-$\dfrac{1}{2}\dfrac{\partial\Lambda^{*}}{\partial p_{2}}=\dfrac{-\delta_{2}G(U_{2},\delta_{2})G(Q,\delta_{1})}{G(Q,\delta_{1})-G(Q,\delta_{2})}.$
-Because the denominator is negative while the numerator remains negative, we have $\dfrac{\partial\Lambda^{*}}{\partial p_{2}}>0$, which implies that $t^{*}$ must decrease. ∎
-It is important to note that the conditions in Proposition 5.3 and Proposition 5.4 are compatible with first-order stochastic dominance.
-One can construct risk aversion distributions such that $\bm{R}_{1}\succeq_{1}\bm{R}_{2}$ holds while the conditions in Proposition 5.3 and 5.4 are simultaneously satisfied.
-For instance, the parameters in Figure 1 satisfy these conditions with $r_{0}=1$, $\delta_{1}=4$, $\delta_{2}=1$, $\tilde{\mathbb{E}}[\bm{R}_{1}]=3>1.5=\tilde{\mathbb{E}}[\bm{R}_{2}]$ and $\lambda\equiv 0.4$.
-Consequently, first-order stochastic dominance fails to rule out the reversal of investment magnitudes characterized by $|\pi_{1}(0,T)|>|\pi_{2}(0,T)|$ for a sufficiently large $T$ and we can also observe the single crossing property.
-However, we demonstrate that the desired monotonicity can be restored under a stronger stochastic order.
-Specifically, we introduce the concept of reverse hazard rate order and show that if $\bm{R_{1}}$ dominates $\bm{R_{2}}$ in this order, the intuition holds, i.e., $|\pi_{1}(t)|\leq|\pi_{2}(t)|$.
+See Appendix B.6. ∎
+Similar to Proposition 5.6, the differentiability of $t^{*}$ established in Proposition 5.7 implies that the crossing time varies smoothly with respect to the probability parameters.
+Specifically, $t^{*}$ is strictly increasing in $p_{1}$ and strictly decreasing in $p_{2}$.
+This indicates that a higher probability of the smallest risk aversion for the first investor brings the crossing closer to maturity, whereas for the second investor, it pushes the crossing farther away from maturity.
+5.2 Reverse Hazard Rate Order
+While a higher RRA does not guarantee a less risky investment under first-order stochastic dominance, we demonstrate that it does guarantee such a reduction under the stronger reverse hazard rate order.
+We first recall the definition of the reverse hazard rate order (see Capéraà (1988) and Shaked and Shanthikumar (2007)).
 Definition 5.8.
 Let $\bm{R}_{1}$} and $\bm{R}_{2}$} be two random variables with cumulative distribution functions $F_{{\bm{R}_{1}}}$ and $F_{{\bm{R}_{2}}}$, respectively.
-We say that $\bm{R_{1}}$ dominates $\bm{R_{2}}$ in the reverse hazard rate order, denoted as $\bm{R_{1}}\succeq_{rh}\bm{R_{2}}$, if the ratio
-$\frac{F_{\bm{R_{2}}}(\gamma)}{F_{\bm{R_{1}}}(\gamma)}$
-is decreasing in $\gamma$ for all $\gamma$ in the union of the supports of $\bm{R_{1}}$ and $\bm{R_{2}}$ where $F_{\bm{R_{1}}}(\gamma)>0$.
-The following proposition compares the equilibrium strategies of two investors ranked by the reverse hazard rate order.
+We say that $\bm{R}_{1}$} dominates $\bm{R}_{2}$} in the reverse hazard rate order, denoted ${\bm{R}_{1}}\succeq_{rh}{\bm{R}_{2}}$, if the ratio $F_{{\bm{R}_{2}}}/F_{{\bm{R}_{1}}}$ is decreasing [^8] Herein, “increasing” means “non-decreasing” and “decreasing” means “non-increasing.” in $\{\gamma:F_{{\bm{R}_{1}}}(\gamma)>0\}\cup\{\gamma:F_{{\bm{R}_{2}}}(\gamma)>0\}$.
+Here, we use the convention $\frac{p}{0}=\infty$ for $p>0$.
+The following proposition establishes a comparison of the risk exposure magnitudes under the reverse hazard rate order.
 Proposition 5.9.
-Let $\bm{R_{1}}$ and $\bm{R_{2}}$} be two non-negative random variables with $\tilde{\mathbb{E}}\left[\bm{R_{i}}\right]<+\infty$ ($i=1,2$), such that $\bm{R_{1}}\succeq_{rh}\bm{R_{2}}$ and their distributions are distinct.
-Let $\pi_{i}(t)=(\sigma^{\top})^{-1}a_{i}(t)$ be the corresponding deterministic equilibrium strategies.
-For any $t\in[0,T)$, the magnitude of $a_{i}(t)$ satisfies
-$|a_{1}(t)|\leq|a_{2}(t)|.$
-Moreover, if $\lambda(t)\neq 0$, the inequality is strict. As a result, the investor with higher risk aversion adopts a less aggressive equilibrium strategy. i.e., $|\pi_{1}(t)|\leq|\pi_{2}(t)|$.
+Let $\bm{R}_{1}$} and $\bm{R}_{2}$} be two non-negative random variables with $\tilde{\mathbb{E}}[{\bm{R}_{i}}]<\infty$ ($i=1,2$) and $F_{\bm{R}_{1}}\neq F_{\bm{R}_{2}}$.
+If ${\bm{R}_{1}}\succeq_{rh}{\bm{R}_{2}}$ and $H_{2}(\infty)>\Lambda(0)$, then $|a_{1}(t)|\leq|a_{2}(t)|$ for any $t\in[0,T)$, where the strict inequality holds if $\lambda(t)\neq 0$.
 Proof.
-Because $h_{1}(x)<h_{2}(x)$ for all positive $s$, it follows that $H_{1}(y)>H_{2}(y)$ for all $y>0$.
-This implies that the inverse functions satisfy $H_{2}^{-1}(z)<H_{2}^{-1}(z)$ for all $z>0$.
-From Theorem 3.2 (1), we have $v_{i}(t)=H_{i}^{-1}(\Lambda(t))$, where $\Lambda(t)=\int_{t}^{T}|\lambda(s)|^{2}ds$ and $v_{i}(t)=\int_{t}^{T}|a_{i}(s)|^{2}\mathrm{d}s$.
-Consequently, monotonicity of the inverse functions implies $v_{1}(t)\leq v_{2}(t)$.
-By (2.5), we have $|a_{i}(t)|=h(v_{i}(t),\bm{R_{i}})|\lambda(t)|$.
-From Proposition A.6, we know that $h_{1}(v_{1}(t))<h_{2}(v_{1}(t))$.
-Because $v_{1}(t)\leq v_{2}(t)$, it holds that $h_{2}(v_{1}(t))\leq h_{2}(v_{2}(t))$.
-Combining these relations yields $h_{1}(v_{1}(t))<h_{2}(v_{2}(t))$.
-Multiplying by $|\lambda(t)|$, we obtain $|a_{1}(t)|\leq|a_{2}(t)|$.
-If $\lambda(t)\neq 0$, the strict inequality $|a_{1}(t)|<|a_{2}(t)|$ is preserved.
-∎
-For the case where the expected risk aversion is infinite, assume that the optimal equilibria exist for both $\bm{R}_{1}$ and $\bm{R}_{2}$.
-Let $a_{i}(t)$ denote the deterministic coefficient corresponding to the optimal equilibrium strategy for $\bm{R}_{i}$ ($i=1,2$).
-Note that $\varphi(0)=\min\{t\in[0,T]\mid\Lambda(t)=0\}$ is determined solely by the market parameters and is therefore identical for both investors.
-Consequently, we can establish the comparative result by examining the magnitudes of $a_{i}(t)$ on the intervals $0,\varphi(0))$ and $\varphi(0),T]$ separately.
-Specifically, on the interval $0,\varphi(0))$, the analysis parallels the finite expectation case derived in Proposition 5.8, and thus the monotonicity of the equilibrium strategies is preserved.
-On the interval $\varphi(0),T]$, the definition of $\varphi(0)$ implies that $\lambda(t)=0$ almost everywhere.
-Consequently, the equilibrium strategies for both investors vanish, trivially satisfying the comparison inequality.
-Based on this decomposition, we have the following corollary.
+See Appendix B.7. ∎
+We now consider the case where the expectation of risk aversion is infinite.
+Assume that the optimal equilibria exist for both $\bm{R}_{1}$ and $\bm{R}_{2}$.
+Note that $\varphi(0)=\min\{t\in[0,T]:\,\Lambda(t)=0\}$ is determined solely by the market parameters and is therefore identical for both investors.
+Consequently, the comparative result can be established by examining the magnitudes of $a_{i}(t)$ on the intervals $0,\varphi(0))$ and $\varphi(0),T]$ separately.
+Specifically, on the interval $0,\varphi(0))$, the analysis parallels the finite expectation case derived in Proposition 5.9, and thus the order of the risk exposure magnitude is preserved.
+On the interval $\varphi(0),T)$, the definition of $\varphi(0)$ implies that $\lambda\equiv\textbf{0}$.
+As a result, the corresponding risk exposure magnitude also satisfies the desired order.
+Therefore, we obtain the following corollary.
 Corollary 5.10.
-Let $\bm{R}_{1}$ and $\bm{R}_{2}$ be two non-negative random variables with $\tilde{\mathbb{E}}[\bm{R}_{i}]=\infty$ ($i=1,2$), such that $\bm{R}_{1}\succeq_{rh}\bm{R}_{2}$ and their distributions are distinct.
-Assume that the optimal equilibria exist for both investors (i.e., $H_{i}(\infty)>\Lambda(0)$, $i=1,2$).
-Let $\pi_{i}(t)$ be the corresponding deterministic optimal equilibrium strategies.
-Then, for any $t\in[0,T)$, the magnitude of $\pi_{i}(t)$ satisfies
-$|\pi_{1}(t)|\leq|\pi_{2}(t)|.$
-Moreover, if $\lambda(t)\neq 0$, the inequality is strict.
-While the previous results focus on comparing distinct distributions ranked by stochastic orders, it is also of theoretical interest to understand the effect of preference aggregation.
-We now investigate the equilibrium strategy when the risk aversion parameter is formed by a convex combination (i.e., a weighted average) of independent and identically distributed components.
+Let $\bm{R}_{1}$ and $\bm{R}_{2}$ be two non-negative random variables with $\tilde{\mathbb{E}}[\bm{R}_{i}]=\infty$ ($i=1,2$) and $F_{\bm{R}_{1}}\neq F_{\bm{R}_{2}}$.
+Assume that $\bm{R}_{1}\succeq_{rh}\bm{R}_{2}$ and the optimal equilibria exist for both investors (i.e., $H_{1}(\infty)\geq H_{2}(\infty)>\Lambda(0)$).
+Let $a_{i}(\cdot)$ be the risk exposure vector corresponding to the optimal equilibrium strategies.
+Then $|a_{1}(t)|\leq|a_{2}(t)|$ for any $t\in[0,T)$, where the strict inequality holds if $\lambda(t)\neq 0$.
+5.3 Aggregation of Risk Aversions
+The preceding results focus on comparisons between investments for distinct risk-aversion distributions ordered by stochastic dominance relations.
+We next study the effect of preference aggregation on the risk exposure magnitude.
+The following proposition shows that a convex combination of i.i.d.
+risk aversions leads to a reduction in the risk exposure magnitude.
 Proposition 5.11.
 Let $\bm{R}_{1},\cdots,\bm{R}_{n}$ be i.i.d.
-non-negative random variables with $\tilde{\mathbb{E}}\left[\bm{R_{i}}\right]<+\infty$ ($i=1,2,\cdots,n$) and $\bm{R}=\sum_{i=1}^{n}w_{i}\bm{R}_{i}$ be a convex combination with $w_{i}\in(0,1)$ and $\sum_{i=1}^{n}w_{i}=1$.
-Let $a(t)$ and $a_{i}(t)$ denote the risk exposure vector associated to the risk aversion $\bm{R}$ and $\bm{R}_{i}$, respectively.
-Then we have
-$|a(t)|\leq|a_{1}(t)|\quad\text{for all }t\in[0,T).$
-If $\bm{R}_{i}$ is non-degenerate and $\lambda(t)\neq 0$, the inequality is strict.
+non-negative random variables with $\tilde{\mathbb{E}}\left[{\bm{R}_{1}}\right]<\infty$ and $\bm{R}\triangleq\sum_{i=1}^{n}w_{i}\bm{R}_{i}$ be a convex combination with $w_{i}\in(0,1)$ ($1\leq i\leq n$) and $\sum_{i=1}^{n}w_{i}=1$.
+Let $H_{1}=H_{\bm{R}_{1}}$ and $H=H_{\bm{R}}$.
+Assume that $H_{1}(\infty)>\Lambda(0)$.
+[^9] Under this assumption, the equilibrium strategy associated with each $\bm{R}_{i}$ exists and is unique by Theorem 3.2(1).
+Moreover, the proof in Appendix B.8 shows that $H(\infty)>\Lambda(0)$ also holds.
+Let $a(\cdot)$ and $a_{1}(\cdot)$ denote the risk exposure vector associated with the risk aversions $\bm{R}$ and $\bm{R}_{1}$, respectively.
+Then $|a(t)|\leq|a_{1}(t)|$ for all $t\in[0,T)$, where the strict inequality holds if $\bm{R}_{1}$ is non-degenerate and $\lambda(t)\neq 0$.
 Proof.
-Let $h(x)$ and $v(t)$ (resp. $h_{i}(x)$ and $v_{i}(t)$) be the functions associated to $\bm{R}$ (resp. $\bm{R_{i}}$ ). From the definition of $\bm{R}$, we have
-$\dfrac{1}{h(x)}=\sum_{j=1}^{n}w_{j}\dfrac{\tilde{\mathbb{E}}[\bm{R}_{j}e^{-\frac{1}{2}w_{j}\bm{R}_{j}x}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}w_{j}\bm{R}_{j}x}]}$ (5.3)
-Because $\bm{R}_{j}$ are identically distributed to $\bm{R}_{1}$, each fraction in the sum is exactly $\dfrac{1}{h_{1}(w_{j}x)}$. Thus (5.3) is actually
-$\dfrac{1}{h(x)}=\sum_{j=1}^{n}w_{j}\dfrac{1}{h_{1}(w_{j}x)}$
-According to Proposition A.6, we know that $\dfrac{1}{h_{1}(x)}$ is decreasing in $x$. Because $w_{j}\in(0,1)$ and $x>0$, we have $w_{j}x<x$ and $\dfrac{1}{h_{1}(w_{j}x)}\geq\dfrac{1}{h_{1}(x)}$ and
-$\dfrac{1}{h(x)}=\sum_{j=1}^{n}w_{j}\dfrac{1}{h_{1}(w_{j}x)}\geq\sum_{j=1}^{n}w_{j}\dfrac{1}{h_{1}(x)}=\dfrac{1}{h_{1}(x)}$
-This implies $h(x)\leq h_{1}(x)$. Analogous to the proof in Proposition 5.8, we have $v(t)\leq v_{1}(t)$. Finally, because $h(x)$ is increasing in $x$, we have
-$|a(t)|=h(v(t))|\lambda(t)|\leq h(v_{1}(t))|\lambda(t)|\leq h_{1}(v_{1}(t))|\lambda(t)|=|a_{1}(t)|$
-If $\bm{R}_{i}$ is non-degenerate and $\lambda(t)\neq 0$, $h$ is strictly increasing, and all inequalities in the above proof become strict. Thus we have $|a(t)|<|a_{1}(t)|$
+See Appendix B.8. ∎
 It is important to note that the inequality established above relies critically on the i.i.d.
 assumption.
-The following example demonstrates that if we retain independence but relax the identical distribution assumption, the equilibrium strategy for the convex combination need not be bounded by the strategies of the individual components.
+The following example demonstrates that, if independence is retained but the identical distribution assumption is relaxed, the conclusion of Proposition 5.11 may no longer hold.
 Example 5.12.
-Assume that there is one stock whose market price of risk is constant, $\lambda(\cdot)\equiv\lambda=0.5$. The time horizon $T=50$. The risk-aversion distributions of the two investors are:
-Investor 1: $\tilde{\mathbb{P}}(\bm{R}_{1}=0.1)=0.2$, $\tilde{\mathbb{P}}(\bm{R}_{1}=8)=0.8$.
+Assume that there is only one stock, whose market price of risk is constant, $\lambda(\cdot)\equiv\lambda=0.5$. The time horizon $T=50$. The risk-aversion distributions of the two investors are:
+  • Investor 1: $\tilde{\mathbb{P}}(\bm{R}_{1}=0.1)=0.2$ and $\tilde{\mathbb{P}}(\bm{R}_{1}=8)=0.8$;
   • Investor 2: $\tilde{\mathbb{P}}(\bm{R}_{2}=1.5)=1$.
-Assume that $\bm{R}_{1}$ and $\bm{R}_{2}$ are independent. We consider a composite investor with risk aversion $\bm{R}=0.5\bm{R}_{1}+0.5\bm{R}_{2}$.
-Figure 2 displays the comparative evolution of $|a(t)|$, $|a_{1}(t)|$ and $|a_{2}(t)|$.
-The red-shaded Reversal Region highlights that the investor with the mixed risk aversion $\bm{R}$ adopts a strictly more aggressive position than both Investor 1 and Investor 2 when the time to maturity is sufficiently long.
-This counterexample shows that without the i.i.d.
-assumption, the magnitudes of equilibrium investment strategies for a convex combination of risk aversions need not be bounded by the strategies of the individual components.
-Figure 2: Comparative evolution of $a(t)$ for a convex combination of risk aversions.
-The black dashed line corresponds to Investor 1, with distribution $\tilde{\mathbb{P}}(\bm{R}_{1}=0.1)=0.2,\tilde{\mathbb{P}}(\bm{R}_{1}=8.0)=0.8$; the blue dash-dotted line corresponds to Investor 2, with degenerate distribution $\tilde{\mathbb{P}}(\bm{R}_{2}=1.5)=1.0$.
-The red solid line corresponds to the mixed investor with risk aversion $\bm{R}=0.5\bm{R}_{1}+0.5\bm{R}_{2}$.
+Obviously, $\bm{R}_{1}$ and $\bm{R}_{2}$ are independent.
+Consider an aggregated investor whose risk aversion is given by $\bm{R}=0.5\bm{R}_{1}+0.5\bm{R}_{2}$.
+Figure 2 illustrates the comparative evolution of the risk exposure magnitudes $|a(\cdot)|$, $|a_{1}(\cdot)|$, and $|a_{2}(\cdot)|$.
+The red-shaded Reversal Region highlights that the investor with the aggregated risk aversion $\bm{R}$ adopts a strictly more aggressive position than both Investor 1 and Investor 2 when the time to maturity is sufficiently long.
+Figure 2: Comparative evolution of risk exposure magnitudes for individual risk aversions and their convex combination.
+The black dashed line corresponds to Investor 1 with $\tilde{\mathbb{P}}(\bm{R}_{1}=0.1)=0.2$ and $\tilde{\mathbb{P}}(\bm{R}_{1}=8)=0.8$; the blue dash-dotted line corresponds to Investor 2 with degenerate distribution $\tilde{\mathbb{P}}(\bm{R}_{2}=1.5)=1$; the red solid line corresponds to the aggregated investor with risk aversion $\bm{R}=0.5\bm{R}_{1}+0.5\bm{R}_{2}$.
 The red-shaded Reversal Region indicates where $|a(t)|>\max\{|a_{1}(t)|,|a_{2}(t)|\}$.
-Parameters: $\lambda=0.5$, horizon $T=50$.
-We now specifically investigate the aggregation of i.i.d.
-risk aversions.
-Distinct from the general result above, the following proposition utilizes the i.i.d.
-property to prove that the equilibrium strategy converges monotonically to its limit.
+Parameters: $\lambda=0.5$, $T=50$.
+Proposition 5.11 establishes a comparison result for convex combinations of i.i.d.
+risk aversions with a fixed population size, but does not address the asymptotic behavior as the number of samples increases.
+We therefore focus on the special case of the sample mean risk aversion.
+The following proposition examines the limit of the risk exposure vector as the sample size $n$ tends to infinity and establishes monotonic convergence by exploiting the structure of the sample mean.
 Proposition 5.13.
-Suppose that $\{\bm{R}_{i}\}_{i=1}^{\infty}$ is a sequence of independent and identically distributed non-negative random variables with $\tilde{\mathbb{E}}\left[\bm{R_{i}}\right]=\mu\in(0,+\infty)$ ($i=1,2,\cdots$,).
-Let $S_{n}=\frac{1}{n}\sum_{i=1}^{n}\bm{R}_{i}$ denote the sample mean risk aversion and $a_{n}(t)$ denote the risk exposure vector associated to the risk aversion $S_{n}$.
+Suppose that $\{\bm{R}_{i}\}_{i=1}^{\infty}$ is a sequence of i.i.d.
+non-negative random variables with $\tilde{\mathbb{E}}\left[{\bm{R}_{1}}\right]=\mu\in(0,\infty)$.
+Let $\bm{Y}_{n}\triangleq\frac{1}{n}\sum_{i=1}^{n}\bm{R}_{i}$ denote the sample mean risk aversion.
+Let $H_{1}$ (resp.
+$\tilde{H}_{n}$) denote the function defined in (3.2) corresponding to $\bm{R}_{1}$(resp.
+$\bm{Y}_{n}$).
+Assume that $H_{1}(\infty)>\Lambda(0)$.
+Then the deterministic equilibrium strategies associated with each $\bm{R}_{i}$} and $\bm{Y}_{n}$ exist and are unique.
+Let $\tilde{a}_{n}(\cdot)$ be the risk exposure vector associated to the risk aversion $\bm{Y}_{n}$.
 Then, as $n\to\infty$, we have
-$|a_{n}(t)|\searrow\frac{|\lambda(t)|}{\mu}\quad\text{for all }t\in[0,T).$ (5.4)
+$|\tilde{a}_{n}(t)|\searrow\frac{|\lambda(t)|}{\mu}\quad\text{for all }t\in[0,T).$ (5.2)
 Proof.
-Let $h_{n}(x)$ and $v_{n}(t)$ be the functions associated to $S_{n}$.
-First, substituting $S_{n}=\frac{1}{n}\sum_{i=1}^{n}\bm{R}_{i}$ into the definition of $h_{n}(x)$ and utilizing the i.i.d.
-property of $\{\bm{R}_{i}\}$, we obtain
-$h_{n}(x)=\frac{\tilde{\mathbb{E}}\left[\prod_{i=1}^{n}e^{-\frac{x}{2n}\bm{R}_{i}}\right]}{\displaystyle\frac{1}{n}\sum_{j=1}^{n}\tilde{\mathbb{E}}\left[\bm{R}_{j}e^{-\frac{x}{2n}\bm{R}_{j}}\prod_{k\neq j}e^{-\frac{x}{2n}\bm{R}_{k}}\right]}$ $=\frac{(\tilde{\mathbb{E}}\left[e^{-\frac{x}{2n}\bm{R}_{1}}\right])^{n}}{\tilde{\mathbb{E}}\left[\bm{R}_{1}e^{-\frac{x}{2n}\bm{R}_{1}}\right](\tilde{\mathbb{E}}\left[e^{-\frac{x}{2n}\bm{R}_{1}}\right])^{n-1}}$
-$=\frac{\tilde{\mathbb{E}}\left[e^{-\frac{1}{2}\bm{R}_{1}(\frac{x}{n})}\right]}{\tilde{\mathbb{E}}\left[\bm{R}_{1}e^{-\frac{1}{2}\bm{R}_{1}(\frac{x}{n})}\right]}=h_{1}\left(\frac{x}{n}\right).$
-According to Proposition A.6, $h_{1}(x)$ is increasing.
-For any fixed $x\geq 0$, the sequence $\frac{x}{n}$ is decreasing in $n$, implying that $h_{n}(x)$ is decreasing in $n$.
-Furthermore, because $H_{n}(y)=\int_{0}^{y}h_{n}^{-2}(z)dz$ and $v_{n}(t)=H_{n}^{-1}(\Lambda(t))$, the monotonicity of $h_{n}$ implies that $v_{n}(t)$ is strictly decreasing in $n$.
-Then we have
-$h_{n+1}(v_{n+1}(t))\leq h_{n}(v_{n+1}(t))\leq h_{n}(v_{n}(t)),$
-equivalently, we have
-$|a_{n+1}(t)|\leq|a_{n}(t)|,\quad\forall t\in[0,T).$
-Now we determine the limit of $|a_{n}(t)|$.
-By the Strong Law of Large Numbers, $S_{n}\to\mu$ almost surely as $n\to\infty$.
-This implies that $S_{n}$ converges in distribution to the degenerate random variable $R\equiv\mu$.
-Moreover, because $\tilde{\mathbb{E}}[S_{n}]=\mu$ for all $n$, the assumptions of Proposition 3.8 are satisfied.
-Thus we conclude that
-$\lim_{n\to\infty}|a_{n}(t)|=h(v(t))|\lambda(t)|=\frac{|\lambda(t)|}{\mu}.$
-Combining this with the monotonicity established above, we obtain (5.4). ∎
+See Appendix B.9. ∎
 Acknowledgement
@@
 $h^{\prime}(x)=\frac{1}{2}\frac{\tilde{\mathbb{E}}\left[e^{-\frac{1}{2}\bm{R}x}\right]\tilde{\mathbb{E}}\left[\bm{R}^{2}e^{-\frac{1}{2}\bm{R}x}\right]-\left(\tilde{\mathbb{E}}\left[\bm{R}e^{-\frac{1}{2}\bm{R}x}\right]\right)^{2}}{\left(\tilde{\mathbb{E}}\left[\bm{R}e^{-\frac{1}{2}\bm{R}x}\right]\right)^{2}}\quad\text{for all }x>0.$
-By Hölder’s inequality,
-$\left(\tilde{\mathbb{E}}\left[\bm{R}e^{-\frac{1}{2}\bm{R}x}\right]\right)^{2}=\left(\tilde{\mathbb{E}}\left[\bm{R}e^{-\frac{1}{2}\bm{R}x}\right]\cdot\tilde{\mathbb{E}}\left[\bm{R}e^{-\frac{1}{2}\bm{R}x}\right]\right)\leq\tilde{\mathbb{E}}\left[e^{-\frac{1}{2}\bm{R}x}\right]\tilde{\mathbb{E}}\left[\bm{R}^{2}e^{-\frac{1}{2}\bm{R}x}\right].$ (A.1)
+By the Cauchy-Schwartz inequality,
+$\left(\tilde{\mathbb{E}}\left[\bm{R}e^{-\frac{1}{2}\bm{R}x}\right]\right)^{2}\leq\tilde{\mathbb{E}}\left[e^{-\frac{1}{2}\bm{R}x}\right]\tilde{\mathbb{E}}\left[\bm{R}^{2}e^{-\frac{1}{2}\bm{R}x}\right].$ (A.1)
 Hence, $h^{\prime}\geq 0$ and $h$ is increasing.
-The inequality in (A.1) becomes an equality if and only if $\tilde{\mathbb{E}}\left[e^{-\frac{1}{2}\bm{R}x}\right]\bm{R}^{2}e^{-\frac{1}{2}\bm{R}x}=e^{-\frac{1}{2}\bm{R}x}\tilde{\mathbb{E}}\left[\bm{R}^{2}e^{-\frac{1}{2}\bm{R}x}\right]$, which is equivalent to $\bm{R}$ following a single-point distribution.
+The equality in (A.1) holds if and only if $\tilde{\mathbb{E}}\left[e^{-\frac{1}{2}\bm{R}x}\right]\bm{R}^{2}e^{-\frac{1}{2}\bm{R}x}=e^{-\frac{1}{2}\bm{R}x}\tilde{\mathbb{E}}\left[\bm{R}^{2}e^{-\frac{1}{2}\bm{R}x}\right]$, which is equivalent to $\bm{R}$ following a single-point distribution.
 Therefore, $h$ is strictly increasing unless $\bm{R}$ is a constant.
@@
 Lemma A.2 (Asymptotic Analysis).
-Let $\bm{R}$ be a non-negative random variable, and let $r_{0}\triangleq\operatorname{essinf}\bm{R}$. Then,
-$\lim_{x\to\infty}h(x)=\frac{1}{r_{0}},$
-where we adopt the convention that $\frac{1}{0}=\infty$.
+Let $\bm{R}$ be a non-negative random variable and let $r_{0}\triangleq\operatorname{essinf}\bm{R}$.
+Then $\lim\limits_{x\to\infty}h(x)=\frac{1}{r_{0}}$, where we adopt the convention $\frac{1}{0}=\infty$.
 Proof.
-Define $\bm{X}\triangleq\bm{R}-r_{0}$, then
+Let $\bm{X}\triangleq\bm{R}-r_{0}$. Then
 $\frac{1}{h(x)}=\frac{\tilde{\mathbb{E}}[\bm{R}e^{-\frac{1}{2}\bm{R}x}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}x}]}=\frac{\tilde{\mathbb{E}}[(r_{0}+\bm{X})e^{-\frac{1}{2}(r_{0}+\bm{X})x}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}(r_{0}+\bm{X})x}]}=r_{0}+Q(x),$ (A.2)
 where $Q(x)\triangleq\frac{\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}\bm{X}x}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}x}]}$.
-For any arbitrary $\varepsilon>0$, because $\operatorname{essinf}\bm{X}=0$, there exists a probability mass $p_{\varepsilon}\triangleq\tilde{\mathbb{P}}(0\leq\bm{X}\leq\varepsilon)>0$.
+By $\operatorname{essinf}\bm{X}=0$, we have $p_{\varepsilon}\triangleq\tilde{\mathbb{P}}(0\leq\bm{X}\leq\varepsilon)>0$ for any $\varepsilon>0$.
 Therefore,
-$Q(x)\leq\frac{\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}\bm{X}x}\mathbb{I}_{\{\bm{X}\leq\varepsilon\}}+\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}\bm{X}x}\mathbb{I}_{\{\bm{X}>\varepsilon\}}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}x}\mathbb{I}_{\{\bm{X}\leq\varepsilon\}}]}\leq\varepsilon+\frac{\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}\bm{X}x}\mathbb{I}_{\{\bm{X}>\varepsilon\}}]}{p_{\varepsilon}e^{-\frac{1}{2}\varepsilon x}}=\varepsilon+\frac{\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}(\bm{X}-\varepsilon)x}\mathbb{I}_{\{\bm{X}>\varepsilon\}}]}{p_{\varepsilon}}.$
-Note that for any $x\geq 1$, $\bm{X}e^{-\frac{1}{2}(\bm{X}-\varepsilon)x}\mathbb{I}_{\{\bm{X}>\varepsilon\}}$ is dominated by $\bm{X}e^{-\frac{1}{2}(\bm{X}-\varepsilon)}\mathbb{I}_{\{\bm{X}>\varepsilon\}}$, which is integrable regardless of whether $\tilde{\mathbb{E}}[\bm{X}]$ is finite.
+$Q(x)\leq$ $\frac{\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}\bm{X}x}\mathbb{I}_{\{\bm{X}\leq\varepsilon\}}+\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}\bm{X}x}\mathbb{I}_{\{\bm{X}>\varepsilon\}}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}x}\mathbb{I}_{\{\bm{X}\leq\varepsilon\}}]}$
+$\leq$ $\varepsilon+\frac{\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}\bm{X}x}\mathbb{I}_{\{\bm{X}>\varepsilon\}}]}{p_{\varepsilon}e^{-\frac{1}{2}\varepsilon x}}=\varepsilon+\frac{\tilde{\mathbb{E}}[\bm{X}e^{-\frac{1}{2}(\bm{X}-\varepsilon)x}\mathbb{I}_{\{\bm{X}>\varepsilon\}}]}{p_{\varepsilon}}.$
+Note that for any $x\geq 1$, $\bm{X}e^{-\frac{1}{2}(\bm{X}-\varepsilon)x}\mathbb{I}_{\{\bm{X}>\varepsilon\}}$ is dominated by $\bm{X}e^{-\frac{1}{2}(\bm{X}-\varepsilon)}\mathbb{I}_{\{\bm{X}>\varepsilon\}}\in L^{\infty}$.
 Hence, by the dominated convergence theorem, $\limsup\limits_{x\to\infty}Q(x)\leq\varepsilon$.
@@
 $\left|\dfrac{\mathrm{d}\left(\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}_{n}x}]\right)}{\mathrm{d}x}\right|=\left|-\frac{1}{2}\tilde{\mathbb{E}}[\bm{R}_{n}e^{-\frac{1}{2}\bm{R}_{n}x}]\right|\leq\frac{1}{2}\tilde{\mathbb{E}}[\bm{R}_{n}]\quad\text{for all}\,\,x\geq 0.$
-Because $\tilde{\mathbb{E}}[\bm{R}_{n}]\to\tilde{\mathbb{E}}[\bm{R}]$, the sequence $\{\tilde{\mathbb{E}}[\bm{R}_{n}]\}_{n\geq 1}$ is bounded.
+The sequence $\{\tilde{\mathbb{E}}[\bm{R}_{n}]\}_{n\geq 1}$ is bounded because $\tilde{\mathbb{E}}[\bm{R}_{n}]\to\tilde{\mathbb{E}}[\bm{R}]$.
 Hence, the family of functions $\{\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}_{n}x}]\}_{n\geq 1}$ is equicontinuous on $\mathcal{K}$.
-Moreover, because $\bm{R}_{n}\xrightarrow{d}\bm{R}$ and the function $r\mapsto e^{-\frac{1}{2}rx}$ is bounded and continuous, we have $\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}_{n}x}]\to\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}x}]$ for any $x\geq 0$.
-It follows that $\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}_{n}x}]\to\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}x}]$ uniformly on $\mathcal{K}$; see, e.g., rudin1976principles.
+Moreover, we have $\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}_{n}x}]\to\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}x}]$ for any $x\geq 0$, because $\bm{R}_{n}\xrightarrow{d}\bm{R}$ and the function $r\mapsto e^{-\frac{1}{2}rx}$ is bounded and continuous.
+It follows that $\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}_{n}x}]\to\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}x}]$ uniformly on $\mathcal{K}$; see, e.g., Rudin (1976, Exercise 7.16).
 Now we show the uniform convergence of $\{\tilde{\mathbb{E}}[\bm{R}_{n}e^{-\frac{1}{2}\bm{R}_{n}x}]\}_{n\geq 1}$ on $\mathcal{K}$.
@@
 Then $f_{n}^{M}$ is continuously differentiable and its derivative is uniformly bounded by $\frac{1}{2}M^{2}$, which implies that the family $\{f_{n}^{M}\}_{n\geq 1}$ is equicontinuous on $\mathcal{K}$.
-Furthermore, because $\bm{R}_{n}\xrightarrow{d}\bm{R}$, we have the pointwise convergence $f_{n}^{M}(x)\to f^{M}(x)\triangleq\tilde{\mathbb{E}}[\bm{R}e^{-\frac{1}{2}\bm{R}x}\mathbb{I}_{\{\bm{R}\leq M\}}]$; see, e.g., follmer2016stochastic.
+Furthermore, because $\bm{R}_{n}\xrightarrow{d}\bm{R}$, we have the pointwise convergence $f_{n}^{M}(x)\to f^{M}(x)\triangleq\tilde{\mathbb{E}}[\bm{R}e^{-\frac{1}{2}\bm{R}x}\mathbb{I}_{\{\bm{R}\leq M\}}]$; see, e.g., Föllmer and Schied (2016, Theorem A.43).
 As a result, $f_{n}^{M}$ converges uniformly to $f^{M}$ on $\mathcal{K}$.
@@
 Observe that
-$|\tilde{\mathbb{E}}[\bm{R}_{n}e^{-\frac{1}{2}\bm{R}_{n}x}]-\tilde{\mathbb{E}}[\bm{R}e^{-\frac{1}{2}\bm{R}x}]|\leq|f_{n}^{M}(x)-f^{M}(x)|+\tilde{\mathbb{E}}[\bm{R}_{n}\mathbb{I}_{\{\bm{R}_{n}>M\}}]+\tilde{\mathbb{E}}[\bm{R}\mathbb{I}_{\{\bm{R}>M\}}].$
+$\left|\tilde{\mathbb{E}}[\bm{R}_{n}e^{-\frac{1}{2}\bm{R}_{n}x}]-\tilde{\mathbb{E}}[\bm{R}e^{-\frac{1}{2}\bm{R}x}]\right|\leq|f_{n}^{M}(x)-f^{M}(x)|+\tilde{\mathbb{E}}[\bm{R}_{n}\mathbb{I}_{\{\bm{R}_{n}>M\}}]+\tilde{\mathbb{E}}[\bm{R}\mathbb{I}_{\{\bm{R}>M\}}].$
 Therefore, for $n$ large enough,
@@
 which implies the uniform convergence of $\{\tilde{\mathbb{E}}[\bm{R}_{n}e^{-\frac{1}{2}\bm{R}_{n}x}]\}_{n\geq 1}$ on $\mathcal{K}$.
-Finally, because $\tilde{\mathbb{E}}[\bm{R}]>0$, the function $\tilde{\mathbb{E}}[\bm{R}e^{-\frac{1}{2}\bm{R}x}]$ is bounded away from zero on $\mathcal{K}$.
+Finally, the function $\tilde{\mathbb{E}}[\bm{R}e^{-\frac{1}{2}\bm{R}x}]$ is bounded away from zero on $\mathcal{K}$, given that $\tilde{\mathbb{E}}[\bm{R}]>0$.
 Therefore, $\{h_{\bm{R}_{n}}\}_{n\geq 1}$ converges uniformly to $h_{\bm{R}}$ on $\mathcal{K}$.
 ∎
-The reverse hazard rate order can be characterized via expectation ratios and connected to first-order stochastic dominance under weighted measures, as established by caperaa1988tail (see also SS2007) and wang2024weighted.
-Lemma A.4 (caperaa1988tail).
-Let $\bm{R_{1}}$ and $\bm{R_{2}}$ be two random variables. Then $R_{1}\succeq_{rh}R_{2}$ if and only if
-$\frac{\tilde{\mathbb{E}}\left[u(\bm{R_{1}})w(\bm{R_{1}})\right]}{\tilde{\mathbb{E}}[w(\bm{R_{1}})]}\geq\frac{\tilde{\mathbb{E}}[u(\bm{R_{2}})w(\bm{R_{2}})]}{\tilde{\mathbb{E}}[w(\bm{R_{2}})]},$
-for all functions $u$ and $w$ for which the expectations exist, such that $w$ is positive, continuous, and decreasing, and $u$ is continuous and strictly increasing.
-Lemma A.5 (wang2024weighted).
-Let $\bm{R_{1}}$ and $\bm{R_{2}}$ be two random variables with cumulative distribution functions $F_{\bm{R_{1}}}$ and $F_{\bm{R_{2}}}$, respectively.
-Assume that $\bm{R_{1}}\succeq_{rh}\bm{R_{2}}$.
-Let $w(\cdot)$ be a positive, continuous, and strictly decreasing function.
-Define the weighted probability measures $\tilde{F}_{i}$ associated with $\bm{R}_{i}$ for $i=1,2$ by
-$\mathrm{d}\tilde{F}_{i}(x)=\frac{w(x)\mathrm{d}F_{\bm{R_{i}}}}{\tilde{\mathbb{E}}\left[w(\bm{R_{i}})\right]}.$ (A.3)
-Then, $\bm{R}_{1}$ dominates $\bm{R}_{2}$ under the weighted measures in the sense of first-order stochastic dominance, i.e.,
-$\tilde{F}_{1}\succeq_{1}\tilde{F}_{2}.$
-[…]
+The reverse hazard rate order can be characterized via expectation ratios, as established by Capéraà (1988) (see also Shaked and Shanthikumar (2007, Theorem 1.B.50)).
 Lemma A.4.
-For fixed $x>0$, $h(x)$ is strictly decreasing with respect to $\bm{R}$ in the sense of the reverse-hazard-rate order among distinct distributions.
+Let $\bm{R}_{1}$} and $\bm{R}_{2}$} be two random variables. Then ${\bm{R}_{1}}\succeq_{rh}{\bm{R}_{2}}$ if and only if
+$\frac{\tilde{\mathbb{E}}\left[u({\bm{R}_{1}})w({\bm{R}_{1}})\right]}{\tilde{\mathbb{E}}[w({\bm{R}_{1}})]}\geq\frac{\tilde{\mathbb{E}}[u({\bm{R}_{2}})w({\bm{R}_{2}})]}{\tilde{\mathbb{E}}[w({\bm{R}_{2}})]}$
+for all functions $u$ and $w$ such that the expectations exist, $w$ is positive and decreasing, and $u$ is increasing.
+The next lemma shows that $h_{\bm{R}}$ is “decreasing” w.r.t. $\bm{R}$ (under the reverse hazard rate order of $\bm{R}$).
+Lemma A.5.
+Let $\bm{R}_{1}$ and $\bm{R}_{2}$ be two non-negative random variables with $F_{{\bm{R}_{1}}}\neq F_{{\bm{R}_{2}}}$.
+If $\bm{R}_{1}\succeq_{rh}\bm{R}_{2}$, then $h_{1}(x)<h_{2}(x)$ for all $x\geq 0$, where $h_{i}=h_{\bm{R}_{i}}$ ($i=1,2$).
 Proof.
-Fix $x>0$.
-Let $\bm{R_{1}}$ and $\bm{R_{2}}$} be two non-negative random variables such that $\bm{R_{1}}\succeq_{rh}\bm{R_{2}}$ and their distributions are distinct.
-For notational simplicity, in the sequel, let $h_{i}(\cdot)$ and $H_{i}(\cdot)$ denote the function $h(\cdot)$ and the function defined in Theorem 3.2 corresponding to the random variable $\bm{R}_{i}$ for $i=1,2$, respectively.
-According to Lemma A.4, let $u(\gamma)=\gamma$ and $w(\gamma)=\exp(-\frac{1}{2}\gamma x)$.
-Note that $u(\gamma)$ is a strictly increasing function, and for $x>0$, $w(\gamma)$ is a positive, continuous, and strictly decreasing function of $\gamma$.
-Substituting them into this inequality, we have
-$\frac{\tilde{\mathbb{E}}[\bm{R_{1}}\exp(-\frac{1}{2}\bm{R_{1}}x)]}{\tilde{\mathbb{E}}[\exp(-\frac{1}{2}\bm{R_{1}}x)]}\geq\frac{\tilde{\mathbb{E}}[\bm{R_{2}}\exp(-\frac{1}{2}\bm{R_{2}}x)]}{\tilde{\mathbb{E}}[\exp(-\frac{1}{2}\bm{R_{2}}x)]}.$
-Moreover, assume that the equality holds.
-Define the weighted probability measures $\tilde{F}_{i}$ for $i=1,2$ by (A.3).
-The given condition is equivalent to $\tilde{\mathbb{E}}_{\tilde{F}_{1}}[\bm{R_{1}}]=\tilde{\mathbb{E}}_{\tilde{F}_{2}}[\bm{R_{1}}]$.
-From Lemma A.5, we have $\tilde{F}_{1}\succeq_{1}\tilde{F}_{2}$ and consequently, $\mathrm{d}\tilde{F}_{1}(\gamma)=\mathrm{d}\tilde{F}_{2}(\gamma)$.
-Given that $w(\gamma)$ is positive, this implies $\frac{\mathrm{d}\tilde{F}_{1}}{\tilde{\mathbb{E}}\left[w(\bm{R_{1}})\right]}=\frac{\mathrm{d}\tilde{F}_{2}}{\tilde{\mathbb{E}}\left[w(\bm{R_{2}})\right]}$.
-Integrating both sides over the entire support and using the property $\int\mathrm{d}\tilde{F}_{i}(\gamma)=1$, we obtain $1=\frac{\tilde{\mathbb{E}}\left[w(\bm{R_{1}})\right]}{\tilde{\mathbb{E}}\left[w(\bm{R_{2}})\right]}$, which implies $\tilde{\mathbb{E}}\left[w(\bm{R_{1}})\right]=\tilde{\mathbb{E}}\left[w(\bm{R_{2}})\right]$.
-Substituting this back yields $\mathrm{d}F_{\bm{R_{1}}}(\gamma)=\mathrm{d}F_{\bm{R_{2}}}(\gamma)$, confirming that $F_{\bm{R_{1}}}(\gamma)=F_{\bm{R_{2}}}(\gamma)$.
-This contradicts the different distributions of $\bm{R_{1}}$ and $\bm{R_{2}}$.
-Thus, we have
-$\frac{\tilde{\mathbb{E}}[\bm{R_{1}}\exp(-\frac{1}{2}\bm{R_{1}}x)]}{\tilde{\mathbb{E}}[\exp(-\frac{1}{2}\bm{R_{1}}x)]}>\frac{\tilde{\mathbb{E}}[\bm{R_{2}}\exp(-\frac{1}{2}\bm{R_{2}}x)]}{\tilde{\mathbb{E}}[\exp(-\frac{1}{2}\bm{R_{2}}x)]},$
-and equivalently, $h_{1}(x)<h_{2}(x)$. ∎
+Fix $x\geq 0$. Let $w(\gamma)=\exp(-\frac{1}{2}\gamma x)$ and $u(\gamma)=\gamma$. Applying Lemma A.4 yields
+$\frac{\tilde{\mathbb{E}}[{\bm{R}_{1}}w({\bm{R}_{1}})]}{\tilde{\mathbb{E}}[w({\bm{R}_{1}})]}\geq\frac{\tilde{\mathbb{E}}[{\bm{R}_{2}}w({\bm{R}_{2}})]}{\tilde{\mathbb{E}}[w({\bm{R}_{2}})]},$ (A.3)
+which implies that $h_{1}(x)\leq h_{2}(x)$.
+To show the strict inequality, suppose on the contrary that the equality holds in (A.3).
+Let $\tilde{F}_{i}$ ($i=1,2$) be the probability distribution functions defined by $\mathrm{d}\tilde{F}_{i}(\gamma)=\frac{w(\gamma)\mathrm{d}F_{\bm{R}_{i}}(\gamma)}{\tilde{\mathbb{E}}[w(\bm{R}_{i})]}$.
+Then $\int_{0}^{\infty}\gamma\mathrm{d}\tilde{F}_{1}(\gamma)=\int_{0}^{\infty}\gamma\mathrm{d}\tilde{F}_{2}(\gamma)$.
+On the other hand, a combination of $\bm{R}_{1}\succeq_{rh}\bm{R}_{2}$, Lemma A.4, and the definition of $\tilde{F}_{i}$ yields $\tilde{F}_{1}\succeq_{1}\tilde{F}_{2}$.
+Therefore, $\tilde{F}_{1}=\tilde{F}_{2}$.
+By $w(\cdot)>0$, we have $F_{{\bm{R}_{1}}}=F_{{\bm{R}_{2}}}$, which contradicts the assumption that $F_{{\bm{R}_{1}}}\neq F_{{\bm{R}_{2}}}$.
+Thus, the inequality must be strict.
+∎
 Appendix B Proofs
 B.1 Proof of Proposition 3.8
-Assume that the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly increasing and $\Lambda_{n}\nearrow H(\infty)$.
-Because the convergence is strictly increasing toward $H(\infty)$, we have $\Lambda_{n}<H(\infty)$ for all $n\geq 1$, ensuring the existence of the equilibrium.
-Thus, for each $n$, $v_{n}(0)$ is well-defined and uniquely determined by $v_{n}(0)=H^{-1}(\Lambda_{n})$.
+Assume that the sequence $\{\Lambda_{n}\}_{n\geq 1}$ strictly increases to $H(\infty)$.
+First, we have that $\Lambda_{n}<H(\infty)$ for all $n\geq 1$, which guarantees the existence of the equilibrium.
+Then, for each $n$, $v_{n}(0)$ is well-defined and uniquely determined by $v_{n}(0)=H^{-1}(\Lambda_{n})$.
 Consequently, the convergence $\Lambda_{n}\to H(\infty)$ implies that $\lim_{n\to\infty}v_{n}(0)=\infty$.
-By Theorem 3.2, we have $a_{n}(0)=h(v_{n}(0))\lambda_{0}$.
-Using the asymptotic property of $h$ established in Lemma A.2, namely $\lim_{x\to\infty}h(x)=\frac{1}{r_{0}}$, we obtain
+Based on Theorem 3.2, we have $a_{n}(0)=h(v_{n}(0))\lambda_{0}$.
+Using the asymptotic property of $h$ established in Lemma A.2, i.e., $\lim_{x\to\infty}h(x)=\frac{1}{r_{0}}$, we obtain
 $\lim_{n\to\infty}a_{n}(0)=\lambda_{0}\lim_{v_{n}(0)\to\infty}h(v_{n}(0))=\frac{\lambda_{0}}{r_{0}}.$
-Furthermore, because $\{\Lambda_{n}\}$ is strictly increasing, $\{v_{n}(0)\}$ is strictly increasing.
-Combined with the monotonicity of $h$ (see Lemma A.1), it follows that $|a_{n}(0)|=|\lambda_{0}|h(v_{n}(0))$ converges monotonically increasing to $\frac{|\lambda_{0}|}{r_{0}}$.
-Assume that the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly decreasing and $\Lambda_{n}\searrow 0$.
-Because $H(\infty)>0$, and $\Lambda_{n}\searrow 0$, there exists an integer $N$ such that for all $n>N$, $\Lambda_{n}<H(\infty)$.
-Thus, the equilibrium exists for all sufficiently large $n$.
+Furthermore, $\{v_{n}(0)\}$ is strictly increasing as $\{\Lambda_{n}\}$ is strictly increasing.
+Combined with the monotonicity of $h$ (see Lemma A.1), it follows that $|a_{n}(0)|=|\lambda_{0}|h(v_{n}(0))$ increases to $\frac{|\lambda_{0}|}{r_{0}}$.
+Assume that the sequence $\{\Lambda_{n}\}_{n\geq 1}$ strictly decreases to 0.
+By $H(\infty)>0$, $\Lambda_{n}<H(\infty)$ for all sufficiently large $n$.
+Then the equilibrium exists for all sufficiently large $n$.
 For such $n$, we have $v_{n}(0)=H^{-1}(\Lambda_{n})$.
@@
 $\lim_{n\to\infty}v_{n}(0)=H^{-1}(0)=0.$
-Recall that $h(0)=\frac{1}{\tilde{\mathbb{E}}[\bm{R}]}$. Taking the limit of $a_{n}(0)$, we have
+Recall that $h(0)=\frac{1}{\tilde{\mathbb{E}}[\bm{R}]}$. Taking the limit of $a_{n}(0)$ yields
 $\lim_{n\to\infty}a_{n}(0)=\lambda_{0}\lim_{v_{n}(0)\to 0}h(v_{n}(0))=\frac{\lambda_{0}}{\tilde{\mathbb{E}}[\bm{R}]}.$
-Finally, because $\{\Lambda_{n}\}$ is strictly decreasing, $\{v_{n}(0)\}$ is strictly decreasing.
-The strict monotonicity of $h$ implies that $|a_{n}(0)|$ converges monotonically decreasing to $\frac{|\lambda_{0}|}{\tilde{\mathbb{E}}[\bm{R}]}$.
+Finally, $\{v_{n}(0)\}$ is strictly decreasing as $\{\Lambda_{n}\}$ is strictly decreasing.
+The strict monotonicity of $h$ implies that $|a_{n}(0)|$ decreases to $\frac{|\lambda_{0}|}{\tilde{\mathbb{E}}[\bm{R}]}$.
 B.2 Proof of Theorem 3.11
@@
 Because the limit function $h$ is strictly positive, it follows that $\{H_{n}\}_{n\geq 1}$ converges uniformly to $H$ on every compact subset of $0,\infty)$ as well.
-By Lemma B.1 below, the sequence of the inverse functions $\{H_{n}^{-1}\}_{n\geq 1}$ converge uniformly to $H^{-1}$ on $0,\Lambda(0)]$.
+By Lemma B.1 below, the sequence of the inverse functions $\{H_{n}^{-1}\}_{n\geq 1}$ converges uniformly to $H^{-1}$ on $0,\Lambda(0)]$.
 We now see the uniform convergence of $\{v_{n}\}_{n\geq 1}$ to $v$ on $[0,T]$:
@@
 Therefore, the uniform convergence of $\{h_{n}\}_{n\geq 1}$ on $\mathcal{K}$ implies $\sup_{t\in[0,T]}|h_{n}(v_{n}(t))-h(v_{n}(t))|\to 0$.
-Moreover, $h$ is continuous and therefore uniformly continuous on the compact set $\mathcal{K}$, which combined with the fact that $\sup_{t\in[0,T]}|v_{n}(t)-v(t)|\to 0$ implies $\sup_{t\in[0,T]}|h(v_{n}(t))-h(v(t))|\to 0$.
+Moreover, $h$ is continuous and therefore uniformly continuous on the compact set $\mathcal{K}$, which combined with the fact that $\sup\limits_{t\in[0,T]}|v_{n}(t)-v(t)|\to 0$ implies $\sup\limits_{t\in[0,T]}|h(v_{n}(t))-h(v(t))|\to 0$.
 Consequently, by (B.1), we have
 $\lim_{n\to\infty}\sup_{t\in[0,T]}|h_{n}(v_{n}(t))-h(v(t))|=0.$ (B.2)
-Finally, recalling that $a_{n}(t)=h_{n}(v_{n}(t))\lambda(t)$ and $a(t)=h(v(t))\lambda(t)$, we get from (B.2) that $\lim_{n\to\infty}a_{n}(t)=a(t)$ for every $t\in[0,T)$.
-Furthermore, if $\lambda\in L^{\infty}([0,T])$, then the above convergence is uniform on $[0,T)$.
+Finally, recalling that $a_{n}(t)=h_{n}(v_{n}(t))\lambda(t)$ and $a(t)=h(v(t))\lambda(t)$, we get from (B.2) that $\lim\limits_{n\to\infty}a_{n}(t)=a(t)$ for every $t\in[0,T)$.
+Furthermore, if $\lambda\in L^{\infty}([0,T))$, then the above convergence is uniform on $[0,T)$.
 Lemma B.1.
@@
 Next, because $H^{-1}$ is continuous on the compact interval $[0,H(M)]$, it is uniformly continuous.
-Thus, for any $\varepsilon>0$, there exists $\delta>0$ such that for any $x^{\prime},x^{\prime\prime}\in[0,H(M)]$ with $|x^{\prime}-x^{\prime\prime}|<\delta$, we have
+Thus, for any $\varepsilon>0$, there exists a $\delta>0$ such that for any $x^{\prime},x^{\prime\prime}\in[0,H(M)]$ with $|x^{\prime}-x^{\prime\prime}|<\delta$, we have
 $|H^{-1}(x^{\prime})-H^{-1}(x^{\prime\prime})|<\varepsilon.$ (B.3)
-By uniform convergence of $H_{n}\to H$ on $[0,M]$, there exists $N\geq N_{0}$ such that for all $n\geq N$ and all $y\in[0,M]$,
+By uniform convergence of $H_{n}\to H$ on $[0,M]$, there exists a $N\geq N_{0}$ such that for all $n\geq N$ and all $y\in[0,M]$,
 $|H_{n}(y)-H(y)|<\delta.$
@@
 $|H(x_{n})-x|=|H(x_{n})-H_{n}(x_{n})|<\delta.$
-Because $x,H(x_{n})\in[0,H(M)]$ and $x_{n}=H^{-1}(H(x_{n}))$, inequality (B.3) yields
+Because $x,H(x_{n})\in[0,H(M)]$ and $x_{n}=H^{-1}(H(x_{n}))$, the inequality (B.3) yields
 $|H_{n}^{-1}(x)-H^{-1}(x)|=|x_{n}-H^{-1}(x)|=|H^{-1}(H(x_{n}))-H^{-1}(x)|<\varepsilon.$
@@
 Lemma B.2.
-Let $\bm{R}_{1}$ and $\bm{R}_{2}$ be two non-negative random variables satisfying the conditions (1) and (2) in Proposition 5.3.
-Recall the definition in Proposition 3.8 that $\bm{X}_{i}=\bm{R}_{i}-r_{0}$ and $Q_{i}(x)=\frac{\tilde{\mathbb{E}}[\bm{X}_{i}e^{-\frac{1}{2}\bm{X}_{i}x}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}_{i}x}]}$, $x\geq 0$.
-Then, the following assertions hold:
+Let $\bm{R}_{1}$ and $\bm{R}_{2}$ be two non-negative random variables satisfying Conditions (c1) and (c2) in Proposition 5.3. Following the notation in the proof of Lemma A.2, let
+$\bm{X}_{i}=\bm{R}_{i}-r_{0}\text{ and }Q_{i}(x)=\frac{\tilde{\mathbb{E}}[\bm{X}_{i}e^{-\frac{1}{2}\bm{X}_{i}x}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}_{i}x}]},\quad x\geq 0.$ (B.4)
+Then we have the following assertions.
   1. (1)
-    For each $i\in\{1,2\}$, there exist constant $M_{i}>0$, such that $Q_{i}(x)\leq M_{i}e^{-\frac{1}{2}\delta_{i}x}$ for any $x\geq 0$. In particular, $Q_{i}(x)\to 0$ as $x\to\infty$.
+    For each $i\in\{1,2\}$, there exists a constant $M_{i}>0$ such that $Q_{i}(x)\leq M_{i}e^{-\frac{1}{2}\delta_{i}x}$ for $x\geq 0$. In particular, $Q_{i}(x)\to 0$ as $x\to\infty$.
   2. (2)
-    For each $i\in\{1,2\}$, the integral $\int_{0}^{\infty}(h_{i}^{-2}(s)-r_{0}^{2})ds$ converges to a finite constant $K_{i}$.
+    For each $i\in\{1,2\}$, the integral $K_{i}=\displaystyle\int_{0}^{\infty}(h_{i}^{-2}(s)-r_{0}^{2})\mathrm{d}s$ is finite.
   3. (3)
-    For any $x_{1},x_{2}\geq 0$ and any $\varepsilon>0$, it holds that $\dfrac{Q_{1}(x_{1})}{Q_{2}(x_{2})}\leq C\dfrac{e^{-\frac{1}{2}\delta_{1}x_{1}}}{e^{-\frac{1}{2}(\delta_{2}+\varepsilon)x_{2}}}$, where $C$ is a constant depending on $\varepsilon$.
+    For any $\varepsilon>0$, there exists a constant $C>0$ such that $\dfrac{Q_{1}(x_{1})}{Q_{2}(x_{2})}\leq C\dfrac{e^{-\frac{1}{2}\delta_{1}x_{1}}}{e^{-\frac{1}{2}(\delta_{2}+\varepsilon)x_{2}}}$ for any $x_{1},x_{2}\geq 0$.
 Proof.
 For $i\in\{1,2\}$.
-By condition (1) in Proposition 5.3, we have $\mathbb{P}\mathbb{(}\bm{X}_{i}=0)=p_{i}$.
-The Dominated Convergence Theorem then implies $\lim_{x\to\infty}\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}_{i}x}]=p_{i}\in(0,1)$.
-Hence, there exists $m_{i}>0$ such that $\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}_{i}x}]\geq m_{0}$ for any $x\geq 0$.
-Moreover, by condition (2) in Proposition 5.3, the support of $\bm{X}_{i}$ on $(0,\infty)$ is contained in $\delta_{i},\infty)$.
-Therefore,
+By Condition (c1) in Proposition 5.3, we have $\tilde{\mathbb{P}}\mathbb{(}\bm{X}_{i}=0)=p_{i}$.
+The dominated convergence theorem then implies $\lim\limits_{x\to\infty}\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}_{i}x}]=p_{i}\in(0,1)$.
+Hence, there exists $m_{i}>0$ such that $\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}_{i}x}]\geq m_{i}$ for any $x\geq 0$.
+Moreover, by Condition (c2) in Proposition 5.3, the support of $\bm{X}_{i}$ on $(0,\infty)$ is contained in $\delta_{i},\infty)$.
+As such,
 $\tilde{\mathbb{E}}[\bm{X}_{i}e^{-\frac{1}{2}\bm{X}_{i}x}]=\int_{\delta_{i}}^{\infty}\gamma e^{-\frac{1}{2}\gamma x}\mathrm{d}F_{\bm{X}_{i}}(\gamma)\leq e^{-\frac{1}{2}\delta_{i}x}\int_{\delta_{i}}^{\infty}\gamma\mathrm{d}F_{\bm{X}_{i}}(\gamma)=e^{-\frac{1}{2}\delta_{i}x}\tilde{\mathbb{E}}[\bm{X}_{i}\mathbb{I}_{\{\bm{X}_{i}\geq\delta_{i}\}}].$
-The desired conclusion then follows by taking $M_{i}\triangleq\frac{\tilde{\mathbb{E}}[\bm{X}_{i}\mathbb{I}_{\{\bm{X}_{i}\geq\delta_{i}\}}]}{m_{i}}$.
+Thus the desired conclusion follows by taking $M_{i}\triangleq\frac{\tilde{\mathbb{E}}[\bm{X}_{i}\mathbb{I}_{\{\bm{X}_{i}\geq\delta_{i}\}}]}{m_{i}}$.
 For $i\in\{1,2\}$, based on (A.2), the conclusion follows immediately from Assertion (1) and the fact that $\delta_{i}>0$.
@@
 $\tilde{\mathbb{E}}[\bm{X_{2}}e^{-\frac{1}{2}\bm{X_{2}}x}]\geq\int_{\delta_{2}}^{\delta_{2}+\varepsilon}xe^{-\frac{1}{2}x\gamma}\mathrm{d}F_{\bm{X_{2}}}(\gamma)\geq\delta_{2}e^{-\frac{1}{2}(\delta_{2}+\varepsilon)x}p_{\varepsilon,2}.$
-Combining this with assertion (1) gives the desired bound with $C\triangleq\frac{M_{1}}{\delta_{2}p_{\varepsilon,2}}$. ∎
-Now we return to prove Proposition B.3.
-By assertion (1) of Lemma B.2 and (A.2), we have $\lim\limits_{x\to\infty}h_{i}(x)=\frac{1}{r_{0}}$.
-Because $r_{0}>0$, it follows that the integral $\int_{0}^{\infty}h_{i}^{-2}(x)\mathrm{d}x$ diverges, which implies $H_{i}(\infty)=\infty$.
-Consequently, Theorem 3.2 (1) ensures the existence of a unique equilibrium $\pi_{i}$ for $\bm{R}_{i}$ for any $T>0$.
-Based on the relation ${\pi}_{i}=(\sigma^{\top})^{-1}a_{i}$, it suffices to compare $a_{i}$.
-By (2.5), we have $|a_{i}(0,T)|=h_{i}(v_{i}(0,T))|\lambda(0)|$.
-First, we establish the asymptotic relationship between $v_{1}(0,T)$ and $v_{2}(0,T)$ as $T\to\infty$.
-By Theorem 3.2 (1), we have $v_{i}(0,T)=H_{i}^{-1}(\Lambda(0,T))$.
-It then follows that $\lim_{T\to\infty}v_{i}(0,T)=\infty$, because $H_{i}(\infty)=\infty$ and $\lim_{T\to\infty}\Lambda(0,T)=\infty$.
-Moreover, the relation $v_{i}(0,T)=H_{i}^{-1}(\Lambda(0,T))$ yields
-$\Lambda(0,T)=\int_{0}^{v_{i}(0,T)}(r_{0}^{2}+(h_{i}^{-2}(x)-r_{0}^{2}))\mathrm{d}x=r_{0}^{2}v_{i}(0,T)+\int_{0}^{v_{i}(0,T)}(h_{i}^{-2}(x)-r_{0}^{2})\mathrm{d}x.$ (B.4)
-Define the tail integral $R_{i}(y)\triangleq\int_{y}^{\infty}(h_{i}^{-2}(x)-r_{0}^{2})\mathrm{d}x$, which satisfies $\lim_{y\to\infty}R_{i}(y)=0$ by Lemma B.2 (2). Then (B.4) can be rewritten as
-$\Lambda(0,T)=r_{0}^{2}v_{i}(0,T)+K_{i}-R_{i}(v_{i}(0,T)).$ (B.5)
+Combining this with Assertion (1) gives the desired bound with $C\triangleq\frac{M_{1}}{\delta_{2}p_{\varepsilon,2}}$. ∎ Now we return to prove Proposition 5.3.
+By Assertion (1) of Lemma B.2 and (A.2), we have $\lim_{x\to\infty}h_{i}(x)=\frac{1}{r_{0}}$.
+Because $r_{0}>0$, it follows that $\int_{0}^{\infty}h_{i}^{-2}(x)\mathrm{d}x=\infty$, which implies $H_{i}(\infty)=\infty$.
+Consequently, Theorem 3.2(1) guarantees the existence and uniqueness of the deterministic equilibrium strategy corresponding to $\bm{R}_{i}$ for any deterministic and right-continuous market price of risk $\lambda(\cdot)$ in $L^{2}([0,T))$.
+First, we establish the asymptotic relationship between the inverse functions $H_{1}^{-1}(\Lambda)$ and $H_{2}^{-1}(\Lambda)$ for large scalar $\Lambda$.
+Because $H_{i}(\infty)=\infty$, it follows that $H_{i}^{-1}(\Lambda)\to\infty$ as $\Lambda\to\infty$.
+Moreover, the identity $\Lambda=H_{i}(H_{i}^{-1}(\Lambda))$ yields
+$\Lambda=\int_{0}^{H_{i}^{-1}(\Lambda)}(r_{0}^{2}+(h_{i}^{-2}(x)-r_{0}^{2}))\mathrm{d}x=r_{0}^{2}H_{i}^{-1}(\Lambda)+\int_{0}^{H_{i}^{-1}(\Lambda)}(h_{i}^{-2}(x)-r_{0}^{2})\mathrm{d}x.$ (B.5)
+Define the tail integral $\Theta_{i}(u)\triangleq\int_{u}^{\infty}(h_{i}^{-2}(x)-r_{0}^{2})\mathrm{d}x$, which satisfies $\lim_{u\to\infty}\Theta_{i}(u)=0$ based on Lemma B.2(2).
+Then (B.5) can be rewritten as
+$\Lambda=r_{0}^{2}H_{i}^{-1}(\Lambda)+K_{i}-\Theta_{i}(H_{i}^{-1}(\Lambda)).$ (B.6)
 Equating (B.6) for $i=1$ and $i=2$ gives
-$v_{1}(0,T)=v_{2}(0,T)+\dfrac{K_{2}-K_{1}}{r_{0}^{2}}+\dfrac{R_{2}(v_{2}(0,T))-R_{1}(v_{1}(0,T))}{r_{0}^{2}}.$
-Let $\Delta K=\dfrac{K_{2}-K_{1}}{r_{0}^{2}}$ and $\xi(T)=\dfrac{R_{2}(v_{2}(0,T))-R_{1}(v_{1}(0,T))}{r_{0}^{2}}$.
-Because $v_{i}(0,T)\to\infty$ as $T\to\infty$, the tail difference term $\xi(T)$ tends to 0.
-Consequently, we obtain the asymptotic relation $v_{1}(0,T)=v_{2}(0,T)+\Delta K+\xi(T)$.
-Second, we compare $h_{1}(v_{1}(0,T))$ and $h_{2}(v_{2}(0,T))$.
-Because $h_{i}(s)=(r_{0}+Q_{i}(s))^{-1}$, this reduces to comparing $Q_{1}(v_{1}(0,T))$ and $Q_{2}(v_{2}(0,T))$.
+$H_{1}^{-1}(\Lambda)=H_{2}^{-1}(\Lambda)+\dfrac{K_{2}-K_{1}}{r_{0}^{2}}+\dfrac{\Theta_{2}(H_{2}^{-1}(\Lambda))-\Theta_{1}(H_{1}^{-1}(\Lambda))}{r_{0}^{2}}.$
+Let $\Delta K\triangleq\dfrac{K_{2}-K_{1}}{r_{0}^{2}}$ and $\xi(\Lambda)\triangleq\dfrac{\Theta_{2}(H_{2}^{-1}(\Lambda))-\Theta_{1}(H_{1}^{-1}(\Lambda))}{r_{0}^{2}}$.
+Because $H_{i}^{-1}(\Lambda)\to\infty$ as $\Lambda\to\infty$, the tail difference term $\xi(\Lambda)$ tends to 0 as $\Lambda\to\infty$.
+Second, we compare $h_{1}(H_{1}^{-1}(\Lambda))$ and $h_{2}(H_{2}^{-1}(\Lambda))$.
+Because $h_{i}(s)=(r_{0}+Q_{i}(s))^{-1}$, this reduces to comparing $Q_{1}(H_{1}^{-1}(\Lambda))$ and $Q_{2}(H_{2}^{-1}(\Lambda))$.
 Let $\varepsilon>0$ be a constant satisfying $\delta_{1}-\delta_{2}-\varepsilon>0$.
 Using the bound established in Lemma B.2(3), we obtain
-$\dfrac{Q_{1}(v_{1}(0,T))}{Q_{2}(v_{2}(0,T))}\leq C\dfrac{e^{-\frac{1}{2}\delta_{1}v_{1}(0,T)}}{e^{-\frac{1}{2}(\delta_{2}+\varepsilon)v_{2}(0,T)}}.$
-Substituting $v_{1}(0,T)=v_{2}(0,T)+\Delta K+\xi(T)$ into the exponent yields
-$\dfrac{Q_{1}(v_{1}(0,T))}{Q_{2}(v_{2}(0,T))}\leq Ce^{-\frac{1}{2}\delta_{1}(\Delta K+\xi(T))}\cdot e^{-\frac{1}{2}(\delta_{1}-\delta_{2}-\varepsilon)v_{2}(0,T)}.$
-Because $\lim_{T\to\infty}\xi(T)=0$, the first exponential term converges to a finite positive constant.
-Moreover, as $v_{2}(0,T)\to\infty$, the second term tends to 0, and hence $\frac{Q_{1}(v_{1}(0,T))}{Q_{2}(v_{2}(0,T))}\to 0$ as $T\to\infty$.
-This implies for sufficiently large $T$, $Q_{1}(v_{1}(0,T))<Q_{2}(v_{2}(0,T))$.
-Because $h_{i}(x)=(r_{0}+Q_{i}(x))^{-1}$ is strictly decreasing in $Q_{i}$, we have $h_{1}(v_{1}(0,T))>h_{2}(v_{2}(0,T))$.
-Consequently, $|a_{1}(0,T)|>|a_{2}(0,T)|$ because $\lambda(0)\neq 0$, and equivalently, $|\pi_{1}(0,T)|>|\pi_{2}(0,T)|$.
+$\dfrac{Q_{1}(H_{1}^{-1}(\Lambda))}{Q_{2}(H_{2}^{-1}(\Lambda))}\leq C\dfrac{e^{-\frac{1}{2}\delta_{1}H_{1}^{-1}(\Lambda)}}{e^{-\frac{1}{2}(\delta_{2}+\varepsilon)H_{2}^{-1}(\Lambda)}}.$
+Substituting $H_{1}^{-1}(\Lambda)=H_{2}^{-1}(\Lambda)+\Delta K+\xi(\Lambda)$ into the exponent yields
+$\dfrac{Q_{1}(H_{1}^{-1}(\Lambda))}{Q_{2}(H_{2}^{-1}(\Lambda))}\leq Ce^{-\frac{1}{2}\delta_{1}(\Delta K+\xi(\Lambda))}\cdot e^{-\frac{1}{2}(\delta_{1}-\delta_{2}-\varepsilon)H_{2}^{-1}(\Lambda)}.$
+As $\Lambda\to\infty$, we have $\xi(\Lambda)\to 0$, so the first exponential term converges to a finite and positive constant.
+Moreover, $H_{2}^{-1}(\Lambda)\to\infty$, so the second term tends to 0.
+Hence, the ratio $\frac{Q_{1}(H_{1}^{-1}(\Lambda))}{Q_{2}(H_{2}^{-1}(\Lambda))}$ tends to 0 as $\Lambda\to\infty$.
+It follows that there exists a constant $\overline{\Lambda}>0$ such that, for all $\Lambda>\overline{\Lambda}$, we have $Q_{1}(H_{1}^{-1}(\Lambda))<Q_{2}(H_{2}^{-1}(\Lambda))$ and thus
+$h_{1}(H_{1}^{-1}(\Lambda))>h_{2}(H_{2}^{-1}(\Lambda))\quad\text{for all }\Lambda>\overline{\Lambda}.$ (B.7)
+Next, we compare the risk exposure magnitudes at time 0.
+For any market price of risk $\lambda(\cdot)$ satisfying $\lambda(0)\neq 0$ and $\Lambda(0)\triangleq\int_{0}^{T}|\lambda(t)|^{2}\,\mathrm{d}t>\overline{\Lambda}$.
+We have $v_{i}(0)=H_{i}^{-1}(\Lambda(0))$ according to Theorem 3.2(1).
+Applying (B.7) with $\Lambda=\Lambda(0)$, we obtain $h_{1}(v_{1}(0))>h_{2}(v_{2}(0))$ because $\Lambda(0)>\overline{\Lambda}$.
+Then, we conclude that $|a_{1}(0)|>|a_{2}(0)|$ because the risk exposure magnitude satisfies $|a_{i}(0)|=h_{i}(v_{i}(0))|\lambda(0)|$ by (2.5).
+Finally, because $v_{i}(T)=0$, we have $h_{i}(v_{i}(T))=\dfrac{1}{\tilde{\mathbb{E}}[R_{i}]}$.
+The assumption $\tilde{\mathbb{E}}[R_{1}]>\tilde{\mathbb{E}}[R_{2}]$ then implies $h_{1}(0)<h_{2}(0)$.
+Combined with $\lambda(T-)\neq 0$, we obtain $|a_{1}(T-)|<|a_{2}(T-)|$.
 B.4 Proof of Proposition 5.4
-For $i=1,2$, let $\mathbb{P}(\bm{R}_{i}=r_{0})=p_{i}$. The function $Q_{i}(\cdot)$ is explicitly given by
+Step 1:
+Fix a right-continuous and deterministic function $\lambda(\cdot)\in L^{2}([0,T))$.
+We show that there exists at most one $t^{*}\in(0,T)$ such that $|a_{1}(t^{*})|=|a_{2}(t^{*}|$.
+Indeed, let $t^{*}\in(0,T)$ be arbitrary and satisfy $|a_{1}(t^{*})|=|a_{2}(t^{*}|$.
+Let $\mathcal{D}(t)\triangleq Q_{1}(v_{1}(t))-Q_{2}(v_{2}(t))$.
+Then, $\mathcal{D}(t^{*})=0$.
+We now investigate the derivative of $\mathcal{D}(\cdot)$ at $t^{*}$.
+Recalling the expression for $Q_{i}(x)$ in (B.4), we have
 $Q_{i}(x)=\frac{\tilde{\mathbb{E}}[\bm{X}_{i}e^{-\bm{X}_{i}\frac{x}{2}}]}{\tilde{\mathbb{E}}[e^{-\bm{X}_{i}\frac{x}{2}}]}=\frac{\delta_{i}(1-p_{i})e^{-\delta_{i}\frac{x}{2}}}{p_{i}+(1-p_{i})e^{-\delta_{i}\frac{x}{2}}},\quad x\in[0,\infty).$
-We first verify the existence of a solution to $|a_{1}(t)|=|a_{2}(t)|$.
-This equality is equivalent to $Q_{1}(v_{1}(t))=Q_{2}(v_{2}(t))$, because $|a_{i}(t)|=|\lambda(t)|(r_{0}+Q_{i}(v_{i}(t)))^{-1}$.
-Define $\mathcal{D}(t)\triangleq Q_{1}(v_{1}(t))-Q_{2}(v_{2}(t))$.
-At $t=0$, by assumption, $\mathcal{D}(0)<0$.
-At $t=T$, because $v_{i}(T)=0$ and $Q_{i}(0)=\tilde{\mathbb{E}}[\bm{R}_{i}]-r_{0}$, the condition $\tilde{\mathbb{E}}[\bm{R}_{1}]>\tilde{\mathbb{E}}[\bm{R}_{2}]$ implies $Q_{1}(0)>Q_{2}(0)$, and therefore $\mathcal{D}(T)>0$.
-Because $\mathcal{D}(t)$ is continuous, there exists at least one $t^{*}\in(0,T)$ such that $\mathcal{D}(t^{*})=0$.
-Next, we show uniqueness by differentiating $\mathcal{D}(\cdot)$. To this end, we compute $Q_{i}^{\prime}(x)$:
+Note that
 $Q_{i}^{\prime}(x)$ $=\frac{-\frac{\delta_{i}^{2}}{2}(1-p_{i})e^{-\delta_{i}\frac{x}{2}}\left(p_{i}+(1-p_{i})e^{-\delta_{i}\frac{x}{2}}\right)-\delta_{i}(1-p_{i})e^{-\delta_{i}\frac{x}{2}}\left(-\frac{\delta_{i}}{2}(1-p_{i})e^{-\delta_{i}\frac{x}{2}}\right)}{\left(p_{i}+(1-p_{i})e^{-\delta_{i}\frac{x}{2}}\right)^{2}}$
@@
 Differentiating $\mathcal{D}(t)$ and using $v_{i}^{\prime}(t)=-|\lambda(t)|^{2}(r_{0}+Q_{i}(v_{i}(t)))^{-2}$ yields
-$\mathcal{D}^{\prime}(t)$ $=Q_{1}^{\prime}(v_{1})v_{1}^{\prime}-Q_{2}^{\prime}(v_{2})v_{2}^{\prime}$
-$=\frac{1}{2}Q_{1}(Q_{1}-\delta_{1})\left(\frac{-|\lambda(t)|^{2}}{(r_{0}+Q_{1})^{2}}\right)-\frac{1}{2}Q_{2}(Q_{2}-\delta_{2})\left(\frac{-|\lambda(t)|^{2}}{(r_{0}+Q_{2})^{2}}\right)$
-$=\frac{|\lambda(t)|^{2}}{2}\left[\frac{Q_{2}(Q_{2}-\delta_{2})}{(r_{0}+Q_{2})^{2}}-\frac{Q_{1}(Q_{1}-\delta_{1})}{(r_{0}+Q_{1})^{2}}\right].$
-At any solution $t^{*}$ where $\mathcal{D}(t^{*})=0$, we have $Q=Q_{1}(v_{1}(t^{*}))=Q_{2}(v_{2}(t^{*}))>0$. Substituting $Q_{1}=Q_{2}=Q$ yields
-$\mathcal{D}^{\prime}(t^{*})=\frac{|\lambda(t^{*})|^{2}Q}{2(r_{0}+Q)^{2}}(\delta_{1}-\delta_{2}).$
-Because $\delta_{1}>\delta_{2}$ and $\lambda(t^{*})\neq 0$, we have $\mathcal{D}^{\prime}(t^{*})>0$. Therefore, the solution $t^{*}$ is unique.
+$\mathcal{D}^{\prime}$ $=Q_{1}^{\prime}(v_{1})v_{1}^{\prime}-Q_{2}^{\prime}(v_{2})v_{2}^{\prime}$
+$=\frac{1}{2}Q_{1}(v_{1})(Q_{1}(v_{1})-\delta_{1})\left(\frac{-|\lambda|^{2}}{(r_{0}+Q_{1}(v_{1}))^{2}}\right)-\frac{1}{2}Q_{2}(v_{2})(Q_{2}(v_{2})-\delta_{2})\left(\frac{-|\lambda|^{2}}{(r_{0}+Q_{2}(v_{2}))^{2}}\right)$
+$=\frac{|\lambda|^{2}}{2}\left[\frac{Q_{2}(v_{2})(Q_{2}(v_{2})-\delta_{2})}{(r_{0}+Q_{2}(v_{2}))^{2}}-\frac{Q_{1}(v_{1})(Q_{1}(v_{1})-\delta_{1})}{(r_{0}+Q_{1}(v_{1}))^{2}}\right].$
+For any $t^{*}\in(0,T)$ such that $\mathcal{D}(t^{*})=0$, we have $Q\triangleq Q_{1}(v_{1}(t^{*}))=Q_{2}(v_{2}(t^{*}))>0$. Therefore,
+$\mathcal{D}^{\prime}(t^{*})=\frac{|\lambda(t^{*})|^{2}Q}{2(r_{0}+Q)^{2}}(\delta_{1}-\delta_{2})>0.$
+Thus, there exists a unique $t^{*}\in(0,T)$ such that $\mathcal{D}(t^{*})=0$.
+Step 2:
+Define
+$\mathcal{Y}(\Lambda)\triangleq h_{1}(H_{1}^{-1}(\Lambda))-h_{2}(H_{2}^{-1}(\Lambda)),\quad\Lambda\in[0,\infty).$ (B.8)
+We show there exists a unique $\widehat{\Lambda}\in(0,\infty)$ such that $\mathcal{Y}(\widehat{\Lambda})=0$.
+At $\Lambda=0$, we have $H_{i}^{-1}(0)=0$ and $h_{i}(0)=\dfrac{1}{\tilde{\mathbb{E}}[\bm{R}_{i}]}$.
+The assumption $\tilde{\mathbb{E}}[\bm{R}_{1}]>\tilde{\mathbb{E}}[\bm{R}_{2}]$ implies $h_{1}(0)<h_{2}(0)$, and thus $\mathcal{Y}(0)<0$.
+Conversely, as $\Lambda\to\infty$, (B.7) implies that $h_{1}(H_{1}^{-1}(\Lambda))>h_{2}(H_{2}^{-1}(\Lambda))$.
+Therefore, $\mathcal{Y}(\Lambda)>0$ for sufficiently large $\Lambda$.
+By the mean value theorem, there exists at least one $\widehat{\Lambda}\in(0,\infty)$ such that $\mathcal{Y}(\widehat{\Lambda})=0$.
+We next prove the uniqueness of $\widehat{\Lambda}$.
+Suppose, by contradiction, that there exist $0<\widehat{\Lambda}_{1}<\widehat{\Lambda}_{2}<\infty$ such that $\mathcal{Y}(\widehat{\Lambda}_{i})=0(i=1,2)$.
+Choose a constant market price of risk $\lambda(t)\equiv\lambda_{0}\neq 0$ such that $\Lambda(0)=|\lambda_{0}|^{2}T>\widehat{\Lambda}_{2}$.
+Then there exist $0<t_{1}<t_{2}<T$ such that $\Lambda(t_{i})=\widehat{\Lambda}_{i}$.
+Consequently, $\mathcal{D}(t_{1})=\mathcal{D}(t_{2})=0$, which contradicts the conclusion in Step 1.
+Therefore, $\widehat{\Lambda}$ is unique.
+Step 3:
+We now apply the conclusions in Steps 1 and 2 to prove the three assertions:
+[…]
+Next, assume $\int_{0}^{T}|\lambda(t)|^{2}\mathrm{d}t>\widehat{\Lambda}$.
+By Proposition 5.3, we have $|a_{1}(T-)|<|a_{2}(T-)|$ .
+Moreover, because $\Lambda(0)>\widehat{\Lambda}$, it follows that $\mathcal{Y}(\Lambda(0))>0$, and hence $|a_{1}(0)|>|a_{2}(0)|$.
+Combining this with the result of Step 1, we conclude that there exists a unique time $t^{*}\in(0,T)$ such that $|a_{1}(t^{*})|=|a_{2}(t^{*})|$.
+The remaining conclusions then follow immediately.
+Finally, assume $\int_{0}^{T}|\lambda(t)|^{2}\mathrm{d}t=\widehat{\Lambda}$.
+Then $\Lambda(0)=\widehat{\Lambda}$, and hence $\mathcal{Y}(\Lambda(0))=0$, which implies $|a_{1}(0)|=|a_{2}(0)|$.
+For any $t\in(0,T)$, we have $\Lambda(t)<\widehat{\Lambda}$.
+Thus $\mathcal{Y}(\Lambda(t))<0$, and hence $|a_{1}(t)|<|a_{2}(t)|$ for all $t\in(0,T)$.
+B.5 Proof of Proposition 5.6
+Throughout the proof, for $i=1,2$ and a fixed $p\in I$, we write $a_{i,p}$, $Q_{i,p}$, etc., to denote the corresponding functions in Propositions 5.3 and 5.4 corresponding to a given $p\in I$.
+We first show that $I$ is an open set.
+Indeed, define $\mathcal{D}(t,p)\triangleq Q_{1,p}(v_{1,p}(t))-Q_{2,p}(v_{2,p}(t))$.
+The crossing time $t^{*}(p)$ is characterized by $\mathcal{D}(t^{*}(p),p)=0$.
+From the proof of Proposition 5.4, we know that $\frac{\partial\mathcal{D}}{\partial t}(t^{*}(p),p)>0$.
+Hence, by the implicit function theorem, for each $p\in I$ there exists an open neighborhood in $(0,1)$ on which $t^{*}(\cdot)$ is uniquely defined and continuously differentiable.
+This implies that $I$ is an open set.
+Recall that $Q_{i,p}(x)=\dfrac{\delta_{i}(1-p)e^{-\delta_{i}\frac{x}{2}}}{p+(1-p)e^{-\delta_{i}\frac{x}{2}}}$, which is strictly decreasing in $x$.
+Define $U_{i}(p)\triangleq Q_{i,p}(0)=\delta_{i}(1-p)$ and $G_{i}(q)\triangleq\dfrac{(r_{0}+q)^{2}}{q(\delta_{i}-q)}$ for $p\in I$ and $q\in(0,\delta_{i})$.
+Then
+$\Lambda(t)=\int_{0}^{v_{i,p}(t)}(r_{0}+Q_{i,p}(x))^{2}\,\mathrm{d}x=2\int_{Q_{i,p}(v_{i,p}(t))}^{U_{i}(p)}G_{i}(q)\,\mathrm{d}q,$ (B.9)
+where the second equality arises from the change of variable $q=Q_{i,p}(x)$, which yields $\mathrm{d}x=\dfrac{2}{q(q-\delta_{i})}\,\mathrm{d}q$.
+For $p\in I$, set $\Lambda^{*}(p)\triangleq\Lambda(t^{*}(p))$ and $Q(p)\triangleq Q_{1,p}(v_{1,p}(t^{*}(p)))=Q_{2,p}(v_{2,p}(t^{*}(p)))$.
+Equation (B.9) then implies
+$\Lambda^{*}(p)=2\int_{Q(p)}^{U_{1}(p)}G_{1}(q)\,\mathrm{d}q=2\int_{Q(p)}^{U_{2}(p)}G_{2}(q)\,\mathrm{d}q.$ (B.10)
+Differentiating (B.10) w.r.t. $p$ and using $U_{i}^{\prime}(p)=-\delta_{i}$, we obtain
+$\dfrac{1}{2}(\Lambda^{*})^{\prime}(p)=-\delta_{1}G_{1}(U_{1}(p))-G_{1}(Q(p))Q^{\prime}(p)=-\delta_{2}G_{2}(U_{2}(p))-G_{2}(Q(p))Q^{\prime}(p).$ (B.11)
+Because $\delta_{1}-Q(p)>\delta_{2}-Q(p)>0$, it follows that $G_{1}(Q(p))<G_{2}(Q(p))$. Solving the second equality in (B.11) for $Q^{\prime}(p)$ and substituting into the first yields
+$\dfrac{1}{2}(\Lambda^{*})^{\prime}(p)=\dfrac{\delta_{2}G_{2}(U_{2}(p))G_{1}(Q(p))-\delta_{1}G_{1}(U_{1}(p))G_{2}(Q(p))}{G_{2}(Q(p))-G_{1}(Q(p))}.$ (B.12)
+Next, a direct computation shows that
+$\dfrac{\delta_{1}G_{1}(U_{1}(p))}{G_{1}(Q(p))}$ $=\dfrac{(r_{0}+\delta_{1}(1-p))^{2}}{p(1-p)\delta_{1}}\cdot\dfrac{Q(p)(\delta_{1}-Q(p))}{(r_{0}+Q(p))^{2}}$
+$=\dfrac{Q(p)}{p(1-p)(r_{0}+Q(p))^{2}}(r_{0}+\delta_{1}(1-p))^{2}\left(1-\dfrac{Q(p)}{\delta_{1}}\right).$
+Similarly,
+$\dfrac{\delta_{2}G_{2}(U_{2}(p))}{G_{2}(Q(p))}$ $=\dfrac{Q(p)}{p(1-p)(r_{0}+Q(p))^{2}}(r_{0}+\delta_{2}(1-p))^{2}\left(1-\dfrac{Q(p)}{\delta_{2}}\right).$
+Because $\delta_{1}>\delta_{2}$, it holds that $(r_{0}+\delta_{1}(1-p))^{2}>(r_{0}+\delta_{2}(1-p))^{2}$ and $1-\dfrac{Q(p)}{\delta_{1}}>1-\dfrac{Q(p)}{\delta_{2}}$.
+Therefore, $\dfrac{\delta_{1}G_{1}(U_{1}(p))}{G_{1}(Q(p))}>\dfrac{\delta_{2}G_{2}(U_{2}(p))}{G_{2}(Q(p))}$.
+Consequently, $(\Lambda^{*})^{\prime}(p)<0$.
+Recall that $\Lambda^{*}(p)=\Lambda(t^{*}(p))$.
+Differentiating both sides w.r.t.
+$p$ yields
+$(\Lambda^{*})^{\prime}(p)=\Lambda^{\prime}(t^{*}(p))\frac{\mathrm{d}t^{*}}{\mathrm{d}p}(p)=-|\lambda(t^{*}(p))|^{2}\frac{\mathrm{d}t^{*}}{\mathrm{d}p}(p),$
+Therefore,
+$\frac{\mathrm{d}t^{*}}{\mathrm{d}p}(p)=-\frac{(\Lambda^{*})^{\prime}(p)}{|\lambda(t^{*}(p))|^{2}}>0\quad\text{for all }p\in I.$
+B.6 Proof of Proposition 5.7
+We retain the notation from the proof of Proposition 5.6.
+Define $\mathcal{D}(t,p_{1},p_{2})\triangleq Q_{1,p_{1}}(v_{1,p_{1}}(t))-Q_{2,p_{2}}(v_{2,p_{2}}(t))$.
+By definition, the crossing time $t^{*}(p_{1},p_{2})$ satisfies $\mathcal{D}(t^{*}(p_{1},p_{2}),p_{1},p_{2})=0$.
+From the proof of Proposition 5.4, we know that $\frac{\partial\mathcal{D}}{\partial t}(t^{*}(p_{1},p_{2}),p_{1},p_{2})>0$.
+Therefore, by the implicit function theorem, for any $(p_{1},p_{2})\in O$, there exists an open neighborhood in $(0,1)^{2}$ on which $t^{*}(p_{1},p_{2})$ is uniquely defined and continuously differentiable.
+In particular, $O$ is an open set and $t^{*}$ is differentiable in $O$.
+For any $(p_{1},p_{2})\in O$, define $\Lambda^{*}(p_{1},p_{2})\triangleq\Lambda(t^{*}(p_{1},p_{2}))$ and $Q(p_{1},p_{2})\triangleq Q_{1,p_{1}}(v_{1,p_{1}}(t^{*}(p_{1},p_{2})))=Q_{2,p_{2}}(v_{2,p_{2}}(t^{*}(p_{1},p_{2})))$.
+Analogous to (B.10), we obtain
+$\Lambda^{*}(p_{1},p_{2})=2\int_{Q(p_{1},p_{2})}^{U_{1}(p_{1})}G_{1}(q)\,\mathrm{d}q=2\int_{Q(p_{1},p_{2})}^{U_{2}(p_{2})}G_{2}(q)\,\mathrm{d}q.$ (B.13)
+Differentiating (B.13) w.r.t. $p_{1}$, and using $\dfrac{\partial U_{1}}{\partial p_{1}}=-\delta_{1}$ and $\dfrac{\partial U_{2}}{\partial p_{1}}=0$, we obtain
+$\dfrac{1}{2}\dfrac{\partial\Lambda^{*}}{\partial p_{1}}=-\delta_{1}G_{1}(U_{1}(p_{1}))-G_{1}(Q(p_{1},p_{2}))\dfrac{\partial Q(p_{1},p_{2})}{\partial p_{1}}=-G_{2}(Q(p_{1},p_{2}))\dfrac{\partial Q(p_{1},p_{2})}{\partial p_{1}}.$ (B.14)
+As shown in the proof of Proposition 5.6, $\delta_{1}>\delta_{2}$ implies $G_{1}(Q(p_{1},p_{2}))<G_{2}(Q(p_{1},p_{2}))$.
+Solving for $\frac{\partial Q(p_{1},p_{2})}{\partial p_{1}}$ from the second equality of (B.14) and substituting into the first yields
+$\dfrac{1}{2}\dfrac{\partial\Lambda^{*}}{\partial p_{1}}=\dfrac{-\delta_{1}G_{1}(U_{1}(p_{1}))G_{2}(Q(p_{1},p_{2}))}{G_{2}(Q(p_{1},p_{2}))-G_{1}(Q(p_{1},p_{2}))}<0.$
+Applying the Chain Rule to the identity $\Lambda^{*}(p_{1},p_{2})=\Lambda(t^{*}(p_{1},p_{2}))$ yields
+$\dfrac{\partial\Lambda^{*}}{\partial p_{1}}=\Lambda^{\prime}(t^{*})\dfrac{\partial t^{*}}{\partial p_{1}}=-|\lambda(t^{*})|^{2}\dfrac{\partial t^{*}}{\partial p_{1}}.$
+Consequently, $\frac{\partial t^{*}}{\partial p_{1}}>0$. Similarly, we obtain $\frac{\partial t^{*}}{\partial p_{2}}<0$.
+B.7 Proof of Proposition 5.9
+From Lemma A.5, we know that $h_{1}(x)<h_{2}(x)$ for all $x\geq 0$.
+it follows that $H_{1}(y)>H_{2}(y)$ for all $y>0$, and in particular, $H_{1}(\infty)\geq H_{2}(\infty)>\Lambda(0)$.
+This guarantees that the deterministic equilibrium strategy associated with each $\bm{R}_{i}$} exists and is unique.
+Moreover, the inverse functions of $H_{i}$ ($i=1,2$) satisfy $H_{1}^{-1}(z)<H_{2}^{-1}(z)$ for all $z>0$.
+From Theorem 3.2(1), we have $v_{i}(t)=H_{i}^{-1}(\Lambda(t))$ ($i=1,2$) .
+Consequently, the monotonicity of $H_{i}^{-1}$ ($i=1,2$) implies $v_{1}(t)\leq v_{2}(t)$.
+By (2.5), it holds that $|a_{i}(t)|=h_{i}(v_{i}(t))|\lambda(t)|$ ($i=1,2$).
+Using Lemma A.5, we further obtain $h_{1}(v_{1}(t))\leq h_{2}(v_{1}(t))\leq h_{2}(v_{2}(t))$.
+Therefore, $|a_{1}(t)|\leq|a_{2}(t)|$.
+If $\lambda(t)\neq 0$, then $v_{1}(t)>0$ and $h_{1}(v_{1}(t))<h_{2}(v_{1}(t))\leq h_{2}(v_{2}(t))$, which yields the strict inequality: $|a_{1}(t)|<|a_{2}(t)|$.
+B.8 Proof of Proposition 5.11
+Let $h(x)$ and $v(t)$ (resp.
+$h_{i}(x)$ and $v_{i}(t)$) denote the functions associated with $\bm{R}$ (resp.
+$\bm{R}_{i}$).
+From the definition of $h$ and the fact that $\bm{R}_{1},\cdots,\bm{R}_{n}$ are i.i.d., we have
+$\frac{1}{h(x)}=\frac{\tilde{\mathbb{E}}[\bm{R}e^{-\frac{1}{2}\bm{R}x}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{R}x}]}=\frac{\tilde{\mathbb{E}}\left[\displaystyle\left(\sum_{i=1}^{n}w_{i}\bm{R}_{i}\right)\prod_{k=1}^{n}e^{-\frac{1}{2}w_{k}\bm{R}_{k}x}\right]}{\displaystyle\tilde{\mathbb{E}}\left[\prod_{k=1}^{n}e^{-\frac{1}{2}w_{k}\bm{R}_{k}x}\right]}$
+$=$ $\frac{\displaystyle\sum_{i=1}^{n}w_{i}\tilde{\mathbb{E}}\left[\bm{R}_{i}e^{-\frac{1}{2}w_{i}\bm{R}_{i}x}\prod_{k\neq i}e^{-\frac{1}{2}w_{k}\bm{R}_{k}x}\right]}{\displaystyle\prod_{k=1}^{n}\tilde{\mathbb{E}}\left[e^{-\frac{1}{2}w_{k}\bm{R}_{k}x}\right]}=\sum_{i=1}^{n}w_{i}\frac{\tilde{\mathbb{E}}[\bm{R}_{i}e^{-\frac{1}{2}w_{i}\bm{R}_{i}x}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}w_{i}\bm{R}_{i}x}]}=\sum_{i=1}^{n}w_{i}\frac{1}{h_{1}(w_{i}x)}.$ (B.15)
+By Lemma A.1, we have $\dfrac{1}{h_{1}(w_{i}x)}\geq\dfrac{1}{h_{1}(x)}$ for any $w_{i}\in(0,1)$ and $x\geq 0$, which combined with (B.15) yields $h(x)\leq h_{1}(x)$ and $H(\infty)\geq H_{1}(\infty)>\Lambda(0)$.
+Consequently, the equilibrium strategy associated with $\bm{R}$ also exists and is unique.
+Analogous to the argument in Proposition 5.9, we further obtain $v(t)\leq v_{1}(t)$.
+Because $h(x)$ is increasing in $x$, it follows that
+$|a(t)|=h(v(t))|\lambda(t)|\leq h(v_{1}(t))|\lambda(t)|\leq h_{1}(v_{1}(t))|\lambda(t)|=|a_{1}(t)|.$
+If $\bm{R}_{1}$ is non-degenerate, then by Lemma A.1, $h_{1}$ is strictly increasing, which in turn yields the strict inequality $h(x)<h_{1}(x)$ for any $x>0$.
+Combined with the condition $\lambda(t)\neq 0$, we obtain $|a(t)|<|a_{1}(t)|$.
+B.9 Proof of Proposition 5.13
+Let $\tilde{h}_{n}(\cdot)$ and $\tilde{v}_{n}(\cdot)$ be the functions associated with the risk aversion $\bm{Y}_{n}$. From (B.15), we have
+$\tilde{h}_{n}(x)=h_{1}\left(\frac{x}{n}\right)\leq h_{1}(x).$
+Then we have $H(\infty)\geq H_{1}(\infty)>\Lambda(0)$ and the deterministic equilibrium strategy of $\bm{Y}_{n}$ exists and is unique.
+Moreover, by Lemma A.5, for any fixed $x\geq 0$, the sequence $\{\tilde{h}_{n}(x)\}_{n\geq 1}$ is decreasing in $n$.
+Because $\tilde{H}_{n}(y)=\displaystyle\int_{0}^{y}\tilde{h}_{n}^{-2}(z)dz$ and $\tilde{v}_{n}(t)=\tilde{H}_{n}^{-1}(\Lambda(t))$, the monotonicity of $\tilde{h}_{n}$ implies that $\tilde{v}_{n}(t)$ is decreasing in $n$.
+Consequently, for all $t\in[0,T)$,
+$\tilde{h}_{n+1}(\tilde{v}_{n+1}(t))\leq\tilde{h}_{n}(\tilde{v}_{n+1}(t))\leq\tilde{h}_{n}(\tilde{v}_{n}(t)),$
+which yields
+$|\tilde{a}_{n+1}(t)|\leq|\tilde{a}_{n}(t)|\quad\text{ for all }t\in[0,T).$
+By the strong law of large numbers, $\bm{Y}_{n}\to\mu$ almost surely as $n\to\infty$.
+Hence, $\bm{Y}_{n}$ converges in distribution to the degenerate random variable $\bm{Y}\equiv\mu$.
+Furthermore, the assumptions of Theorem 3.11 are satisfied because $\tilde{\mathbb{E}}[\bm{Y}_{n}]=\mu$ for all $n$.
+We therefore conclude that
+$\lim_{n\to\infty}|\tilde{a}_{n}(t)|=\frac{|\lambda(t)|}{\mu}.$
+Combining this limit with the monotonicity established above yields (5.2).